### v2

# Scalable Private Search with Wally

Abstract
This paper presents Wally  a private search system that supports efficient semantic and keyword search queries against large databases. When sufficient clients are making the queries, Wally performance is significantly better than previous systems. In previous private search systems, for each client query, the server must perform at least one expensive cryptographic operation per database entry. As a result, performance degraded proportionally with the number of entries in the database.
In Wally we get rid of this limitation. Specifically, for each query the server performs cryptographic operations only against a few database entries. We achieve these results by requiring each client to add a few fake queries, and sends each query via an anonymous network to the server at independently chosen random instants. Additionally, each client also uses somewhat homomorphic encryption (SHE) to hide whether a query is real or fake, Wally provides $(\epsilon,\delta)$-differential privacy guarantee, which is an accepted standard for strong privacy.
The number of fake queries each client makes depends inversely on the number of clients making queries. Therefore, the fake queries’ overhead vanishes as the number of clients increases, enabling scalability to millions of queries and large databases. Concretely, Wally can serve $8$M requests at a rate of 3,000 queries per second. That is around 60x higher than the state-of-the-art scheme.
1 Introduction
Consider a scenario where a client holding a search query wants to retrieve relevant information from a large database hosted on a server. This scenario reflects the flow of current search engines such as Google, Bing, and DuckDuckGo. These search engines improve the quality of our lives by promptly providing helpful information. Nevertheless, they require learning the client’s query to respond accurately to each query, hence providing no query privacy. In some situations, learning the query could reveal sensitive and personal information about the client. For example, consider the queries “High blood pressure with late-stage HIV?”, “Closest diabetes clinic from Albuquerque”, and “Best divorce lawyers ”. Revealing these queries to search engines could be damaging to the client. However, even if the queries do not appear sensitive, they may allow the server to learn personal information like medical information, marital status, sexuality, etc. Therefore, enabling query privacy is critical for search engines.
Private search engines opt to provide query privacy. However, these systems have significantly higher computation and communication overhead than the state-of-the-art insecure solutions. For example, in Tiptoe [HDCZ23], a state-of-the-art private search engine, a single client query is 21 MB and requires 339 core-seconds of server computing against a database of $400$ million entries. The high computation is because, for each client query, the server must scan the entire database; otherwise, it will learn which database entries the client is not interested in. Similarly, the client must send cryptographic material for each database entry, which results in high communication cost [^1] In reality, these systems trade off response size with request size by using a standard database clustering trick. We detail this trade off in a later section.. Additional limitation of Tiptoe is that it requires each client to store a database-dependent state of around 50 MB for each client query.
In this paper, we present Wally, a privacy preserving search engine that efficiently scales to performance constraints of large-scale search engines. Wally supports semantic search, where the client wants to fetch the most relevant documents to a query, and keyword search, where the client wants to retrieve documents associated with a private keyword. To provide privacy, our key observation is that in search engines, there are always enough clients making queries. We exploit this observation to gain privacy by devising a way to hide a particular query in the crowd of queries. In our protocol, the server computation and the communication get amortized over the number of clients making the requests; therefore, when the number of clients is large, the per-query cost becomes minimal.
Wally achieves these results by relaxing its privacy notion to $(\epsilon,\delta)$-differential privacy. In Wally the server learns noised differentially private histogram of anonymized queries from all clients. We emphasize that Tiptoe, is fully oblivious, i.e., the server learns nothing besides that a particular client is making a query. Therefore, the privacy guarantee of Tiptoe is stronger than Wally  however differential privacy is an accepted standard for strong privacy [DR+14]. Specifically, in Wally every client is guaranteed that all the statistical inferences are insensitive to their data. In other words, their participation does not significantly alter the server’s view, except with probability $\delta$, and thus provides strong privacy protection.
At a high level, Wally works in epochs. Within each epoch, a fixed number of users make queries. The epoch size is picked to ensure enough clients are present to make queries, and their queries will be processed by the end of an epoch. The server database is divided into $C$ clusters. Each query consists of a relevant cluster index, and the server response is all the entries in the cluster. The queries arrive at the server via an anonymous network, ensuring all identifiable information like IP address is removed from the queries. The server then generates a response to each query independently and sends it back to the respective client via an anonymous network.
Even though queries are anonymized, they do not provide query privacy. The server can infer information about the client query by exploiting the access pattern (if a particular cluster is accessed) or arrival time (if queries arrive earlier than the others). To prevent this leakage, Wally implements the following two changes at the client.
To ensure that the server does not learn information by observing the access patterns, in Wally each client also makes fake queries. For each fake query, the client picks a randomly selected bucket. This ensures that the server cannot differentiate between real and fake queries. The number of fake queries is sampled from a particular distribution to ensure the revealed histogram of queries is $(\epsilon,\delta)$-differentially private. In expectation, each client makes $O(\frac{C\Delta\log(1/\delta)}{U\epsilon})$ fake queries where $U$ is the number of clients within an epoch, and $\Delta$ is the maximum number of queries each client can make in an epoch. Note that increasing the number of clients decreases the average number of fake queries. The overhead due to fake queries can be traded with privacy because more fake queries are required as $\epsilon$ decreases. We emphasize that fake queries do not affect the accuracy or correctness seen by any client, as the client can discard the responses due to these queries.
Wally also ensures that arrival time and order of the queries leak no information about the client to the server. To achieve this, instead of making a query instantly, the clients make queries at random one second slots within an epoch. Also, an anonymous network collects the requests received within a particular slot and sends them as a batch to the server.
Anonymizing the queries, adding fake queries that are indistinguishable from real ones, and randomizing query order and time guarantees differential privacy.
As mentioned above, for each query, the server response contains all the entries in the cluster. For large clusters, this would result in very high response overhead. In Wally  we utilize lattice-based, somewhat homomorphic encryption (SHE) to reduce the response overhead. Specifically, for semantic search, each query additionally includes a semantic information (dependent on the client’s input) encrypted under SHE. The server computes the distance function between the encrypted information and the cluster entries under SHE and returns encrypted scores. This reduces the response because the size of encrypted scores is significantly smaller than the entries. For a keyword search, the client includes encrypted PIR query, and the server runs an SHE-based keyword private information retrieval (PIR) protocol within the cluster, which returns data associated with the client keyword. To ensure that the server cannot differentiate between real and fake queries, for every fake query the client includes encrypted semantic, or keyword queries based on fake data. We propose various optimizations that significantly improve performance for both kinds of searches.
Our contributions are as follows:
  • We describe Wally’s design, which enables privacy-preserving large-scale search engines. Its novelty stems from the combination of anonymity, fake queries added by the clients, and the client’s use of SHE to encrypt their queries.
  • We also describe the implementation of Wally, showing efficiency and scalability along with a robust privacy guarantee. Generally, SHE computations have a high performance overhead. To achieve desired performance constraints, we propose several optimizations to SHE computation (distance computation and PIR) that could be of independent interest.
  • We evaluate Wally for private image search, demonstrating both strong privacy and low overhead.
Limitation.
Wally requires decent number of clients always present in an epoch to make queries. We argue that this assumption aligns with current search engines, for example a hundred of millions of people make queries in Google search daily [Fle24]. We remark that for systems with few users, overhead due to fake queries could get quite high, therefore for these systems fully oblivious protocols like Tiptoe will be suitable.
2 Background
Let $[n]:=\{1,\dots,n\}$ be the set of the first $n$ natural numbers starting from one. Matrices and vectors are represented as capital boldface letters and capital lowercase, respectively, e.g., $\mathbf{M}$ and $\mathbf{v}$. The entry-wise, Hadamard, product of two vectors is denoted as $\mathbf{v}\odot\mathbf{v}^{\prime}$. All logarithms are the natural log unless noted otherwise. We denote a random variable being sampled from a probability distribution $D$ as $a\leftarrow D$. All distributions in this work are over finite sets and $a\leftarrow S$ denotes uniformly sampling from $S$ when $S$ is a discrete set. We denote the integers as $\mathbb{Z}$ and the integers modulo a positive integer $Q$ as $\mathbb{Z}_{Q}$. Let $\mathcal{NB}(r,p)$ denote the negative binomial distribution parameterized by $r,p$. Its probability mass function is $Pr(X=k)=\binom{k+r-1}{k}(1-p)^{k}p^{r}$, $X\sim\mathcal{NB}(r,p)$, for every non-negative integer $k$. An important fact about the negative binomial distribution is that it is infinitely divisible, Definition 1.
Definition 1.
A distribution $\mathcal{D}$ is infinitely-divisible if for all $n\in\mathbb{N}$, it can be expressed as the sum of $n$ iid variables. That is, there exists a distribution $\mathcal{D}_{n}$ such that $\mathcal{X}_{1}+\cdots+\mathcal{X}_{n}\sim\mathcal{D}$ and $\mathcal{X}_{i}\sim\mathcal{D}_{n}$.
2.1 Homomorphic encryption
We use the BFV [Bra12, FV12] SHE scheme [^2] BFV is also a fully homomorphic scheme, which can homomorphically compute any function, with large enough parameter sets. We only focus on its simpler SHE parameter regime in our paper. where each plaintext is a polynomial in $\mathbb{Z}_{t}[X]/(X^{n}+1)$ (with degree at most $n-1$) and $t$ is a prime or prime power. Wally can also be implemented with the BGV [BGV14] SHE scheme with similar performance. Both schemes are post-quantum SHE schemes based on the ring-learning-with-errors (RLWE) problem [Reg05, LPR10]. A BFV ciphertext is a pair of polynomials $\mathsf{ct}=(c_{0},c_{1})\in R_{Q}^{2}$ reduced modulo $Q$ and $X^{n}+1$ for a power of two $n$, $R_{Q}:=\mathbb{Z}_{Q}[X]/(X^{n}+1)$. The number $n$ is a ring dimension, $n\in\{2^{12},2^{13}\}$ in our use cases, and $Q$ is a ciphertext modulus with $\log_{2}Q=83$ or $\log_{2}Q=155$ since we use 128-bit secure post-quantum security parameters [ADPS16].
A BFV ciphertext satisfies its decryption formula: $c_{0}+c_{1}s\mod Q=\lfloor Q/t\rceil m+e$ where $s$ is a secret key polynomial with small coefficients, in $\{0,\pm 1\}$, $e$ is small error polynomial, $t$ is the plaintext modulus, and $m\in\mathbb{Z}_{t}[X]/(X^{n}+1)$ is the plaintext polynomial. This error polynomial’s entries grow as we compute more homomorphic operations on the ciphertext and we can only decrypt if $\|e\|_{\infty}<Q/2t$. We call $\log_{2}(Q)-\log_{2}(2t)$ the parameter’s noise budget, measured in bits.
SHE parameters.
We choose parameters by first choosing a homomorphic computation, plaintext modulus, and a security parameter. This leads to a ring dimension and modulus $(n,Q)$ which 1) satisfies the security requirement and 2) allows for enough noise budget to allow decryption after homomorphic computation. Larger $Q$ gives more noise budget, but it also necessitates a larger dimension $n$ for security. The ratio of the ring dimension to the modulus bits, $n/\log_{2}Q$, stays roughly constant for a fixed security level.
Vectorized HE.
Vectorized plaintext spaces are used via polynomial interpolation: $X^{n}+1\mod t$ splits when $t$ is a prime satisfying $t=1\mod 2n$. In the vectorized case, each ciphertext encrypts a vector $\mathbf{v}\in\mathbb{Z}_{t}^{n}$ and the homomorphic operations are single-instruction multiple-data (SIMD) instructions. Vectorized SHE is crucial for applications requiring homomorphic linear algebra.
SHE operations and costs.
BFV’s $\mathsf{Eval}$ calls five subroutines:
  • $\mathsf{PtCtAdd}(\mathsf{ct},\mathbf{v}^{\prime})$ returns a ciphertext encrypting $\mathbf{v}+\mathbf{v}^{\prime}$ if $\mathsf{ct}=\mathsf{Enc}(\mathbf{v})$.
  • $\mathsf{CtCtAdd}(\mathsf{ct},\mathsf{ct}^{\prime})$ returns a ciphertext encrypting $\mathbf{v}+\mathbf{v}^{\prime}$ if $\mathsf{ct}=\mathsf{Enc}(\mathbf{v})$ and $\mathsf{ct}^{\prime}=\mathsf{Enc}(\mathbf{v}^{\prime})$
  • $\mathsf{PtCtMult}(\mathsf{ct},\mathbf{v}^{\prime})$ returns a ciphertext encrypting $\mathbf{v}\odot\mathbf{v}^{\prime}$ if $\mathsf{ct}^{\prime}=\mathsf{Enc}(\mathbf{v})$.
  • $\mathsf{CtCtMult}(\mathsf{ct},\mathsf{ct}^{\prime})$ returns a ciphertext encrypting $\mathbf{v}\odot\mathbf{v}^{\prime}$ if $\mathsf{ct}=\mathsf{Enc}(\mathbf{v})$ and $\mathsf{ct}^{\prime}=\mathsf{Enc}(\mathbf{v}^{\prime})$.
  • $\mathsf{CtRotate}(\mathsf{ct},r)$ for $r\in[0,n/2)$ returns a ciphertext encrypting
    (((( $\mathsf{Rot}(\mathbf{v}),\mathsf{Rot}(\mathbf{\bar{v}}))\in\mathbb{Z}_{p}^{n/2}\times\mathbb{Z}_{p}^{n/2}$
    where $\mathsf{Rot}(\mathbf{v})$ ($\mathsf{Rot}(\mathbf{\bar{v}})$) is $\mathbf{v}$ ($\mathbf{\bar{v}}$) cyclically shifted to the left by r positions, if $\mathsf{ct}$ originally encrypted $\mathbf{v}^{\prime}=(\mathbf{v},\mathbf{\bar{v}})$. We can homomorphically swap $(\mathbf{v},\mathbf{\bar{v}})$ with the same operation, called “conjugation”.
Ciphertext rotation and ciphertext multiplication require an evaluation key since the secret changes with each rotation and ciphertext multiplication. This operation is called “key-switching” for rotations and “relinearization” for ciphertext multiplications.
Table 1 has concrete run times and noise growth for BFV. In general, addition and plaintext multiplication are efficient and addition has minimal noise growth. The latter has larger growth proportional to $t$. Ciphertext-ciphertext multiplication has large noise growth and is somewhat inefficient since it requires multiplying the ciphertext polynomials over the integers then rounding down back to $\mod Q$. Ciphertext rotation adds little noise but it is somewhat slow due to key-switching. Therefore, minimizing the number of ciphertext multiplications and ciphertext rotations is crucial for scalable applications.
Operation Time (ms) Noise added (bits)
$\mathsf{CtCtAdd}$ $070$absent0$\approx$ 0≈ 0
$\mathsf{PtCtMult}$ $0.09$ $22$
$\mathsf{CtCtMult}$ $12.1$ $20$
$\mathsf{CtRotate}$ $3.6$ $\approx 0$
Table 1: Experimental computation cost and noise growth of each BFV homomorphic operation. The polynomial degree $n$ is $8192$, the ciphertext modulus $Q$ has $150$ bits, and the plaintext modulus $t$ has $20$ bits. Time costs are measured with the SEAL library [CLP17] version 3.7.2 on a single core in a AWS t2.2xlarge instance.
Differential privacy
We use the standard notion of differential privacy (DP) [DMNS06, DKM+06]: a randomized mechanism $M$ is $(\epsilon,\delta)$-differentially private, if for all $\epsilon\geq 0$ and $\delta\in[0,1]$, any two datasets $X,X^{\prime}$ differing on one element, and for any subset $S$ of possible transcripts output by $M$, $\Pr[M(X)\in S]\leq e^{\epsilon}\cdot\Pr[M(X^{\prime})\in S]+\delta$ where the probability is over the randomness of $M$. Let $\zeta$ be an output of a mechanism $M$, then the privacy loss in observing $\zeta$ between two datasets $X,X^{\prime}$ as input is
$\log\left(\frac{\Pr[M(X)=\zeta]}{\Pr[M(X^{\prime})=\zeta]}\right).$
Essentially, standard DP ensures that neighboring datasets have a privacy loss of at most $\epsilon$ with probability at least $1-\delta$ for all observations. We use basic composition in DP [Vad17, Lemma 2.3].
Lemma 1.
If $M_{1}(X),\dots,M_{k}(X)$ are each $(\epsilon,\delta)$-DP with independently sampled randomness, then $(M_{1}(X),\dots,M_{k}(X))$ is $(k\epsilon,k\delta)$-DP.
Anonymous networks.
Anonymous networks (ANs), also called anonymization networks and anonymous communication protocols, are protocols which offer user anonymity by using cryptographic techniques such as onion routing [GRS99], or mixnets [Cha81]. Throughout the paper, we assume that ANs 1) strip all identifying information from incoming messages from clients, and 2) permute all received messages within a batch of messages from clients before sending them to the server.
Private information retreival.
Though the paper is mostly presented for approximate nearest neighbor retrieval, Wally applies to exact retrieval as well by using keyword PIR for its homomorphic computation. Private information retrieval (PIR) protocols are cryptographic client-server protocols where a server holds a large database and the client obliviously retrieves an entry from the database. PIR protocols can be for index retrieval or retrieval by keyword [CGN98]. In this paper, we focus on single-server PIR protocols [KO97] since we can shard the database and the client can run Wally to anonymously retrieve entries and hide its downloaded entry on the shard. We use and optimize PIR protocols based on SHE with minimal preprocessing, such as [ACLS18] and [ALP+21]. There is a notion of differentially-private PIR in the multi-server setting [TDG16, AIVG22], where the database is stored across non-colluding servers while guaranteeing differential privacy. This is in contrast to Wally which stores the database on one server or servers controlled by a single entity.
3 System Design
Figure 1: High-level overview of Wally. Throughout an epoch, clients sends encrypted real and fake queries at random time slots to an anonymization network which “washes” queries of identifying information.
A high-level design of Wally is in Figure 1. It consists of three entities: the server, the clients, and the anonymization network. Wally runs in epochs of time; within each epoch, a fixed number of clients make queries. Wally does not require any synchronization across clients. Each client can decide to participate or skip an epoch. However, we require a minimum of clients stay online throughout an epoch and that a participating client can only make at most $\Delta$ queries within each epoch. Overall, privacy loss is the sum of the loss due to each epoch in which the client has participated. The correctness guarantee is that all clients’ queries are processed by the end of the epoch. Wally’s privacy notion also requires that each client will make at most $M$ queries in the system’s lifetime; therefore, a can participate at max $M$ epochs.
Each epoch is divided into second-long time slots, and privacy is guaranteed by the following:
  • Hiding the real queries of clients among fake queries (differential privacy).
  • Hiding the content of real and fake queries (somewhat homomorphic encryption).
  • Removing the identifiable information from each query and shuffling the batch of queries received within a second (anonymization network).
  • Hiding the information revealed by queries order and time of arrival (random query schedule)
Within each epoch, Wally performs a semantic search that outputs the most relevant results to client queries, similar to insecure search engines.
Embedding-based search.
Wally utilizes (vector) embeddings for semantic search. Embeddings are a machine learning-based technique that map unstructured objects like documents, images, and videos to $d$-dimensional vectors of floats. These embeddings retain the semantic relationship between objects by placing similar objects close together in a vector space. Hence, it is used to find relevant entries in search systems. Embeddings are usually much smaller than the objects themselves, which results in a significant reduction in client-to-server communication. Embeddings can be generated for different data types, including text, image, and video. Therefore, Wally can support all these data types. Wally is compatible with all the recent embedding models, such as BERT [DCLT19].
In the following sections, we introduce the insecure version of Wally. We then highlight the privacy issues in this version. Finally, we discuss how we overcome these privacy issues.
3.1 Insecure Search Pipeline
Once the client query and the server index are represented as embeddings, Wally uses an approximate nearest neighbor algorithm to find the documents most relevant to the client’s query. Our two-step approximate nearest neighbor protocol has the following flow:
  • Server initialization. At initialization, the server divides the entries into $K$ clusters using $K$-means clustering. A heuristic-based technique partitions data into $K$ disjoint clusters $C_{1},\cdots,C_{K}$ and outputs these clusters with their centroid embeddings $c_{1},\cdots,c_{K}$. The server sends the centroid embeddings to the client.
  • Pick nearest centroid. To make a query, the client first finds the cluster nearest to the query embedding. For this, the client computes a cosine similarity between the query and each centroid embedding and picks a cluster whose centroid has high cosine similarity.
  • Server computation. The client then sends the nearest cluster ID $C^{\prime}$ and query embedding $q^{\prime}$ to the server. The server computes cosine similarity between the query embedding and each embedding in cluster $C^{\prime}$. The server returns the similarity scores to the client.
  • Pick nearest entry. The client sorts these scores and picks the entry closest to the query.
  • Metadata retrieval. The client then fetches the metadata associated with the closest entry from the server.
The clustering technique provides a trade-off between performance and accuracy; generally, increasing the number of clusters improves the performance (reduces server computation and response size). However, it degrades the accuracy because there is a higher chance that a potential nearest neighbor is in a different cluster than the cluster with the closest centroid. Therefore, finding a balance is important for the overall usability of the system.
The above protocol is not private. To ensure that the protocol does not reveal any information about the client, it must hide the following information from the server:
[…]
  • Nearest cluster. The client reveals to the server the nearest cluster id. Recall that clustering puts highly correlated documents together in a cluster. Therefore, cluster id reveals the kind of data the client holds. For example, if the client query is an image, the cluster ID might reveal whether the image is a landmark, e.g., a castle or a bridge.
  • Closest entry index. To fetch the relevant metadata, the client sends the index of the closest entry to the server. Learning this could reveal server information about client queries.
We show how we hide the above data below and in Section 5.
Figure 2: A single slice computation in Wally using SHE operations. The server only require aligning the query once for all the slices.
3.2 Hiding the Query Embedding with SHE
To hide a query embedding, we rely on vectorized, RLWE-based SHE introduced in Section 2. The client query is an encrypted vector, and the server database is encoded as an array of plaintext vectors. The output is a single response ciphertext with at most $n$ encrypted similarity scores, where $n$ is the RLWE polynomial dimension.
Server database.
The server divides the database, of size $N$, into $K$ clusters each with at most $N^{\prime}$ embeddings. For simplicity, we assume that for a given cluster $C_{i}$, the dataset is represented as a cube $D_{C_{i}}$ of dimensions $N^{\prime}/d*d*d$, recall that $d$ is the embedding size. Each cube consists of $N^{\prime}/d$ slices, each of which is a matrix of dimensions $d*d$, and each column within a slice is an embedding. Denote elements of $i$’th embedding as $e_{i}=[e_{i}^{0},\cdots,e_{i}^{d-1}]$. For now, we assume that each diagonal of a slice is a separate plaintext vector of size $n=d$. That is each slice is encoded as $d$ plaintexts vectors and $j$’th vector consists of entries $p_{j}=[e_{0}^{j},e_{1}^{j+1},e_{2}^{j+2},\cdots,e_{d-1}^{j+d-1}]$. This means each plaintext vector consists of a single index from each embedding in a slice in increments of one. The client encrypts the query embedding $\hat{q}=[q^{0},q^{1},\cdots,q^{d-1}]$ in a single SHE ciphertext.
Dot product computation.
Recall, from Section 2, for cosine similarity, the server needs to compute a dot-product between two embeddings. To achieve that, the server must align the query to ensure that elements of a query vector only get multiplied with corresponding element of plaintext vectors, i.e., every element-wise multiplication should be like $e_{j}^{i}*q^{i}$. The server will use homomorphic rotation to move the elements within the encrypted query vector to the correct alignment. Specifically, for a given slice, starting with the first plaintext vector $j=0$, the server performs vectorized ciphertext-plaintext multiplication between the query ciphertext and the first plaintext vector $j=0$. This results in encrypted $p_{0}*\hat{q}=[e_{0}^{0}*q^{0},e_{1}^{1}*q^{1},e_{2}^{2}*q^{2},\cdots,e_{d-1}^{d-1}*q^{d-1}]$. No rotation is required for first multiplication because the query is already aligned with the first vector. The elements of the second plaintext vector $j=1$ are $p_{1}=[e_{0}^{1},e_{1}^{2},e_{2}^{3},\cdots,e_{d-1}^{0}]$. Hence, the server cannot directly multiply it with the query vector. The server first homomorphically rotates the query vector one slot to the left, resulting in query vector $\hat{q}^{1}=[q^{1},q^{2},\cdots,q^{0}]$ then multiplies the result: $p_{1}*\hat{q}^{1}=[e_{0}^{1}*q^{1},e_{1}^{2}*q^{2},e_{2}^{3}*q^{3},\cdots,e_{d-1}^{0}*q^{0}]$. The server then repeats this for all the remaining plaintext vectors in a slice: rotating the query vector in increments of one and multiplying it with the plaintext vector. By the end the server has $p_{0}*\hat{q},p_{1}*\hat{q}^{1},\cdots,p_{d-1}*\hat{q}^{d-1}$ encrypted multiplications. The server then sums all of these vectors together, resulting in a single encrypted vector in which each element is a dot product between the client query vector and one of the embeddings in the slice. Figure 2 shows an example server computation for a single slice for $d=4$. This diagonal method is to minimize homomorphic rotations [HS14, HS18].
Database and query packing.
In the above description, we assume that the polynomial dimension $n$ equals embedding dimension $d$. However, $n$ is often significantly larger than $d$. In this setting, each plaintext vector could hold multiple diagonals. Our scheme packs diagonals across the slices in a single plaintext vector. Concretely, the $j$-th plaintext vector in a cube holds $(j\mod d)$-th diagonals of $\lfloor{n/d}\rfloor$ slices. In all the practical databases that we have considered, the number of slices per cube $N^{\prime}/d$ is smaller than $\lfloor{n/d}\rfloor$. Therefore, by using packing, the entire cube can be encoded with only $d$ plaintext vectors.
The client query encrypts a plaintext vector of $\lfloor{n/d}\rfloor$ repetitions of the query embedding. This query still allows the server to align the entries of query embeddings by homomorphically rotating left in increments of one. The number of ciphertext-plaintext multiplications is now $\lceil N^{\prime}/n\rceil*d$ while the response consists of $\lceil N^{\prime}/n\rceil$ ciphertexts. Again, for the databases we consider, $\lceil N^{\prime}/n\rceil\approx 1$ results in $d$ multiplications, and each response is a single ciphertext.
Recall from Section 2 that homomorphic rotation is a costly operation. Even after packing, the server still performs $d$ rotations. Looking ahead to Section 6, we explain how we reduce the number of rotations to $2\sqrt{d}$ by using the standard the baby-step giant-step optimization [HS18].
Float-to-integer encoding.
Recall from Section 2 that vectorized SHE supports operations over vectors of integers while embeddings are vectors of floats. Wally converts each float into an integer by multiplying with the appropriate scaling factor and only taking the integer part of the result. The choice of scaling factor depends on the size of SHE plaintext modulus $t$ and the required precision.
Figure 3: Differential privacy related changes to hide real queries. Changes are highlighted in gray.
3.3 Hiding the nearest cluster
By using SHE, the content of the query is hidden from the server, but the server learns the nearest cluster. Hiding nearest cluster is a significant performance bottleneck in state-of-the-art schemes like Tiptoe. Specifically, to hide the nearest cluster for a single query, Tiptoe requires the server to perform SHE operations against the entire database. Otherwise, skipping clusters will reveal to the server that the client is not interested in them. Computing over the entire database hides the nearest cluster in the strongest possible sense, but this results in a prohibitively high server computation.
Instead, we propose an approach based on differential privacy. The key advantage is that, in expectation, the server computes encrypted similarity scores against a few clusters per query. Our approach results in a significant improvement in server computation. An additional advantage of our approach is that the request size is only $O(d)$, around $\sqrt{N}$ smaller than Tiptoe. Conversely, differential privacy is a weaker privacy notion than the full, cryptographic, obliviousness Tiptoe achieves.
The starting point of our solution is that in large-scale systems, sufficiently many clients are making the queries at any given time. Therefore, we propose a novel mechanism allowing clients to hide their queries among the crowd of all queries. Our proposed approach is based on $(\epsilon,\delta)$-differential privacy. The approach guarantees to every client that the maximum privacy loss due to their queries is essentially bounded by $\epsilon$ with all but $\delta$ probability. In other words, a client’s queries do not significantly influence the outcome of any statistical inference run by the server.
Intuitively, one would expect that routing all the traffic through an anonymization network would provide sufficient privacy as queries do not carry identifiable information. However, note that only using an anonymization network is not sufficient because by observing traffic patterns, the server can potentially learn information about a client. For example, if the server has side information that a particular client is interested in a specific cluster and all the other clients are known to not access the cluster, then by observing that the cluster is accessed, the server can infer whether a request is from the client or not.
Our protocol randomizes queries to guarantee differential privacy. Specifically, we make the following two changes:
  • Fake queries. In addition to real queries, each client makes a few fake queries. Specifically, for each cluster, the client makes $Z\leftarrow\mathcal{D}$ fake queries, where distribution $\mathcal{D}$ is chosen to guarantee differential privacy after fake queries from many clients are aggregated. In expectation, the client makes $K\mathbb{E}[\mathcal{D}]$ fake queries. For each fake query, the client creates an SHE encryption of zero and pseudorandomly picks on of the clusters.
  • Random query schedule. The client makes queries at random instances within the epoch. For this, we assume that each epoch is divided into a slots, each one second long. For each query (real or fake), the client independently picks a random slot and only makes the query at that slot.
In Section 5, we show that using the negative binomial distribution for fake queries $\mathcal{D}$ along with random query schedule guarantees differential privacy.
3.4 Hiding Metadata Index
After learning the index of the nearest entry, the client might want to retrieve metadata associated with it. When each entry has a small amount of metadata, the server returns metadata of the entire cluster with each query. Note that this leaks no extra information because the server already learns the cluster ID for each query. The client then locally picks the metadata associated with the most relevant entry. However, if the metadata is quite large, the client can run PIR to fetch the metadata of only the most relevant entry.
3.5 Keyword Search
Wally also supports keyword search in which the client is interested in retrieving data associated with a keyword. Using a hash function, the server database is divided into random $K^{\prime}$ buckets (instead of clusters). For each query, the client locally finds the bucket in which the keyword falls and then uses the keyword PIR scheme to fetch a specific record within a bucket.
Wally uses a keyword PIR construction based on cuckoo hashing and index PIR. At a high level, at initialization, the server uses the cuckoo hashing to map the $N$ database elements into a table of size $O(N)$ using two random hash functions. The guarantee is that if a keyword is present in the database, it must be at index $h_{1}(\textit{keyword})$ or $h_{2}(\textit{keyword})$ in the table. The client then uses index PIR to privately fetch entries at these indices and locally find the index containing the entry and associated data. Wally uses MulPIR [ALP+21] as the underlying index PIR, an efficient single-server PIR scheme based on SHE. We emphasize that even though we initialize Wally with MulPIR, it is compatible with any recent efficient PIR schemes
Each bucket consisting of $N/K^{\prime}$ entries represented as $\sqrt{N/K^{\prime}}\times\sqrt{N/K^{\prime}}$ matrix $\mathbf{D}$. MulPIR uses a polynomial version of SHE, where each plaintext is a polynomial. For this explanation, we assume that each entry has the same size as SHE plaintext. If entries are larger than a plaintext, then we split each entry across multiple plaintexts, and if entries are smaller, then we pack multiple entries into each plaintext polynomial when applicable. A client query in MulPIR is a single ciphertext $\hat{q}$ encrypting row and column indices. Given a query, the server does the following.
  1. 1.
    First expand $\hat{q}$ to two $\sqrt{N/K^{\prime}}$-length ciphertext vectors. Encrypting row indicator vector $\mathbf{r}$ and column indicator vector $\mathbf{c}$. To expand, the server uses the oblivious expansion algorithm given in [ALP+21]. This algorithm requires $2\sqrt{N/K^{\prime}}$ homomorphic substitutions.
  2. 2.
    Then compute the matrix-vector product $\mathbf{a}:=\mathbf{D}\mathbf{c}$ with $N/K^{\prime}$ plaintext-ciphertext multiplications and $N/K^{\prime}-\sqrt{N/K^{\prime}}$ ciphertext additions.
  3. 3.
    Finish by computing $\mathbf{r}^{T}\mathbf{a}$ that results a single ciphertext that is a response. This step uses $\sqrt{N/K^{\prime}}$ ciphertext-ciphertext multiplications and $\sqrt{N/K^{\prime}}-1$ ciphertext additions. to compute.
The differential privacy part remains the same as in semantic search. That is, clients make fake queries for random buckets at a random schedule. Section 6 discusses various optimizations we implemented for MulPIR. Overall, our optimizations result in around a three times reduction in server computation and a two-three times reduction in request size.
[…]
Figure 4: Simplified protocol for Wally. Each client generates queries with Algorithm 2 and sends them to the anonymization network, $\mathcal{AN}$, according to the randomized schedule described in Algorithm 4. Cluster $j$ receives anonymous queries denoted as $(\tilde{q}^{j}_{i})$ and computes each response $(r^{j}_{i})$ via SHE.
Algorithm 1 $\mathsf{ServerInit}$
Input: Semantic search database $\{(\mathbf{e}_{i},m_{i})\}_{l\in[N]}$, Number of clusters $K$
Output: Processed semantic $D_{\text{sem}}$,
Clusters centroids $c$.
1:$(\mathbf{c},\mathbf{C})\leftarrow\textsf{K-MeansClustering}(\{(\mathbf{e}_{i},m_{i})\}_{l\in[N]},K)$triangleright dividing semantic data into $K$ clusters
2:for $i\in[K]$ do
3:     for $j\in[d]$ do $\triangleright$ Diagonal packing
4:         for $l\in[|C[i]|/d]$ do
5:              $\mathbf{p}[{j}]=\mathbf{p}[{j}]||[C[i,0]_{l*C[i]/d+j}^{j},$
6:
7:              $\cdots,C[i,0]_{l*C[i]/d+j+d-1}^{d-1}]$
8:         end for
9:         $\mathbf{p}[j]=\mathsf{EncodePtVec}(\mathbf{p}[j])$
10:     end for
11:     for $l\in[h]$ do $\triangleright$ Rotations for BSGS
12:         for $j\in[g]$ do
13:              $\mathbf{p}[gl+j]=\textsf{PtxtRotate}(\mathbf{p}[gl+j],-gk)$
14:         end for
15:     end for
16:     $D_{\text{sem}}[i]=p$
17:end for
18:Output $D_{\text{sem}},c$
4 Wally Protocol
In this section, we describe the complete Wally protocol for a single epoch with $U$ clients.
Server initialization.
The server uses Algorithm 1 to encode the database with $N$ embeddings into $K*d$ vectorized plaintexts. Specifically, the algorithm first divides the input embeddings into $K$ clusters using K-means clustering. We pick $K$ such that the maximum cluster size is less than SHE polynomial dimension $n$ with high probability, which ensures that for each cluster all the embeddings can be encoded into $d$ plaintext vectors. The algorithm then iterates over each cluster separately. Recall each cluster consists of $|C[i]|/d$ column-wise $d*d$ matrices. The algorithm then fills $d$ plaintext vectors for each cluster. In a plaintext vector $j$, the algorithm packs $j$’th diagonal of all $|C[i]|/d$ matrices. Once $d$ plaintext vectors are generated, the algorithm iterates over a group of $g=\sqrt{d}$ plaintexts and rotating each group in increasing multiples of $-g$. Note that these rotations are performed on plaintexts so their cost are negligible.
Query generation.
The client uses Algorithm 2 to generate queries and their schedule at the start of an epoch. The client locally stores the set of cluster centroids $c$ that the server generates during initialization. The algorithm first generates independent queries for at most $\Delta$ real query embeddings. To generate a $j$’th real query, the algorithm picks a centroid $\mathsf{id}_{l}$ nearest to $j$’th query embedding. Then, the algorithm copies the embedding $n/d$ times into a plaintext vector $p$ of size $n$. This is done to take advantage of SHE packing. The algorithm then generates a fresh SHE secret key $\mathsf{sk}$ and evaluation key $\mathsf{evk}$, and encrypts $p$ using $\mathsf{sk}$.
Next, the algorithm generates fake queries. Here, the algorithm iterates over each cluster, sampling a number of fake queries from the negative binomial distribution $\mathcal{NB}(r/U,p)$. To generate a single fake query, the algorithm picks a random centroid, generates a fresh SHE secret key $\mathsf{sk}$ and evaluation keys $\mathsf{evk}$, and encrypts $\mathbf{0}$ using $\mathsf{sk}$.
The algorithm, then permutes the real and fake queries list $\mathbf{Q}$ and generates a random schedule for the queries. To generate the schedule $\mathbf{S}$, for each query $i\in Q$, the algorithm picks a random slot $t$ independently and appends $i$ to $S[t]$. In other words, the schedule $\mathbf{S}$ maps slot indices to the query indices. At a particular time slot $t$ within an epoch, the client will make queries (independently) indicated by $S[t]$.
Server computation.
The server uses Algorithm 4 to generate a response for every query received. The algorithm first aligns the query ciphertext $\mathsf{ct}$. That is, first it copies the query ciphertext $g$ times and then rotates each copy left in increments of one, in total $g$ rotations. Then, the algorithm performs a dot-product between rotated query ciphertexts and the plaintext vectors for cluster $\mathsf{id}$. Recall that we ensure each cluster is encoded into $d$ plaintexts. The algorithm iterates in groups of $g$ plaintexts; within each group, multiply $i$-th plaintext with the $i$-th copy of the rotated query and sum the resulting ciphertexts. These iterations yield $h=\sqrt{d}$ ciphertexts. After that, the algorithm iterates over the $h$ resulting ciphertexts, rotating right each in increments of $g$ and summing all of them into a single ciphertext $R$. This step requires a total of $h$ rotations. The algorithm returns $R$ and cluster $\mathsf{id}$’s metadata.
Algorithm 2 $\mathsf{ClientQuery}$
Input: client queries $q=\{q_{i}\}_{i\in\Delta}$
Clusters centroids $c=\{\mathbf{c}_{l}\}_{l\in[K]}$
$U$ is the number of clients in the system
Output: Real and fake queries list $\mathbf{Q}$.
1:for $l\in[\Delta]$ do $\triangleright$ generate real queries
2:     $\mathbf{p}[1:N/K]=\mathbf{e}_{l}||\mathbf{e}_{l}||\cdots||\mathbf{e}_{l}$
3:     $(\mathsf{sk}_{l},\mathsf{evk}_{l})\leftarrow\mathsf{KeyGen}(1^{\lambda})$
4:     $\mathsf{ct}_{l}\leftarrow\mathsf{EncSemPt}({\mathsf{sk}_{l}},\mathbf{p})$
5:     Pick cluster $\mathsf{id}_{l}$ nearest to $\mathbf{e}_{l}$ from $c$
6:     $Q\leftarrow Q||(\mathsf{ct}_{l},\mathsf{evk}_{l},\mathsf{id}_{l})$
7:end for
8:for $i\in[K]$ do
9:     $F_{i}\leftarrow\mathcal{NB}(r/U,p)$triangleright number of fake queries
10:     for $j\in[F_{i}]$ do $\triangleright$ generate fake queries
11:         $(\mathsf{sk}^{\prime}_{j},\mathsf{evk}^{\prime}_{j})\leftarrow\mathsf{KeyGen}(1^{\lambda})$
12:         $\mathsf{ct}_{j}^{\prime}\leftarrow\mathsf{EncSemPt}(\mathsf{sk}_{j}^{\prime},\mathbf{0})$
13:         Pick random cluster $\mathsf{id}_{j}$ from $c$
14:         $Q\leftarrow Q||(\mathsf{ct}_{j}^{\prime},\mathsf{evk}_{j}^{\prime},\mathsf{id}_{j})$
15:     end for
16:end for
17:Randomly permute $Q$
18:$\mathbf{S}\leftarrow\mathsf{RandScheduleGen}(Q)$triangleright generate schedule
Algorithm 3 $\mathsf{RandScheduleGen}$
Input: List $\mathbf{Q}$ of real and fake queries
1:Epoch length is $T$
2:Initialize a schedule $\mathbf{S}$triangleright $T$ empty lists
3:for $j\in[|\mathbf{Q}|]$ do
4:     $i\leftarrow[T]$triangleright Sample random slot
5:     $\mathbf{S}[i]\leftarrow\mathbf{S}[i]||j$
6:end for
7:return $\mathbf{S}$
Algorithm 4 $\mathsf{ServerComputation}$
Input: Query $(\mathsf{ct},\mathsf{evk},\mathsf{id})$
Output: Encrypted scores and data $r$.
1:for $j\in[g]$ do $\triangleright$ Align query
2:     $\mathsf{ct}^{\prime}[j]=\mathsf{SHERotate}_{\mathsf{evk}}(\mathsf{ct},j)$
3:end for
4:$D=D_{\text{sem}}[\mathsf{id}]$triangleright Dot-Product
5:$R=\sum_{k\in[h]}\mathsf{SHERotate}_{\mathsf{evk}}(\sum_{j\in[g]}D[j+kg]*\mathsf{ct}^{\prime}[j],gk)$
6:$\textsf{data}\leftarrow\textsf{data}_{\mathsf{id}}$triangleright Cluster $\mathsf{id}$’s metadata
7:return $r:=(R,\textsf{data})$
5 Security
In this section, we prove that if honest clients follow Algorithm 2 to generate queries, then the overall system achieves $(2\epsilon,2\Delta\delta)$-differential privacy in each epoch. Recall from Algorithm 2 that each query consists of the nearest centroid and the embedding. We only focus on proving the privacy of centroids because the IND-CPA security of SHE protects query embeddings. Therefore, we model each query as a nearest centroid without an embedding in this section. During each epoch, the server receives multiple queries for each cluster. Towards the end of epoch, the server’s view can be considered as a noisy histogram over all clusters. We consider this noisy histogram as an output of the DP mechanism.
We prove this argument in two steps: First, we show that a noisy histogram output by a curator, which gets users’ queries as input and uses negative binomial $\mathcal{NB}$ mechanism to sample fake queries, is $(2\epsilon,2\Delta\delta)$-DP in the central model. Second, we show that the view of the server in an epoch can be simulated only using the curator’s output.
5.1 DP Security in the Central Model
We define the central model as follows:
Central model.
The curator performs the following:
  1. 1.
    Collects the real queries from all the clients, $\{R_{u}\}_{u\in[U]}$, where $U$ is the total number of clients, and $R_{u}$ is each client’s set of queries and will be at most $\Delta$. Let $R$ denote the list of all real queries.
  2. 2.
    For each cluster $b$, samples the number of fake queries $F_{j}\leftarrow\mathcal{NB}(r,p)$. Add it to the fake queries list $F$.
  3. 3.
    Given the list of real and fake queries, randomly permute them: $\mathsf{list}\leftarrow\mathsf{permute}(R,F)$.
  4. 4.
    Send $\mathsf{list}$ to the server.
Another way to view it is that the curator sends a noisy histogram of queries to each cluster to the server. Recall that each client can contribute at most $\Delta$ queries.
We then prove the following theorem:
Theorem 1.
For any query database of size $N$, number of clusters $K$, differential privacy parameters $\epsilon,\delta\in(0,1)$, and $\Delta\in\mathbb{N}$, let $p=e^{-0.2\epsilon/\Delta}$ and $r=3(1+\log(1/\delta))$. Then, the negative binomial mechanism $\mathcal{NB}(r,p)$ is $(2\epsilon,2\Delta\delta)$-DP for $\Delta$-histogram.
Proof.
We define the curator as Algorithm $M$ that takes input from a database $X$ of requests from $N$ clients and outputs a noisy histogram. We prove the theorem by investigating the affected buckets when we change the input $X$ by replacing a client with another.
For a particular cluster $b$ in the histogram, we define $M_{b}(\cdot):=R_{b}+\mathcal{NB}(r,p)$, where $R_{b}$ is the total real queries for the cluster $b$ in the database. Note that this is the curator’s exact output for each cluster. We break down the proof into two edge cases: 1) the simpler case where $X$ and $X^{\prime}$ differ by a client which sends all their messages to a single cluster and 2) they differ by a client which sends $\Delta$ messages to $\Delta$ different clusters.
Simple case. Consider two neighboring databases $X$ and $X^{\prime}$ which differ in queries contributed by a single client. Concretely, in $X$ the client contributes $\Delta$ queries to a cluster $b$ and in $X^{\prime}$ instead contributes to another cluster $b^{\prime}\not=b$. Note that when $j\notin\{b,b^{\prime}\}$ then $M_{j}(R)$ and $M_{j}(R^{\prime})$ are identically distributed. Thus, the privacy provided by $M_{j}$ is 0-DP for all $j\notin\{b,b^{\prime}\}$.
Now consider the cluster $b$; the two databases differ by $\Delta$ real queries for this cluster. By the following, lemma $M_{b}$ provides $(\epsilon,\delta)$ privacy on cluster $b$.
Lemma 2 (Theorem 13 [GKMP20]).
[…]
We can make a similar argument for a cluster $b^{\prime}$. Therefore, by basic composition, Lemma 1, the total privacy provided by $M$ for these neighboring databases is $(2\epsilon,2\delta)$—DP.
General case. Now, we consider a general case where two neighboring databases that differ in queries from a single client. In database $X$ the client contributes queries $\gamma_{j}$ for cluster $j$ and in database $X^{\prime}$ the client contributes queries $\gamma_{j}^{\prime}$ for the cluster $j$. Define $\Delta_{j}:=|\gamma_{j}-\gamma_{j}^{\prime}|$. The total difference between two neighboring databases is given by $\sum_{j}^{B}\Delta_{j}\leq 2\Delta$.
Let $\epsilon_{j}:=\frac{\epsilon\Delta_{j}}{\Delta}\leq\epsilon$. Further, we only consider the clusters where $\Delta_{j}>0$ since the clusters where $\Delta_{j}=0$ are already 00-DP. (Equivalently, $\epsilon_{j}=0$ for these $j$.) Then, Lemma 2 implies cluster $j$ is $(\epsilon_{j},\delta)$-DP for $\Delta_{j}$-summation for the same mechanism since
$\frac{\epsilon}{\Delta}=\frac{\epsilon}{\Delta}\cdot\frac{\Delta_{j}}{\Delta_{j}}=\frac{\epsilon_{j}}{\Delta_{j}}.$
Simple composition over all the clusters yields
$\sum_{j}\epsilon_{j}=\frac{\epsilon}{\Delta}\sum_{j}\Delta_{j}\leq 2\epsilon.$
Basic composition also yields that the composed protocol has $\delta^{\prime}:=2\Delta\delta$ since at most $2\Delta$ clusters differ.
We emphasize that the proof is over a particular epoch $S$. If clients participate in many epochs, say $l$, then basic composition yields $(2l\epsilon,2l\cdot\Delta\delta)$-DP.
5.2 Generating view of the server
We show how the curator’s output is sufficient to simulate the server’s view generated by distributed DP algorithm defined in Algorithm 2 together with an anonymization network (AN). We also require pseudorandom ciphertexts and evaluation keys, a common assumption for RLWE-based SHE schemes. The latter is also why we cannot reuse evaluation keys.
Theorem 2.
Let $\mathcal{AN}$ be an anonymous network that 1) removes all identifying information from messages it receives and 2) randomly permutes all the messages it receives over a second-long slot. Let each honest client run the distributed mechanism in Algorithm 2 with messages sent to the server via $\mathcal{AN}$. Then, the view of the server in an epoch can be simulated using the curator noisy histogram output described above in the central model.
Proof.
We use infinite divisibility of $\mathcal{NB}$ distribution to produce a standard hybrid argument in cryptography.
Hybrid 0: Every client is the same as the real client given in Algorithm 2 except it does not include encrypted embeddings with each request (real or fake). Hybrid 0 is equivalent to the real client finding the nearest centroid for each real query and a fake random centroid for each fake query. The client then picks a random slot for each query and send associated queries at each slot.
Hybrid 1: In this hybrid, we replace all the clients with a single simulator. The input of a simulator is a list $R=\{R_{u}\}_{u}$, where $R_{u}$ is real queries for $u$-th client. The simulator then simulates each client $u$ with input $R_{u}$ as in Hybrid 0.
As the simulator internally runs each client $u$ as in Hybrid 0 with input $R_{u}$. From the adversary’s point of view, the distribution of requests is the same as Hybrid 0.
Hybrid 2: In this Hybrid, the simulator’s input is the same as in the previous Hybrid 1. The simulator is defined as follows.
  1. 1.
    Collects the real queries from all the clients, $\{R_{u}\}_{u\in[U]}$. Call it a list of real queries $R$.
  2. 2.
    For each cluster $b$, samples the number of fake queries $F_{j}\leftarrow\mathsf{NB}(p,r)$. Add it to the fake queries list $F$.
  3. 3.
    Given the list of real queries $R$ and fake queries $F$, randomly permutes them: $\mathsf{list}\leftarrow\mathsf{permute}(R,F)$
  4. 4.
    Initialize schedule list $S$ of size $z$ (where z is the total number of slots in an epoch). For each query $q\in\mathsf{list}$, pick a random slot $p\leftarrow z$ and set $S[p]=S[p]\cup q$.
  5. 5.
    For each slot $p\in S$ send queries in $S[p]$ to the server.
There are two differences between Hybrid 2 and Hybrid 1:
  1. 1.
    In Hybrid 1, each simulated client samples fake queries for each cluster from a distribution $\mathcal{NB}(r/U,p)$. While in Hybrid 2 (Line 2) the simulator samples fake queries per cluster from $\mathcal{NB}(r,p)$.
  2. 2.
    In Hybrid 1, each simulated client generates a schedule for its queries independently. While in Hybrid 2, the simulator generates a schedule for queries.
To show that Hybrid 2 and 1 are indistinguishable, we use the definition of infinitely divisible distribution, Definition 1.
Note that negative binomial distribution is infinite divisible because $\mathcal{NB}(r,p)\ \sum_{i=1}^{U}\mathcal{NB}(r/U,p)$. Thus, the distribution of per cluster fake queries per cluster in Hybrid 2 (Line 2) is the same as in Hybrid 1. Observe that a random slot is picked for each query independently in both Hybrids. Therefore, the query schedule generated in both Hybrids is indistinguishable.
To complete the proof, note that in Line 3, the simulator generates $\mathsf{list}$ in the same way as the curator in the central model. After that, the simulator uses this $\mathsf{list}$ to generate a random query schedule, which results in a server’s view indistinguishable from the view in Hybrid 0.
6 Optimizations
Here we present our optimizations for semantic and keyword search. Our optimizations revolve around the [Bra12, FV12] and BGV [BGV14] SHE schemes. Further, we describe several optimizations to the MulPIR scheme [ALP+21]. First, we describe the standard RNS optimization in SHE [GHS12] and a method to compress ciphertexts by dropping LSBs as described in Cheetah [HLHD22].
RNS form.
Often $Q$ is larger than machine-size words and we use $Q$ as the product of machine-sized primes $Q=q_{0}q_{1}\cdots q_{l}$. This allows us to store a ciphertext as two polynomials each stored as an $n\times(l+1)$ matrix of integers by the Chinese remainder theorem (CRT) on $Q$. Further, we require each $q_{i}$ to be NTT friendly, $q_{i}=1\mod 2n$. This allows us to use the number theoretic transform (NTT) to efficiently switch between each evaluation form and coefficient form in time $O(n\log_{2}n)$, since the NTT is a modular version of the fast Fourier transform over a prime modulus. We can efficiently compress a ciphertext by modulus-switching down to a smaller modulus, $Q^{\prime}$ where $Q^{\prime}|Q$, e.g., $Q^{\prime}=q_{0}$, without affecting the ciphertext’s noise budget.
Dropping ciphertext LSBs.
We use Cheetah’s [HLHD22] method to compress ciphertexts after homomorphic computation: we modulus switch down to the smallest RNS modulus, $Q^{\prime}=q_{0}$, then drop the least significant bits of the ciphertext. This technique compresses the response size further than simply modulus-switching down, but it also adds noise. Cheetah incorrectly models the ciphertext as noiseless (Appendix F in [HHC+23]), but we derive the proper analysis as follows. Dropping the LSBs of a BFV ciphertext $\mathsf{ct}=(c_{0},c_{1})$ can be analyzed by expressing the polynomials as high and low bit polynomials $c_{0}=2^{b_{0}}c_{0}^{h}+c_{0}^{l},\ c_{1}=2^{b_{1}}c_{1}^{h}+c_{1}^{l}$. Then, the new decryption equation is
$2^{b_{0}}c_{0}^{h}+2^{b_{1}}c_{1}^{h}s=\lfloor Q/t\rceil m+e+e_{drop}$
for
$e_{drop}=-c_{0}^{l}-c_{1}^{l}s.$
We have correctness as long as $\|e+e_{drop}\|_{\infty}<Q^{\prime}/2t$. Concretely, we are able to drop around $8$kB from the response size in approximate nearest neighbors search to a response size of $47$kB when combined our BFV plaintext CRT optimization described in the next subsection.
6.1 Approximate Nearest Neighbors Optimizations
The homomorphic computation in the approximate nearest neighbors use case is $C$ inner products over real embedding vectors satisfying $\mathbf{e}$, normalized to $\|\mathbf{e}\|_{2}=1$ in our use case, where $C$ is the maximum cluster size. We scale $\mathbf{e}$’s entries to integers, round them to a vector $\mathbf{\widetilde{e}}$, and perform the inner product over the integers. The computation is correct as long as the inner product does not wrap around mod $t$. We define plaintext precision as $\approx\log_{2}(t)/2$ since the inner product has a multiplicative depth of one.
Minimal rotation keys and bandwidth.
Wally uses the “baby-step giant-step” (BSGS) optimization [HS18] to minimize the number of rotation keys: for $d$-dimensional inner products, BSGS requires $2\sqrt{d}$ rotations for step sizes of $1$ (baby) and $\sqrt{d}$ (giant). See Algorithm 1 in [HS18] for the BSGS algorithm. Additionally, we are able to use two rotation keys since $d<n/2$ in our use case. This optimization leads us to a query size of one ciphertext $\mathsf{ct}$ and two rotation keys $\mathsf{evk}=(\mathsf{evk}_{1},\mathsf{evk}_{\sqrt{d}})$. For key-switching, and relinearization in PIR below, we use the hybrid-GHS [GHS12] key-switching strategy described in-detail by Kim et al. [KPZ21].
Say we require a plaintext modulus of $t\approx 2^{29}$ for $\sim 14$ bits of plaintext precision. The smallest BFV ring dimension supporting this is $n=8192$ which requires a evaluation key modulus of $Q\approx 2^{165}$, ciphertext modulus of $Q^{\prime}\approx 2^{110}$ for 128 bits of post-quantum security [APS15] [^4] https://github.com/malb/lattice-estimator. A query consists of $2$ polynomials for the ciphertext and $2$ polynomials for the evaluation keys, or $225$kB and $676$kB respectively, for a total query size of $901$kB for these parameters.
Re-using BFV evaluation keys via the plaintext CRT.
Wally achieves minimal bandwidth per client query by a novel use of the Chinese remainder theorem (CRT) on the BFV plaintext space. The CRT on plaintexts allows us to keep a set amount of plaintext bits while re-using one set of evaluation keys across all plaintext spaces since BFV’s evaluation keys are independent of the ciphertext modulus, unlike BGV evaluation keys. Therefore, we encrypt a vector $\mathbf{\widetilde{v}}$ as $\mathbf{\widetilde{v}}\mod t_{i}$ for NTT-friendly primes $\{t_{i}\}$. The computation is correct if the inner products do not wrap around modulo $t=t_{0}t_{1}\cdots t_{l}$, but the individual ciphertexts can have their computation wrap around modulo each $t_{i}$. Therefore, we can drop the ring dimension $n^{\prime}=4096$ and use two ciphertexts, modulo $t_{0}=40961\approx 2^{15}$ and modulo $t_{1}=65537\approx 2^{16}$ for $t_{0}t_{1}\approx 2^{31}$, half the bandwidth from ciphertexts, and use evaluation keys in dimension $n^{\prime}$, or one fourth the bandwidth from evaluation keys. Concretely, this leaves us with $113$kB for ciphertexts and $170$kB for evaluation keys, $283$kB in total, over a 3×3$\times$3 × improvement in query size without sacrificing plaintext precision.
In general, the plaintext CRT is preferable to increasing the ring dimension since the evaluation key sizes dominate the ciphertext size: six polynomials versus 2 polynomials. This is true in the bandwidth-optimal key-switching strategy for hybrid GHS key-switching. Another advantage of the plaintext CRT is that it allows for very high precision computations which do not wrap around mod $t$. One limitation of this optimization is that we run out of NTT-friendly plaintext primes for small parameter sets, like $n=4096$ or $2048$.
6.2 PIR Optimizations
Uneven dimensions.
In general, the cost of a ciphertext-ciphertext multiplication is much higher than a ciphertext rotation and both are much higher than addition or plaintext multiplication. For example, SEAL [CLP17] on an AWS t2.2xlarge AWS instance has ciphertext addition and plaintext multiplication as less than $100\mu s$ while ciphertext rotation is $3.6ms$ and ciphertext multiplication is $12.1ms$ [MR23].
Therefore, we structure the database (cluster) as a rectangle $\mathbf{D}\in R_{t}^{d_{1}\times d_{2}}$ with uneven dimensions $d_{1}$ and $d_{2}$. For any fixed HE parameter set, the server’s compute time is dominated by ciphertext multiplications and rotations and that the former is a constant factor $\gamma$ more than the latter. This simplification yields a compute time of $2d_{1}t_{r}+2d_{2}t_{r}+d_{2}t_{\times}$ where $t_{\times}$ and $t_{r}$ are the times to multiply and rotate ciphertexts, respectively. More simply, the total compute time, measured in terms of rotations, is
$(\gamma+2)d_{1}+2d_{2}=(\gamma+2)C/d_{2}+2d_{2}$
rotations. Minimizing this as a function in $d_{2}$ gives $d_{2}=\sqrt{1+\gamma/2}\sqrt{C}$.
This optimization saw an improvement of $30-40\%$ in MulPIR run times. We chose this concrete analysis since SHE parameter regimes are restricted per use case and it is much simpler than an asymptotic analysis.
Linearizing parts of query expansion
MulPIR is an optimized query expansion which originated with Angel et al. [ACLS18]. In short, a query index is encoded in a plaintext polynomial then expanded into $k$ ciphertexts by calling $k$ ciphertext rotation operations (Galois automorphisms followed by key-switching). We noticed that some of these expansions can be substituted with linear operations: call $\mathsf{ct}\pm\mathsf{ct}^{\prime}$ where $\mathsf{ct}^{\prime}$ is $\mathsf{ct}$ rotated. In general, we saw a $10-25\%$ improvement in MulPIR’s expansion from our optimization.
Lazy rescaling in BFV multiplication.
We apply the lazy rescaling technique of Kim et al [KPZ21] in MulPIR’s last step, the inner product between ciphertexts $\mathbf{r}^{T}\mathbf{a}$. The main idea here is that BFV multiplications first multiply polynomials over the integers, then scale and round back to integers modulo $Q$. Kim et al [KPZ21, Appendix F] noticed that one can multiply over the integers, add over the integers, then scale and round after the additions, from the dimension number of scaling operations to one. Lazy rescaling yielded a $15-20\%$ improvement in our MulPIR implementation.
Handling Very Large Entries
In MulPIR the database is represented as a plaintext matrix $\mathbf{D}\in R_{t}^{d_{1}\times d_{2}}$. Conventionally, the response is computed as
[…]
This approach results in $n\cdot t$ ciphertext-plaintext multiplications and $d_{2}\cdot t$ ciphertext-ciphertext multiplications, where $t$ is the number of plaintexts used to encode one entry. The number of ciphertext-ciphertext multiplication grows linearly with $t$ and remains a concretely small value when entries are not large. But it will blow up with extremely large entries (e.g. when they are high-resolution photos).
To overcome this challenge, we can swap the order of the computation as
[…]
The outer-product takes $d_{1}\cdot d_{2}$ ciphertext-ciphertext mutliplications to compute, which is independent of the entry size and $t$. This optimization reduce the computation overhead when $t$ is larger than the first dimension $d_{1}$.
Keyword PIR.
MulPIR’s keyword PIR functionality is given by using Cuckoo hashing to find the entry’s proper index. We noticed for our cluster sizes $C$, we are able to split the cluster into two hash tables and perform $1$-hash Cuckoo hashing on each without increasing the (empirical) failure probability. This resulted in a 2×2$\times$2 × improvement in query size and server computation time.
Low-level SHE Optimizations
BFV’s multiplication step over the integers is done by extending the modulus $Q$ to a larger modulus $PQ$ so the multiplication does not wrap around. We use Bajard et al.’s method to extend the basis, “fast basis conversion” in [BEHZ16] and re-use the basis elements in $Q$ for a $\sim 15\%$ improvement in BGV multiplication.
We add a conditional, lazy modular reduce to Harvey’s NTT [Har14]. The main idea is that we only reduce during the butterfly if the integers would exceeds the machine word size. This saw a $20-25\%$ improvement in forward NTT times and $15-20\%$ improvement in inverse NTT times for $61$-bit primes. Note, we saw no improvement on primes near $64$-bits.
7 Performance
In this section, we demonstrate Wally’s scalability and concrete performance as seen in experiments. We also track Wally’s performance tradeoffs: fake query overhead versus number of clients, fake query overhead versus the DP parameters, and the difference in low-precision performance versus high-precision performance.
Experimental setup.
We ran server-side computations on an AWS c5.24xlarge instance with 192GB of RAM, single-threaded. We extrapolate throughput from the single-threaded times seen on this AWS instance.
End-to-end performance.
Metric Low prec. Mid prec. Tiptoe
Exp. query (MB) 0.67 0.84 16.2
Exp. response (MB) 0.16 0.18 3.2
Queries/s 2998 1499 47
8M queries (hrs) 0.4 1.5 47.5
Fake query 3×3$\times$3 × 3×3$\times$3 × --
overhead
Precision (bits) 6 13 4
Privacy DP DP SHE
Network req. AN AN --
Table 2: Wally performance for a parameter set with BFV dimension $n=4096$ on a 35M entry dataset with the maximum cluster size as $4096$ compared to Tiptoe. Query and response times and sizes for Wally include the number fake query overhead (3×3$\times$3 × or two fake for each real query). Throughput for Wally is measured for 8M real queries, or 24M in total. All Wally responses are computed as the encrypted scores plus $44$kB of metadata.
Table 2 describes our main experiments in comparison with Tiptoe. Here, we use two parameter sets for Wally: first, we use a small BFV parameter set ($n=4096$, $\log_{2}t=15$, and $\log_{2}q=83$) and second, a plaintext CRT (Section 6) parameter set which includes the former. The plaintext precision is $(\log_{2}t)/2-(\log_{2}d)/4$ since the inner-product computation grows heuristically as $\sqrt{d}\|\mathbf{\tilde{e}}\|_{\infty}^{2}$ where $\mathbf{\tilde{e}}=\alpha\cdot\mathbf{e}$ is a scaled (unit) embedding vector, $\|\mathbf{e}\|_{2}=1$. The former parameter set enables six bits of plaintext precision whereas the latter enables 13 bits.
We set $\Delta=2$ messages per client per epoch and calculated the DP overhead by using $(\epsilon,\delta)=(1,2^{-20})$ after the loss in Theorem 2 with four million clients per epoch. The fake query overhead is derived by the expected number of fake queries, for the negative binomial mechanism, $\mathcal{NB}(r,p)$, as
$\mathbb{E}[\text{total fake queries per client}]=\frac{rp}{1-p}\cdot\frac{K}{U}$ (1)
where $K$ is the total number of clusters and $U$ is the total number of clients.
Wally’s ability to cluster the dataset and keep server computation time proportional to the maximum size of a cluster is demonstrated in its small per-query server computation time, $95$ms for each real or fake query, and its relatively small expected query size, $679$kB. We keep each cluster size less than the total number of plaintext slots in each BFV ciphertext, $n=4096$. Otherwise, Wally can compute similarity scores in parallel if the cluster size is larger than $n$. Large meta-data sizes can be retrieved by adding another round of PIR after the similarity computation. The throughput times represent how many queries per second can be processed on $K=8545$ clusters during an epoch needed to serve $24$M total queries, real and fake. This throughput calculation leads to a minimum epoch size of $44$ minutes serving roughly 3k queries per second. This compares to the $47.5$ hours Tiptoe would take to serve $8$M queries.
Comparing to Tiptoe.
Wally has the following two extra assumptions compared to Tiptoe:
  • Wally requires sufficiently many honest clients in the system during an epoch.
  • It also needs an anonymization network that removes all identifying information from incoming messages and randomly permutes messages over a short amount of time, e.g., one second time slots.
In other words, Wally trades off scalability for privacy. Such a tradeoff is needed since oblivious, cryptographically secure systems require the server to compute over the entire database and cannot scale to many clients.
Tiptoe ran its encrypted similarity computation on a cluster of $80$ r5.xlarge AWS instances, each with 4 vCPUs and 32 GB of DRAM, and saw a latency of $11.2$ seconds per query, of which is a significant amount of per-query preprocessing (8.7 seconds). Each server held $8$GB to $12$GB of the plaintext database. We emphasize that each Tiptoe query computes of the entire database (all clusters), and we estimate their total single threaded time as 80×80$\times$80 × their per query time, or $183$ seconds per query. We obtained this estimate by dividing the pre-processing (token) times and the download bandwidths in Table 7 in [HDCZ23] by six since our embedding dimension is 3×3$\times$3 × higher and Tiptoe generates two pre-processing elements per query, “tokens”: one each for a similarity computation and PIR. The non-pre-processing times and bandwidths, as well as the token upload bandwidth, are divided by 3×3$\times$3 × again due to the difference in embedding dimension and that they use the same token upload for both the similarity computation and PIR. Next, the server compute times are multiplied by $80$ since they use $80$ clusters, yielding an estimated $163$ seconds, single-threaded, for a homomorphic similarity computation, i.e., a Wally query. Given this time, we divide it by the number of clusters we expect in Wally, $8545$, to estimate throughput in Table 2.
In comparison, Wally can process each query in parallel since queries, real and fake, have their computation restricted to a single cluster. If there are $U$ clients making at most $\Delta$ queries, then Wally can set a epoch time as the minimal amount of time needed to compute $\phi\times U\times\Delta$ queries. Since each client sends a query at a random time over an epoch, the average response time is half the epoch whereas the worst case time is the epoch length. We derived Wally’s throughput numbers in Table 2 using $95$ms per query, $U=4$M, $\Delta=2$, and $\phi=2$. This leads to $24$M queries in 44 minutes, of which $8$M are real queries, on $K=8545$ clusters.
Scaling DP with $\epsilon$.
Figure 5 shows how the expected fake queries per cluster for each client scales with $\epsilon$ for one million clients and four million clients. This is derived by dividing Equation 1 by the number of clusters $K$. Note that the total number of clients has a linear effect while the dependence on $\epsilon$ is nonlinear as we decrease $\epsilon$.
Figure 5: The relationship between the number of queries per cluster per client for $500$k and $1$M total clients. We fix the number of real queries to $\Delta=2$.
Precision.
Table 3 shows how each query grows as we increase precision, i.e., as we increase the BFV parameters using our plaintext CRT optimization described in Section 6. The six bit precision parameters are the same as we used in Table 2 while the $13$ bit parameters use the plaintext CRT on $\log_{2}(t_{1})=15.3$ and $\log_{2}(t_{2})=16$, $\log_{2}(t)=\log_{2}(t_{1}t_{2})=31.3$. Lastly, the 29 precision parameters use two $n=8192$ BFV parameter sets with $\log_{2}t_{1}=29$ and $\log_{2}t_{2}=30$.
Precision (bits) 6 13 29
Query (KB) 226 283 901
Response (KB) 53 63 196
Server comp. (ms) 95 190 580
Table 3: Here we list the query and response size per each query (real and fake). Each response size includes $44$kB of metadata. Multiply all numbers by $\phi+1$ if DP necessitates $\phi$ fake queries per real query.
Scaling with the number of clients.
Table 4 shows how the fake query overhead varies with the number of messages per epoch $\Delta$ and the number of clients in the system. We emphasize that the number of fake queries increases linearly with the number of buckets and decreases proportionally to the number of clients. As mentioned above, this can be offset by increasing the size of the clusters to a multiple of $n$ and we can perform the similarity computation in parallel with more threads. Table 4 emphasizes the importance of having many clients in a system since the overhead vanishes as the number of clients becomes large.
Number of Clients 0.5M 1M 2M 10M
Fake queries ($\Delta=1$) $7.5$ $4.8$ $2.9$ $1.4$
Exp. MB ($\Delta=1$) $1.9$ $1.08$65.65.65.65 $3.11$
Fake queries ($\Delta=4$) $35.3$ $18$ $9.6$ $2.7$
Exp. MB ($\Delta=4$) $8$ $4.1$ $2.2$61.61.61.61
Table 4: This table shows how the fake query overhead changes with the number of clients and the number of messages. We fix the number of clusters to $K=8545$ to match our 35M dataset, Table 2.
7.1 Keyword Query Performance
Table 5 shows single-threaded keyword PIR server performance, measured on an M1 Max laptop. Note, however, that each dataset size requires manual tuning for optimal encryption and cuckoo hashing parameters. The request size and the response size for each entry in the table is 728KB and 39KB, respectively.
Database Entries Entry Size (bytes) Time (s)
100K 10 0.2
100K 100 0.5
1M 10 0.7
1M 100 2
10M 10 2.7
100M 100 6.3
Table 5: Keyword-PIR, server run times for Wally. These times include the optimizations in Section 6.
8 Conclusion
Wally demonstrates that private search systems can be scalable to systems with many users. Wally carefully balances somewhat homomorphic encryption together with differential privacy to hide the client’s query and reveal a differentially private histogram over all clients’ traffic to database clusters. Further, Wally’s overhead from differential privacy vanishes as the number of users in the system increases. This vanishing overhead is crucial to scalability and is potentially useful in many private systems deployable at-scale.