### v4

# Scalable Private Search with Wally

Abstract
This paper presents Wally, a private search system that supports efficient semantic and keyword search queries against large databases. When sufficiently many clients are making queries, Wally’s performance is significantly better than previous systems. In previous private search systems, for each client query, the server must perform at least one expensive cryptographic operation per database entry. As a result, performance degraded proportionally with the number of entries in the database.
In Wally, we remove this limitation. Specifically, for each query the server performs cryptographic operations against only a few database entries. We achieve these results by requiring each client to add a few fake queries, and send each query via an anonymous network to the server at independently chosen random instants. Additionally, each client also uses somewhat homomorphic encryption (SHE) to hide whether a query is real or fake. Wally provides $(\epsilon,\delta)$-differential privacy guarantee, which is an accepted standard for strong privacy.
The number of fake queries each client makes depends inversely on the number of clients making queries. Therefore, the fake queries’ overhead vanishes as the number of clients increases, enabling scalability to millions of queries and large databases. Concretely, Wally can process eight million queries in 117 minutes, or just under two hours. That is around four orders of magnitude faster than the state of the art.
1 Introduction
Consider a scenario where a client holding a search query wants to retrieve relevant information from a large database hosted on a server. This scenario reflects the flow of current search engines such as Google, Bing, and DuckDuckGo. These search engines improve the quality of our lives by promptly providing helpful information. Nevertheless, they require learning the client’s query in order to respond accurately, hence providing no query privacy. In some situations, learning the query could reveal sensitive and personal information about the client. For example, consider the queries “High blood pressure with late-stage HIV?”, “Closest diabetes clinic from Albuquerque”, and “Best divorce lawyers”. Revealing these queries to search engines could be damaging to the client. However, even if the queries do not appear sensitive, they may allow the server to learn personal information like medical information, marital status, sexuality, etc. Therefore, enabling query privacy is critical for search engines.
Private search engines opt to provide query privacy. However, these systems have significantly higher computational and communication overhead than the state-of-the-art insecure solutions. For example, in Tiptoe [26], a state-of-the-art private search engine, a single client query is 21 MB and requires 339 core-seconds of server computing against a database of $400$ million entries. The high computational cost is because, for each client query, the server must scan the entire database; otherwise, it will learn which database entries the client is not interested in. Similarly, the client must send cryptographic material for each database entry, which results in high communication cost [^1] In reality, these systems trade off response size with request size by using a standard database clustering trick. We detail this trade off in a later section.. An additional limitation of Tiptoe is that it requires each client to store a database-dependent state of around 50 MB for each client query.
In this paper, we present Wally, a privacy preserving search engine that efficiently scales to the performance constraints of large-scale search engines. Wally supports semantic search, where the client wants to fetch the most relevant documents to a query, and keyword search, where the client wants to retrieve documents associated with a private keyword. To provide privacy, our key observation is that in search engines, there are always enough clients making queries. We exploit this observation to gain privacy by devising a way to hide a particular query in the crowd of queries. In our protocol, the server computation and the communication are amortized over the number of clients making the requests; therefore, when the number of clients is large, the per-query cost becomes minimal.
Wally achieves these results by relaxing its privacy notion to $(\epsilon,\delta)$-differential privacy. In Wally, the server learns a noised differentially private histogram of anonymized queries from all clients, over a partition of the server database. We emphasize that previous works, e.g., Tiptoe, are fully oblivious. That is, the server learns nothing besides that a particular client is making a query. Therefore, the privacy guarantee of schemes like Tiptoe is stronger than Wally. However, differential privacy is an accepted standard for strong privacy [16]. Specifically, every client in Wally is guaranteed that any statistical inferences are insensitive to their data. In other words, their participation does not significantly alter the server’s view, except with probability $\delta$, and thus provides strong privacy protection.
At a high level, Wally works in epochs. Within each epoch, a fixed number of users make queries. The epoch length is picked to ensure enough clients are present to make queries, and their queries will be processed by the end of an epoch. The server divides the database into $K$ clusters. Each query consists of a relevant cluster index, and the server response is all the entries in the cluster. The queries arrive at the server via an anonymous network, ensuring all identifiable information like IP address is removed from the queries. The server then generates a response to each query independently and sends it back to the respective client via an anonymous network.
Even though queries are anonymized, they do not provide query privacy. The server can infer information about the client query by exploiting the access pattern (if a particular cluster is accessed) or arrival time (when the server receives queries). To prevent this leakage, Wally implements the following two changes at the client.
First, to ensure that the server does not learn information by observing the access patterns, each client in Wally makes fake queries. For each fake query, the client picks a randomly selected bucket. This ensures that the server cannot differentiate between real and fake queries. The number of fake queries is sampled from a particular distribution to ensure the revealed histogram of queries is $(\epsilon,\delta)$-differentially private. In expectation, each client makes $O(\frac{C\Delta\log(1/\delta)}{U\epsilon})$ fake queries where $U$ is the number of clients within an epoch, and $\Delta$ is the maximum number of queries each client can make in an epoch. Note that increasing the number of clients decreases the average number of fake queries. The overhead due to fake queries can be traded with privacy because more fake queries are required as $\epsilon$ decreases. We emphasize that fake queries do not affect the accuracy or correctness seen by any client, as the client can discard the responses due to these queries.
Second, Wally also ensures that arrival time and order of the queries leak no information about the client to the server. To achieve this, instead of making a query instantly, the clients make queries at random one second slots within an epoch. Also, an anonymous network collects the requests received within a particular slot and sends them as a batch to the server.
Anonymizing the queries, adding fake queries that are indistinguishable from real ones, and randomizing query order and time guarantees differential privacy.
As mentioned above, for each query, the server response contains all the entries in the cluster. For large clusters, this would result in very high response overhead. In Wally, we utilize lattice-based, somewhat homomorphic encryption (SHE) to reduce the response overhead. Specifically, for semantic search, each query additionally includes a semantic information (dependent on the client’s input) encrypted under SHE. The server computes the distance function between the encrypted information and the cluster entries under SHE and returns encrypted scores. This reduces the response because the size of encrypted scores is significantly smaller than the entries.
For a keyword search, the client sends an encrypted PIR query, and the server runs an SHE-based keyword private information retrieval (PIR) protocol within the cluster, which returns data associated with the client keyword. To ensure that the server cannot differentiate between real and fake queries, for every fake query the client includes encrypted semantic, or keyword queries based on fake data. We propose various optimizations that significantly improve performance for both kinds of searches.
Overall, for a database with 35 million entries Wally can process eight million real queries in 117 minutes on 320 cores, a drastic improvement over the state-of-the-art, which would have required years to process the same queries. Additionally, the size of each request is 53 times smaller and the size of each response is 3.5 times smaller compared to Tiptoe.
Our contributions are as follows:
  • We describe Wally’s design, which enables privacy-preserving large-scale search engines. Its novelty stems from the combination of anonymity, fake queries added by the clients, and the client’s use of SHE to encrypt its queries.
  • We also describe the implementation of Wally, showing efficiency and scalability along with a robust privacy guarantee. Generally, SHE computations have a high performance overhead. To achieve desired performance constraints, we propose several optimizations to SHE computation (distance computation and PIR) that could be of independent interest.
  • We evaluate Wally for private image search, demonstrating both strong privacy and low overhead.
Limitation.
Wally requires a set number of clients always present in an epoch to make queries. This assumption aligns with current search engines. For example, hundreds of millions of people make queries in Google search daily [19]. We remark that for systems with few users, the overhead due to fake queries could get quite high, therefore fully oblivious protocols, like Tiptoe, for these systems would be more suitable.
2 Background
Let $[n]:=\{1,\dots,n\}$ be the set of the first $n$ natural numbers starting from one. Matrices and vectors are represented as capital boldface letters and capital lowercase, respectively, e.g., $\mathbf{M}$ and $\mathbf{v}$. The entry-wise, Hadamard, product of two vectors is denoted as $\mathbf{v}\odot\mathbf{v}^{\prime}$. All logarithms are the natural log unless noted otherwise. We denote a random variable being sampled from a probability distribution $D$ as $a\leftarrow D$. All distributions in this work are over finite sets and $a\leftarrow S$ denotes uniformly sampling from $S$ when $S$ is a discrete set. We denote the integers as $\mathbb{Z}$ and the integers modulo a positive integer $Q$ as $\mathbb{Z}_{Q}$. Let $\mathcal{NB}(r,p)$ denote the negative binomial distribution parameterized by $r,p$. Its probability mass function is $Pr(X=k)=\binom{k+r-1}{k}(1-p)^{k}p^{r}$, $X\sim\mathcal{NB}(r,p)$, for every non-negative integer $k$. An important fact about the negative binomial distribution is that it is infinitely divisible, Definition 1.
Definition 1.
A distribution $\mathcal{D}$ is infinitely-divisible if for all $n\in\mathbb{N}$, it can be expressed as the sum of $n$ i.i.d. variables. That is, there exists a distribution $\mathcal{D}_{n}$ such that $\mathcal{X}_{1}+\cdots+\mathcal{X}_{n}\sim\mathcal{D}$ and $\mathcal{X}_{i}\sim\mathcal{D}_{n}$.
2.1 (Somewhat) Homomorphic Encryption
Fully homomorphic encryption (FHE) is a special cryptosystem allowing arbitrary computation over ciphertexts. FHE for arbitrary computation is still very expensive. To achieve practical performance, somewhat homomorphic encryption (SHE, also called leveled FHE) is often used, which only supports a limited number of computations.
We focus on SHE schemes based on the Ring Learning with Errors (RLWE) problem. Concretely, Wally can be implemented using any SHE scheme such as BFV [18] or BGV [10] that share the following structure. A plaintext $m$ is a polynomial in a ring $R_{t}=\mathbb{Z}_{t}[X]/(X^{n}+1)$ (with degree at most $n-1$) and plaintext modulus $t$ is a prime or prime power. The secret key $s$ is a polynomial of degree $n-1$ with small coefficients, in $\{0,\pm 1\}$. A ciphertext is a pair of polynomials $\mathsf{ct}=(c_{0},c_{1})=(a,as+m+e)\in R_{Q}^{2}$ where $Q$ is ciphertext modulus and $R_{Q}:=\mathbb{Z}_{Q}[X]/(X^{n}+1)$. Here $a$ is picked uniform randomly and $e$ is a noise polynomial with coefficients sampled from a bounded Gaussian distribution. A scheme satisfies its decryption formula: $c_{0}+c_{1}s\mod Q=m+e$. This noise polynomial’s coefficients grow as we compute more homomorphic operations on the ciphertext, and we can only decrypt correctly if $\|e\|_{\infty}<Q/2t$. We call $\log_{2}(Q)-\log_{2}(2t)$ the parameter’s noise budget, measured in bits.
SHE parameters.
We choose parameters by first choosing a homomorphic computation, plaintext modulus, and a security parameter. This leads to a ring dimension and modulus $(n,Q)$ which 1) satisfies the security requirement and 2) allows for enough noise budget to allow decryption after homomorphic computation. Larger $Q$ gives more noise budget, but it also necessitates a larger dimension $n$ for security. The ratio of the ring dimension to the modulus bits, $n/\log_{2}Q$, stays roughly constant for a fixed security level.
Vectorized SHE.
When a plaintext modulus $t$ is a prime satisfying $t=1\mod 2n$, these SHE schemes support operations on plaintext vectors as well as polynomials. In the vectorized case, each ciphertext encrypts a plaintext vector $\mathbf{v}\in\mathbb{Z}_{t}^{n}$, and the homomorphic operations are single-instruction multiple-data (SIMD) instructions, that is, the same operation is performed on each component of the plaintext vectors. Vectorized SHE is crucial for applications requiring homomorphic linear algebra.
SHE operations and costs.
SHE schemes we consider support the following operations. Given $\mathsf{ct}=\mathsf{Enc}(\mathbf{v})$, $\mathsf{ct}^{\prime}=\mathsf{Enc}(\mathbf{v})$,
  • $\mathsf{PtCtAdd}(\mathsf{ct},\mathbf{v}^{\prime})$ returns a ciphertext encrypting $\mathbf{v}+\mathbf{v}^{\prime}$.
  • $\mathsf{CtCtAdd}(\mathsf{ct},\mathsf{ct}^{\prime})$ returns a ciphertext encrypting $\mathbf{v}+\mathbf{v}^{\prime}$.
  • $\mathsf{PtCtMult}(\mathsf{ct},\mathbf{v}^{\prime})$ returns a ciphertext encrypting $\mathbf{v}\odot\mathbf{v}^{\prime}$.
  • $\mathsf{CtCtMult}(\mathsf{ct},\mathsf{ct}^{\prime})$ returns a ciphertext encrypting $\mathbf{v}\odot\mathbf{v}^{\prime}$.
  • $\mathsf{CtRotate}(\mathsf{ct},r)$ for $r\in[0,n/2)$ returns a ciphertext encrypting
    (((( $\mathsf{Rot}(\mathbf{v}),\mathsf{Rot}(\mathbf{\bar{v}}))\in\mathbb{Z}_{p}^{n/2}\times\mathbb{Z}_{p}^{n/2}$
    where $\mathsf{Rot}(\mathbf{v})$ ($\mathsf{Rot}(\mathbf{\bar{v}})$) is $\mathbf{v}$ ($\mathbf{\bar{v}}$) cyclically shifted to the left by r positions, if $\mathsf{ct}$ originally encrypted $\mathbf{v}^{\prime}=(\mathbf{v},\mathbf{\bar{v}})$. We can homomorphically swap $(\mathbf{v},\mathbf{\bar{v}})$ with the same operation, called “conjugation”.
In Table 1 we show concrete run times and noise growth for BFV, an efficient SHE scheme. In general, addition and plaintext multiplication are efficient and addition has minimal noise growth. The latter has larger growth proportional to $t$. Ciphertext-ciphertext multiplication has large noise growth and is somewhat inefficient since it requires multiplying the ciphertext polynomials over the integers then rounding down back to $\mod Q$. Ciphertext rotation adds little noise but it is somewhat slow due to key-switching. Therefore, minimizing the number of ciphertext multiplications and ciphertext rotations is crucial for scalable applications.
Evaluation keys.
Ciphertext rotation and multiplication require evaluation keys that depend on a secret key. Additionally, each rotation requires a separate evaluation key. Usually, these keys can be reused, but as mentioned later, in Wally  the client must send fresh keys for each query. Therefore, these keys contribute to the client’s request size.
Operation Time (ms) Noise added (bits)
$\mathsf{CtCtAdd}$ $0.004$ $0.5$
$\mathsf{PtCtMult}$ $0.02$ $20$
$\mathsf{CtCtMult}$ $2.5$ $26$
$\mathsf{CtRotate}$ $0.5$ $0.5$
Table 1: Experimental computation cost and noise growth of each BFV homomorphic operation. The polynomial degree $n$ is $4096$, the ciphertext modulus $Q$ has $83$ bits, and the plaintext modulus $t=40961$ (16 bits). $\mathsf{CtRotate}$ noise is measured on repeated rotations and the others are from two input ciphertexts with the same noise levels. Time costs are measured with the swift-homomorphic-encryption library [6] (git commit b70d927) on an Intel Xeon w3-2423 using a single core.
Differential privacy.
We use the standard notion of differential privacy (DP) [15, 14]: a randomized mechanism $M$ is $(\epsilon,\delta)$-differentially private if for all $\epsilon\geq 0$ and $\delta\in[0,1]$, any two datasets $X,X^{\prime}$ differing on one element, and for any subset $S$ of possible transcripts output by $M$, $\Pr[M(X)\in S]\leq e^{\epsilon}\cdot\Pr[M(X^{\prime})\in S]+\delta$ where the probability is over the randomness of $M$. Let $\zeta$ be an output of a mechanism $M$, then the privacy loss in observing $\zeta$ between two datasets $X,X^{\prime}$ as input is
$\log\left(\frac{\Pr[M(X)=\zeta]}{\Pr[M(X^{\prime})=\zeta]}\right).$
Essentially, standard DP ensures that neighboring datasets have a privacy loss of at most $\epsilon$ with probability at least $1-\delta$ for all observations. We use basic composition in DP [36, Lemma 2.3].
Lemma 1.
If $M_{1}(X),\dots,M_{k}(X)$ are each $(\epsilon,\delta)$-DP with independently sampled randomness, then $(M_{1}(X),\dots,M_{k}(X))$ is $(k\epsilon,k\delta)$-DP.
Anonymous networks.
Anonymous networks (ANs), also called anonymization networks and anonymous communication protocols, are protocols which offer user anonymity by using cryptographic techniques such as onion routing [23], or mixnets [11]. Throughout the paper, we assume that ANs 1) strip all identifying information from incoming messages from clients, and 2) permute all received messages within a batch of messages from clients before sending them to the server.
Private information retrieval.
Though the paper is mostly presented for approximate nearest neighbor retrieval, Wally applies to exact retrieval as well by using keyword PIR for its homomorphic computation. Private information retrieval (PIR) protocols are cryptographic client-server protocols where a server holds a large database and the client obliviously retrieves an entry from the database. PIR protocols can be for index retrieval or retrieval by keyword [12]. In this paper, we focus on single-server PIR protocols [30] since we can shard the database and the client can run Wally to anonymously retrieve entries and hide its downloaded entry on the shard. We use and optimize PIR protocols based on SHE with minimal preprocessing, such as [5] and [4]. There is a notion of differentially-private PIR in the multi-server setting [34, 2], where the database is stored across non-colluding servers while guaranteeing differential privacy. This is in contrast to Wally which stores the database on one server or servers controlled by a single entity.
3 System Overview
Figure 1: High-level overview of Wally. Throughout an epoch, clients send encrypted queries, real and fake, at random time slots to an anonymization network which “washes” queries of identifying information.
Figure 1 shows a high-level design of Wally. It consists of three entities: the server, the clients, and the anonymization network. Wally runs in epochs of time; within each epoch, a fixed number of clients make queries. Within each epoch, Wally performs a search that outputs the most relevant results to client queries, similar to insecure search engines. Wally does not require any synchronization across clients. Each client can independently decide to participate or skip an epoch. However, we require a minimum number of honest clients stay online throughout an epoch and that a participating client can only make at most $\Delta$ queries within each epoch. Overall, the privacy loss for a particular client is the sum of the loss in each epoch in which the client has participated. Wally’s privacy notion also requires that each client makes at most $M$ queries in the system’s lifetime; therefore, a client can participate at most $M$ epochs. The correctness guarantee is that all the queries within epoch are processed by the end of the epoch.
In Wally each epoch is divided into time slots, each a second long, e.g., and privacy is guaranteed by the following:
  • Hiding the content of the queries (somewhat homomorphic encryption).
  • Hiding the real queries of clients among fake queries (differential privacy).
  • Removing the identifiable information from each query and shuffling the batch of queries received within a time slot (anonymization network).
  • Hiding the information revealed by queries order and time of arrival (random query schedule).
Embedding-based search.
Similar to state of the art insecure semantic search and previous secure search, Wally utilizes (vector) embeddings for semantic search. Embeddings are a machine learning-based technique that map unstructured objects like documents, images, and videos to $d$-dimensional vectors of floats. These embeddings retain the semantic relationship between objects by placing similar objects close together in a vector space. Hence, they are used to find relevant entries in search systems. Embeddings are usually much smaller than the objects themselves, which results in a significant reduction in client-to-server communication. Embeddings can be generated for different data types, including text, image, and video. Therefore, Wally can support all these data types. Wally is compatible with all the recent embedding models, such as BERT [13].
In the following sections, we introduce the insecure version of semantic search in Wally. We then highlight the privacy issues in this version. Finally, we discuss how we overcome these privacy issues.
3.1 Insecure Search Pipeline
In the offline phase, the server maps each document to the corresponding embedding. At the time of a query the client embeds the query into an embedding. Once the client query and the server documents are represented as embeddings, Wally uses an approximate nearest neighbor algorithm to find the documents most relevant to the client’s query. Our two-step approximate nearest neighbor protocol has the following flow:
  • Server initialization. At initialization, the server divides the documents (using their embeddings) into $K$ clusters using K-means clustering, a heuristic-based technique that partitions data into $K$ disjoint clusters $C_{1},\cdots,C_{K}$, and outputs these clusters and their centroid embeddings $c_{1},\cdots,c_{K}$. The server sends the centroid embeddings to the client.
  • Pick nearest centroid. To make a query, the client first locally finds the cluster nearest to the query embedding. For this, the client computes a similarity between the query and each of the centroid embeddings and picks a cluster whose centroid has high similarity.
  • Server computation. The client then sends the nearest cluster index $C^{\prime}$ and query embedding $q^{\prime}$ to the server. The server computes similarity between the query embedding and each embedding in cluster $C^{\prime}$. The server returns the similarity scores to the client.
  • Pick nearest entry. The client sorts these scores and picks the entry closest to the query.
  • Metadata retrieval. The client then fetches the metadata associated with the closest entry from the server.
The clustering technique provides a trade-off between performance and accuracy; generally, increasing the number of clusters improves the performance (reduces server computation and response size). However, it degrades the accuracy because there is a higher chance that a potential nearest neighbor is in a different cluster than the cluster with the closest centroid. Therefore, finding a balance is important for the overall usability of the system.
The above protocol is not private. To ensure that the protocol does not reveal any information about the client’s query, it must hide the following information from the server:
[…]
  • Nearest cluster. The client reveals to the server the nearest cluster ID. Recall that clustering puts highly correlated documents together in a cluster. Therefore, the cluster ID can reveal the kind of data the client holds. For example, if the client query is a song embedding, the cluster ID might reveal the genre of the song.
  • Closest entry index. To fetch the relevant metadata, the client sends the index of the closest entry to the server. Learning this could reveal server information about the query.
We show how we hide the above data below and in Section 6.
3.2 Hiding the Query Embedding
We rely on efficient vectorized, RLWE-based SHE described in Section 2 to hide a query embedding. At a high level, the client sends an SHE encryption of its embedding to the server, and the server homomorphically (under encryption) computes similarity scores between the query embedding and all the embeddings in the nearest cluster. The server then returns encrypted scores to the client. SHE computation is inherently expensive; we utilize various techniques to make similarity computation efficient. Overall, in our implementation, the client query and server response are a single RLWE ciphertext of sizes 226kB, and 53kB, respectively, (the former with evaluation keys) and it only takes 19 ms for the server to compute similarity scores for a single cluster homomorphically.
3.3 Hiding the nearest cluster
By using SHE, the content of the query is hidden from the server, but the server learns the nearest cluster. Hiding the nearest cluster is a significant performance bottleneck in previous schemes like Tiptoe. Specifically, to conceal the nearest cluster for a single query, Tiptoe requires the server to perform SHE operations against the entire database (all the clusters). Otherwise, skipping clusters will reveal to the server that the client is not interested in them. Computing over the entire database hides the nearest cluster in the most robust possible sense, but this results in a prohibitively high server computation cost.
Instead, we propose an approach based on differential privacy. The key advantage is that, in expectation, the server computes encrypted similarity scores only against a few clusters per query. Our approach results in a significant improvement in server computation. An additional advantage of our approach is a reduction in request size; concretely, the request size is only $O(d)$, around $\sqrt{N}$ smaller than Tiptoe. Conversely, differential privacy is a weaker privacy notion than the full, cryptographic obliviousness Tiptoe achieves.
The starting point of our solution is an observation that in large-scale search engines, sufficiently many clients are making the queries at any given time. Therefore, we propose a novel mechanism to hide a particular client query among the crowd of queries from all the clients. Our proposed approach provides $(\epsilon,\delta)$-differential privacy. This approach guarantees to every client that the maximum privacy loss due to its queries is essentially bounded by $\epsilon$ with all but $\delta$ probability. In other words, a client’s queries do not significantly influence the outcome of any statistical inference run by the server.
As the first step, we require all the traffic (queries and responses) to and from the server to route through the anonymization network. Intuitively, one would expect that doing this would hide the query in the crowd of queries, as queries do not carry identifiable information. However, note that only using an anonymization network is not sufficient because by observing traffic patterns, the server can learn information about a client. For example, if the server has side information that a particular client is interested in a specific cluster and all the other clients are known not to access the cluster then by observing that the cluster is accessed, the server can infer whether a request is from the client or not. Similarly, the server can exploit the arrival times of queries to infer if queries belong to a particular client. For example, the queries towards each epoch’s start could belong to a particular user.
To achieve the desired privacy, we make the following two additional changes, highlighted in Figure 3:
  • Fake queries. In addition to real queries, each client makes a few fake queries. Specifically, for each cluster, the client makes $Z\leftarrow\mathcal{D}$ fake queries, where distribution $\mathcal{D}$ is chosen to guarantee differential privacy after fake queries from many clients are aggregated. In expectation, the client makes $K\mathbb{E}[\mathcal{D}]$ fake queries. For each fake query, the client creates an SHE encryption of zero and pseudorandomly picks one of the clusters.
  • Random query schedule. The client makes queries at random instances within the epoch. For this, we assume that each epoch is divided into slots, each one second long. For each query (real or fake), the client independently picks a random slot and only makes the query at that slot.
Intuitively, these two changes suppress the leakage, as mentioned above. In Section 6, we show that using the negative binomial distribution for fake queries $\mathcal{D}$ along with random query schedule guarantees differential privacy.
3.4 Hiding Metadata Index
After learning the index of the nearest entry, the client retrieves metadata associated with it. When metadata per entry is small, the server returns metadata of the entire cluster with each query. Note that this leaks no extra information because the server already learns the cluster ID for each query. The client then locally selects the metadata associated with the most relevant entry. However, if the metadata is quite large, the client can run Keyword PIR to fetch the metadata of only the most relevant entry within a cluster.
3.5 Keyword Search
Wally also supports keyword searches in which the client is interested in retrieving data associated with a keyword. Using a random hash function, the server divides the database into $K^{\prime}$ random buckets (instead of clusters). For each query, the client uses the same hash function to find the bucket in which the keyword falls locally and then uses the efficient keyword PIR scheme to fetch a specific record within a bucket. The differential privacy part remains the same as in semantic search. That is, clients make fake queries for random buckets at a random schedule.
4 Efficient SHE-Based Instantiation
In Wally, for semantic search, SHE is used to calculate similarity scores between the query and nearest cluster embeddings, and for the keyword search, an SHE based Keyword PIR is used to fetch the entry associated with query keyword among the entries in the bucket in which query keyword falls. In this section, we discuss the details of our efficient SHE-based instantiation.
4.1 SHE details in semantic search
The server database is encoded as an array of plaintext vectors. The output is a single response ciphertext with at most $n$ encrypted similarity scores, where $n$ is the RLWE polynomial dimension.
Server database.
Recall that the server database is divided into K clusters. Each cluster consists of at most $N^{\prime}$ embeddings. For simplicity, we assume that for a given cluster $C_{i}$, the dataset is represented as a cube $D_{C_{i}}$ of dimensions $N^{\prime}/d\times d\times d$, recall that $d$ is the embedding size. Each cube consists of $N^{\prime}/d$ slices, each of which is a matrix of dimensions $d\times d$, and each column within a slice is an embedding. Denote elements of $i$’th embedding as $e_{i}=[e_{i}^{0},\cdots,e_{i}^{d-1}]$. For now, we assume that the plaintext dimension size $n=d$ and each diagonal within a slice is a separate plaintext. That is each slice is represented as $d$ plaintexts vectors, and $j$’th vector consists of entries $p_{j}=[e_{0}^{j},e_{1}^{j+1},e_{2}^{j+2},\cdots,e_{d-1}^{j+d-1}]$. This means each plaintext vector consists of a single index from each embedding in a slice in increments of one. The client encrypts the query embedding $\hat{q}=[q^{0},q^{1},\cdots,q^{d-1}]$ in a single SHE ciphertext.
Figure 2: A single slice computation in Wally using SHE operations. The server only require aligning the query once for all the slices.
SHE dot-product computation.
Recall, from Section 2, to calculate cosine similarity between two embeddings, the server needs to compute a dot-product between them. To achieve that, the server must align the query to ensure that elements of a query vector only get multiplied with corresponding element of plaintext vectors, i.e., every element-wise multiplication should be of the form $e_{j}^{i}*q^{i}$. The server will use homomorphic rotation to move the elements within the encrypted query vector to the correct alignment. Specifically, for a given slice, starting with the first plaintext vector $j=0$, the server performs vectorized ciphertext-plaintext multiplication between the query ciphertext and the first plaintext vector $j=0$. This results in encrypted $p_{0}*\hat{q}=[e_{0}^{0}*q^{0},e_{1}^{1}*q^{1},e_{2}^{2}*q^{2},\cdots,e_{d-1}^{d-1}*q^{d-1}]$. No rotation is required for first multiplication because the query is already aligned with the first vector. The elements of the second plaintext vector $j=1$ are $p_{1}=[e_{0}^{1},e_{1}^{2},e_{2}^{3},\cdots,e_{d-1}^{0}]$. Hence, the server cannot directly multiply it with the query vector. The server first homomorphically rotates the query vector one slot to the left, resulting in query vector $\hat{q}^{1}=[q^{1},q^{2},\cdots,q^{0}]$ then multiplies the result: $p_{1}*\hat{q}^{1}=[e_{0}^{1}*q^{1},e_{1}^{2}*q^{2},e_{2}^{3}*q^{3},\cdots,e_{d-1}^{0}*q^{0}]$. The server then repeats this for all the remaining plaintext vectors in a slice: rotating the query vector in increments of one and multiplying it with the plaintext vector. By the end the server has $p_{0}*\hat{q},p_{1}*\hat{q}^{1},\cdots,p_{d-1}*\hat{q}^{d-1}$ encrypted multiplications. The server then sums all of these vectors together, resulting in a single encrypted vector in which each element is a dot product between the client query vector and one of the embeddings in the slice. Figure 2 shows an example server computation for a single slice for $n=d=4$.
Database and query packing.
The above description assumes that the polynomial dimension $n$ equals embedding dimension $d$. However, $n$ is often significantly larger than $d$. In this setting, each plaintext vector could hold multiple diagonals. Our scheme packs diagonals across the slices in a single plaintext vector. Concretely, a cube’s $j$-th plaintext vector holds $(j\mod d)$-th diagonals of $\lfloor{n/d}\rfloor$ slices. In all the practical databases that we have considered, the number of slices per cube $N^{\prime}/d$ is smaller than $\lfloor{n/d}\rfloor$. Therefore, by using packing, the entire cube can be encoded with only $d$ plaintext vectors.
The client query encrypts a plaintext vector of $\lfloor{n/d}\rfloor$ repetitions of the query embedding. This query still allows the server to align the entries of query embeddings by homomorphically rotating left in increments of one. The number of ciphertext-plaintext multiplications is now $\lceil N^{\prime}/n\rceil*d$ while the response consists of $\lceil N^{\prime}/n\rceil$ ciphertexts. Again, for the databases we consider, $\lceil N^{\prime}/n\rceil\approx 1$ results in $d$ multiplications, and each response is a single ciphertext.
Recall from Section 2 that homomorphic rotation is costly. In the above explanation, the server performs $d$ rotations even after packing to align the query. In Wally  we use the standard baby-step giant-step (BSGS) optimization [24] to minimize the number of rotations: for $d$-dimensional inner products, BSGS requires only $\sqrt{d}$ rotations for step sizes of $1$ (baby step) and $\sqrt{d}$ (giant step). See Algorithm 1 in [24] for the BSGS algorithm.
Float-to-integer encoding.
Recall from Section 2 that vectorized SHE supports operations over vectors of integers while embeddings are vectors of floats. In Wally we convert each float into an integer by multiplying it with the appropriate scaling factor having the plaintext modulus be large enough to perform the computation over the integers. The choice of scaling factor depends on the size of SHE plaintext modulus $t$ and the required precision.
Figure 3: Differential privacy related changes to hide real queries. Changes are highlighted in gray.
4.2 SHE details in keyword search
The server database consists of $N$ (keyword, value) entries. Recall that the server divides these entries into $B$ buckets using a random hash function. Within each bucket, in Wally  we use a keyword PIR construction based on cuckoo hashing and index PIR.
At a high level, for each bucket independently, the server uses the cuckoo hashing to map the $N^{\prime}$ entries into a table of size $O(N^{\prime})$ using two random hash functions. The guarantee is that if a keyword is present in the database, it must be at bucket index $h_{1}(\textit{keyword})$ or $h_{2}(\textit{keyword})$ in the table. The client then uses index PIR to privately fetch entries at these indices and locally find the index containing the entry and associated data. Wally uses MulPIR [4] as the underlying index PIR, an efficient single-server PIR scheme based on SHE. We emphasize that even though we initialize Wally with MulPIR, it is compatible with any recent efficient PIR scheme.
Each server bucket consisting of $N/B$ entries is represented as a $\sqrt{N/B}\times\sqrt{N/B}$ matrix $\mathbf{D}$. MulPIR uses a polynomial version of SHE, where each plaintext is a polynomial. For this explanation, we assume that each entry has the same size as SHE plaintext. If entries are larger than a plaintext, then we split each entry across multiple plaintexts, and if entries are smaller, then we pack multiple entries into each plaintext polynomial when applicable. A client query in MulPIR is a single ciphertext $\hat{q}$ encrypting row and column indices. Given a query, the server does the following.
  1. 1.
    First expand $\hat{q}$ to two $\sqrt{N/B}$-length ciphertext vectors. Encrypting row indicator vector $\mathbf{r}$ and column indicator vector $\mathbf{c}$. The server uses the oblivious expansion algorithm given in [4] to expand. This algorithm requires $2\sqrt{N/B}$ homomorphic substitutions.
  2. 2.
    Then compute the matrix-vector product $\mathbf{a}:=\mathbf{D}\mathbf{c}$ with $N/B$ plaintext-ciphertext multiplications and $N/B-\sqrt{N/B}$ ciphertext additions.
  3. 3.
    Finish by computing $\mathbf{r}^{T}\mathbf{a}$ that results a single ciphertext that is a response. This step uses $\sqrt{N/B}$ ciphertext-ciphertext multiplications and $\sqrt{N/B}-1$ ciphertext additions. to compute.
Section A discusses various optimizations we implemented for MulPIR. Overall, our optimizations resulted in a three-fold reduction in server computation and a three-fold reduction in request size.
Algorithm 1 $\mathsf{ServerInit}$
Input: Semantic search database $\{(\mathbf{e}_{i},m_{i})\}_{l\in[N]}$, number of clusters $K$
Output: Processed semantic $D_{\text{sem}}$, cluster centroids $c$.
1:$(\mathbf{c},\mathbf{C})\leftarrow\textsf{K-MeansClustering}(\{(\mathbf{e}_{i},m_{i})\}_{l\in[N]},K)$triangleright dividing semantic data into $K$ clusters
2:for $i\in[K]$ do
3:     for $j\in[d]$ do $\triangleright$ Diagonal packing
4:         for $l\in[|C[i]|/d]$ do
5:              $\mathbf{p}[{j}]=\mathbf{p}[{j}]||[C[i,0]_{l*C[i]/d+j}^{j},$
6:
7:              $\cdots,C[i,0]_{l*C[i]/d+j+d-1}^{d-1}]$
8:         end for
9:         $\mathbf{p}[j]=\mathsf{EncodePtVec}(\mathbf{p}[j])$
10:     end for
11:     for $l\in[h]$ do $\triangleright$ Rotations for BSGS
12:         for $j\in[g]$ do
13:              $\mathbf{p}[gl+j]=\textsf{PtxtRotate}(\mathbf{p}[gl+j],-gk)$
14:         end for
15:     end for
16:     $D_{\text{sem}}[i]=p$
17:end for
18:Output $D_{\text{sem}},c$
5 Wally Protocol
In this section, we describe the complete Wally protocol for a single epoch in semantic search. We assume that $U$ honest clients are present to query within that epoch. Due to the space limitation, we only describe the semantic search protocol in detail, but keyword search follows a similar protocol with changes mentioned in Section A.
Server initialization.
The server uses Algorithm 1 to encode the database with $N$ embeddings into $K*d$ vectorized plaintexts. Specifically, the algorithm divides the input embeddings into $K$ clusters using K-means clustering. The algorithm then iterates over each cluster separately. Each cluster consists of $|C[i]|/d$ column-wise $d\times d$ slices. The algorithm then fills $d$ plaintext vectors for each cluster. In a plaintext vector $j$, the algorithm packs $j$’th diagonal of all $|C[i]|/d$ matrices. Once $d$ plaintext vectors are generated, the algorithm iterates over a group of $g=\sqrt{d}$ plaintexts, rotating each group in increasing multiples of $-g$. These rotations are required for the baby-step giant step (BSGS) optimization [24]. Note that these rotations are performed on plaintexts, so their cost are negligible.
Query generation.
Each client uses Algorithm 2 to generate queries and their schedule at the start of an epoch. The client locally has a set of cluster centroids $c$ that the server generates during initialization. The algorithm first generates independent queries for at most $\Delta$ real query embeddings. To generate a $j$’th real query, the algorithm picks a centroid $\mathsf{id}_{l}$ nearest to $j$’th query embedding. Then, the algorithm copies the embedding $n/d$ times into a plaintext vector $p$ of size $n$. This is done to take advantage of SHE packing. The algorithm then generates a fresh SHE secret key $\mathsf{sk}$ and rotation evaluation key $\mathsf{evk}$, and encrypts $p$ using $\mathsf{sk}$.
Next, the algorithm generates fake queries. Here, the algorithm iterates over each cluster, sampling a number of fake queries from the negative binomial distribution $\mathcal{NB}(r/U,p)$. To generate a single fake query, the algorithm picks a random centroid, generates a fresh SHE secret key $\mathsf{sk}$ and evaluation keys $\mathsf{evk}$, and encrypts $\mathbf{0}$ using $\mathsf{sk}$. Note that the client must generate a fresh evaluation key to keep each query anonymous. Also, each evaluation key must include a rotation key for steps $1$ and $g$, because the server computation involves rotating by these steps.
The algorithm then permutes the real and fake queries list $\mathbf{Q}$ and generates a random schedule for the queries. To generate the schedule $\mathbf{S}$, for each query $i\in Q$, the algorithm picks a random slot $t$ independently and appends $i$ to $S[t]$. In other words, the schedule $\mathbf{S}$ maps slot indices to the query indices. At a particular time slot $t$ within an epoch, the client will make queries (independently) indicated by $S[t]$.
Server computation.
The server uses Algorithm 4 to generate a response for every query received. The algorithm first aligns the query ciphertext $\mathsf{ct}$. That is, first it copies the query ciphertext $g$ times and then rotates each copy left in increments of one, in total $g$ rotations. Then, the algorithm performs a dot-product between rotated query ciphertexts and the plaintext vectors for cluster $\mathsf{id}$. Recall that we ensure each cluster is encoded into $d$ plaintexts. The algorithm iterates in groups of $g$ plaintexts; within each group, multiply the $i$-th plaintext with the $i$-th copy of the rotated query and sum the resulting ciphertexts. These iterations yield $h=\sqrt{d}$ ciphertexts. After that, the algorithm iterates over the $h$ resulting ciphertexts, rotating right each in increments of $g$ and summing all of them into a single ciphertext $R$. This step requires a total of $h$ rotations. The algorithm returns $R$ and cluster $\mathsf{id}$’s metadata.
Algorithm 2 $\mathsf{ClientQuery}$
Input: client queries $q=\{q_{i}\}_{i\in\Delta}$
Clusters centroids $c=\{\mathbf{c}_{l}\}_{l\in[K]}$
$U$ is the number of clients in the system
Output: Real and fake queries list $\mathbf{Q}$.
1:for $l\in[\Delta]$ do $\triangleright$ generate real queries
2:     $\mathbf{p}[1:N/K]=\mathbf{e}_{l}||\mathbf{e}_{l}||\cdots||\mathbf{e}_{l}$
3:     $(\mathsf{sk}_{l},\mathsf{evk}_{l})\leftarrow\mathsf{KeyGen}(1^{\lambda})$
4:     $\mathsf{ct}_{l}\leftarrow\mathsf{EncSemPt}({\mathsf{sk}_{l}},\mathbf{p})$
5:     Pick cluster $\mathsf{id}_{l}$ nearest to $\mathbf{e}_{l}$ from $c$
6:     $Q\leftarrow Q||(\mathsf{ct}_{l},\mathsf{evk}_{l},\mathsf{id}_{l})$
7:end for
8:for $i\in[K]$ do
9:     $F_{i}\leftarrow\mathcal{NB}(r/U,p)$triangleright number of fake queries
10:     for $j\in[F_{i}]$ do $\triangleright$ generate fake queries
11:         $(\mathsf{sk}^{\prime}_{j},\mathsf{evk}^{\prime}_{j})\leftarrow\mathsf{KeyGen}(1^{\lambda})$
12:         $\mathsf{ct}_{j}^{\prime}\leftarrow\mathsf{EncSemPt}(\mathsf{sk}_{j}^{\prime},\mathbf{0})$
13:         Pick random cluster $\mathsf{id}_{j}$ from $c$
14:         $Q\leftarrow Q||(\mathsf{ct}_{j}^{\prime},\mathsf{evk}_{j}^{\prime},\mathsf{id}_{j})$
15:     end for
16:end for
17:Randomly permute $Q$
18:$\mathbf{S}\leftarrow\mathsf{RandScheduleGen}(Q)$triangleright generate schedule
Algorithm 3 $\mathsf{RandScheduleGen}$
Input: List $\mathbf{Q}$ of real and fake queries
1:Epoch length is $T$
2:Initialize a schedule $\mathbf{S}$triangleright $T$ empty lists
3:for $j\in[|\mathbf{Q}|]$ do
4:     $i\leftarrow[T]$triangleright Sample random slot
5:     $\mathbf{S}[i]\leftarrow\mathbf{S}[i]||j$
6:end for
7:return $\mathbf{S}$
Algorithm 4 $\mathsf{ServerComputation}$
Input: Query $(\mathsf{ct},\mathsf{evk},\mathsf{id})$
Output: Encrypted scores and data $r$.
1:for $j\in[g]$ do $\triangleright$ Align query
2:     $\mathsf{ct}^{\prime}[j]=\mathsf{SHERotate}_{\mathsf{evk}}(\mathsf{ct},j)$
3:end for
4:$D=D_{\text{sem}}[\mathsf{id}]$triangleright Dot-Product
5:$R=\sum_{k\in[h]}\mathsf{SHERotate}_{\mathsf{evk}}(\sum_{j\in[g]}D[j+kg]*\mathsf{ct}^{\prime}[j],gk)$
6:$\textsf{data}\leftarrow\textsf{data}_{\mathsf{id}}$triangleright Cluster $\mathsf{id}$’s metadata
7:return $r:=(R,\textsf{data})$
6 Security
In this section, we prove that if honest clients follow Algorithm 2 to generate queries, then the overall system achieves $(2\epsilon,2\Delta\delta)$-differential privacy in each epoch. Recall from Algorithm 2 that each query consists of the nearest centroid and the embedding. We only focus on proving the privacy of centroids because the IND-CPA security of SHE protects query embeddings. Therefore, we model each query as a nearest centroid without an embedding in this section. During each epoch, the server receives multiple queries for each cluster. Towards the end of the epoch, the server’s view can be considered as a noisy histogram over all clusters. We consider this noisy histogram as an output of the DP mechanism.
We prove this argument in two steps: First, we show that a noisy histogram output by a curator, which gets users’ queries as input and uses negative binomial $\mathcal{NB}$ mechanism to sample fake queries, is $(2\epsilon,2\Delta\delta)$-DP in the central model. Second, we show that the view of the server in an epoch can be simulated only using the curator’s output.
6.1 DP Security in the Central Model
We define the central model as follows:
Central model.
The curator performs the following:
  1. 1.
    Collects the real queries from all the clients, $\{R_{u}\}_{u\in[U]}$, where $U$ is the total number of clients, and $R_{u}$ is each client’s set of queries and will be at most $\Delta$. Let $R$ denote the list of all real queries.
  2. 2.
    For each cluster $b$, samples the number of fake queries $F_{j}\leftarrow\mathcal{NB}(r,p)$. Add it to the fake queries list $F$.
  3. 3.
    Given the list of real and fake queries, randomly permute them: $\mathsf{list}\leftarrow\mathsf{permute}(R,F)$.
  4. 4.
    Send $\mathsf{list}$ to the server.
Another way to view it is that the curator sends a noisy histogram of queries to each cluster to the server. Recall that each client can contribute at most $\Delta$ queries.
We then prove the following theorem:
Theorem 1.
For any query database of size $N$, number of clusters $K$, differential privacy parameters $\epsilon,\delta\in(0,1)$, and $\Delta\in\mathbb{N}$, let $p=e^{-0.2\epsilon/\Delta}$ and $r=3(1+\log(1/\delta))$. Then, the negative binomial mechanism $\mathcal{NB}(r,p)$ is $(2\epsilon,2\Delta\delta)$-DP for $\Delta$-histogram.
Proof.
We define the curator as Algorithm $M$ that takes input from a database $X$ of requests from $N$ clients and outputs a noisy histogram. We prove the theorem by investigating the affected buckets when we change the input $X$ by replacing one client with another.
For a particular cluster $b$ in the histogram, we define $M_{b}(\cdot):=R_{b}+\mathcal{NB}(r,p)$, where $R_{b}$ is the total real queries for the cluster $b$ in the database. Note that this is the curator’s exact output for each cluster. We break down the proof into two edge cases: 1) the simpler case where $X$ and $X^{\prime}$ differ by a client which sends all their messages to a single cluster and 2) they differ by a client which sends $\Delta$ messages to $\Delta$ different clusters.
Simple case. Consider two neighboring databases $X$ and $X^{\prime}$ which differ in queries contributed by a single client. Concretely, in $X$ the client contributes $\Delta$ queries to a cluster $b$ and in $X^{\prime}$ instead contributes to another cluster $b^{\prime}\not=b$. Note that when $j\notin\{b,b^{\prime}\}$ then $M_{j}(R)$ and $M_{j}(R^{\prime})$ are identically distributed. Thus, the privacy provided by $M_{j}$ is 0-DP for all $j\notin\{b,b^{\prime}\}$.
Now consider the cluster $b$; the two databases differ by $\Delta$ real queries for this cluster. By the following, lemma $M_{b}$ provides $(\epsilon,\delta)$ privacy on cluster $b$.
Lemma 2 (Theorem 13 [22]).
[…]
We can make a similar argument for a cluster $b^{\prime}$. Therefore, by basic composition, Lemma 1, the total privacy provided by $M$ for these neighboring databases is $(2\epsilon,2\delta)$—DP.
General case. Now, we consider a general case where two neighboring databases that differ in queries from a single client. In database $X$ the client contributes queries $\gamma_{j}$ for cluster $j$ and in database $X^{\prime}$ the client contributes queries $\gamma_{j}^{\prime}$ for the cluster $j$. Define $\Delta_{j}:=|\gamma_{j}-\gamma_{j}^{\prime}|$. The total difference between two neighboring databases is given by $\sum_{j}^{B}\Delta_{j}\leq 2\Delta$.
Let $\epsilon_{j}:=\frac{\epsilon\Delta_{j}}{\Delta}\leq\epsilon$. Further, we only consider the clusters where $\Delta_{j}>0$ since the clusters where $\Delta_{j}=0$ are already 00-DP. (Equivalently, $\epsilon_{j}=0$ for these $j$.) Then, Lemma 2 implies cluster $j$ is $(\epsilon_{j},\delta)$-DP for $\Delta_{j}$-summation for the same mechanism since
$\frac{\epsilon}{\Delta}=\frac{\epsilon}{\Delta}\cdot\frac{\Delta_{j}}{\Delta_{j}}=\frac{\epsilon_{j}}{\Delta_{j}}.$
Simple composition over all the clusters yields
$\sum_{j}\epsilon_{j}=\frac{\epsilon}{\Delta}\sum_{j}\Delta_{j}\leq 2\epsilon.$
Basic composition also yields that the composed protocol has $\delta^{\prime}:=2\Delta\delta$ since at most $2\Delta$ clusters differ. ∎
We emphasize that the theorem is stated for a particular epoch $S$. If clients participate in many epochs, say $l$, then basic composition yields $(2l\epsilon,2l\cdot\Delta\delta)$-DP.
6.2 Generating view of the server
We show how the curator’s output is sufficient to simulate the server’s view generated by distributed DP algorithm defined in Algorithm 2 together with an anonymization network (AN). We also require pseudorandom ciphertexts and evaluation keys, a common assumption for RLWE-based SHE schemes. The latter is also why we cannot reuse evaluation keys.
Theorem 2.
Let $\mathcal{AN}$ be an anonymous network that 1) removes all identifying information from messages it receives and 2) randomly permutes all the messages it receives over a second-long slot. Let each honest client run the distributed mechanism in Algorithm 2 with messages sent to the server via $\mathcal{AN}$. Then, the view of the server in an epoch can be simulated using the curator noisy histogram output described above in the central model.
Proof.
We use infinite divisibility of $\mathcal{NB}$ distribution to produce a standard cryptography hybrid argument.
Hybrid 0: Every client is the same as the real client given in Algorithm 2 except it does not include encrypted embeddings with each request (real or fake). Hybrid 0 is equivalent to the real client finding the nearest centroid for each real query and a fake random centroid for each fake query. The client then picks a random slot for each query and send associated queries at each slot.
Hybrid 1: In this hybrid, we replace all the clients with a single simulator. The input of a simulator is a list $R=\{R_{u}\}_{u}$, where $R_{u}$ are the real queries for $u$-th client. The simulator then simulates each client $u$ with input $R_{u}$ as in Hybrid 0.
As the simulator internally runs each client $u$ as in Hybrid 0 with input $R_{u}$, from the adversary’s point of view, the distribution of requests is the same as Hybrid 0.
Hybrid 2: In this Hybrid, the simulator’s input is the same as in the previous Hybrid 1. The simulator is defined as follows.
  1. 1.
    Collects the real queries from all the clients, $\{R_{u}\}_{u\in[U]}$. Call it a list of real queries $R$.
  2. 2.
    For each cluster $b$, samples the number of fake queries $F_{j}\leftarrow\mathsf{NB}(p,r)$. Add it to the fake queries list $F$.
  3. 3.
    Given the list of real queries $R$ and fake queries $F$, randomly permutes them: $\mathsf{list}\leftarrow\mathsf{permute}(R,F)$
  4. 4.
    Initialize schedule list $S$ of size $z$ (where z is the total number of slots in an epoch). For each query $q\in\mathsf{list}$, pick a random slot $p\leftarrow z$ and set $S[p]=S[p]\cup q$.
  5. 5.
    For each slot $p\in S$ send queries in $S[p]$ to the server.
There are two differences between Hybrid 2 and Hybrid 1:
  1. 1.
    In Hybrid 1, each simulated client samples fake queries for each cluster from a distribution $\mathcal{NB}(r/U,p)$. While in Hybrid 2 (Line 2) the simulator samples fake queries per cluster from $\mathcal{NB}(r,p)$.
  2. 2.
    In Hybrid 1, each simulated client generates a schedule for its queries independently, while in Hybrid 2, the simulator generates a schedule for queries.
To show that Hybrid 2 and 1 are indistinguishable, we use the definition of infinitely divisible distribution, Definition 1.
Note that negative binomial distribution is infinite divisible because $\mathcal{NB}(r,p)=\sum_{i=1}^{U}\mathcal{NB}(r/U,p)$. Thus, the distribution of per cluster fake queries per cluster in Hybrid 2 (Line 2) is the same as in Hybrid 1. Observe that a random slot is picked for each query independently in both Hybrids. Therefore, the query schedule generated in both Hybrids is indistinguishable.
To complete the proof, note that in Line 3, the simulator generates $\mathsf{list}$ in the same way as the curator in the central model. After that, the simulator uses this $\mathsf{list}$ to generate a random query schedule, which results in a server’s view indistinguishable from the view in Hybrid 0.
7 Implementation and Performance Evaluation
In this section, we demonstrate Wally’s concrete performance via a set of experiments. We also showcase Wally’s performance tradeoffs: fake query overhead versus number of clients, fake query overhead versus the DP parameters, and the difference in low-precision performance versus high-precision performance.
7.1 Implementation
We have implemented Wally in Swift atop the Apple swift-homomorphic-encryption library [6]. Following the library, we use BFV [18] to implement the SHE parts of Wally. However, our protocol could easily be implemented using other SHE schemes such as BGV [10].
BFV parameters
The choice of BFV parameters provides a trade-off between computation and communication in Wally. At a high level, larger parameters provide better computation but result in high request size. We tried multiple sets of parameters and picked the one that provided the best performance. We set polynomial modulus $n=4096$, ciphertext modulus $\log_{2}{Q}=83$, and plaintext modulus $\log_{2}{t}=15$ bits. These parameters allow a precision of six bits. To increase the precision, we use plaintext CRT optimization described in Section A. This optimization’s key benefit is increasing precision without increasing the evaluation key size (which is a dominating factor in each query). For this we set first plaintext modulus $\log_{2}{t_{0}}=15$ bits and second plaintext $\log_{2}{t_{1}}=16$. This gives us a precision of 13 bits.
Differential privacy parameters
Differential privacy parameters provide a trade-off between privacy and performance overhead due to fake queries. For a fixed failure probability $\delta$, smaller $\epsilon$ provides stronger privacy but worse performance because each client is required to make more fake queries. We set $\delta=2^{-20}$, which means DP will fail only once a million and $\epsilon=1$, providing a decent privacy guarantee.
7.2 Performance Evaluation
Experimental setup.
We ran server-side computations on an Intel Xeon w3-2423 instance with 32GB of RAM and 6 cores.
End-to-end performance.
Metric Low prec. Mid prec. Tiptoe
Query Size (MB) 0.67 0.84 16.2
Response Size
(MB) 0.16 0.18 3.2
Time to process
queries 117 mins 139 mins 2.8 years
Precision (bits) 6 13 4
Privacy DP DP SHE
Table 2: Performance comparison of Wally and Tiptoe for database of 35 millions entries. Query and response times and sizes for Wally include the number fake query overhead (3×3$\times$3 × or two fake for each real query). Throughput for Wally is measured for 8 million real queries, or 24 million in total. All Wally responses are computed as the encrypted scores plus $44$ KB of metadata.
Table 2 describes our main experiments compared to Tiptoe. For an end-to-end system, we estimate the network time on a 30Mbps channel. This comes out to 74ms for transmitting a query and response over the network for the small precision setting (first column in Tables 2 and 3).
The server database consists of 35 million entries divided into around 8500 clusters. We keep each cluster size less than the number of plaintext slots in each BFV ciphertext, $n=4096$. We assume each client makes $\Delta=2$ queries per epoch for these experiments. Moreover, there are four million honest clients present in an epoch. This results in eight million real queries in an epoch.
For Wally, we calculated the overhead by using Theorem 2. The fake query overhead is derived by the expected number of fake queries for the negative binomial mechanism, $\mathcal{NB}(r,p)$, as
$\mathbb{E}[\text{total fake queries per client}]=\frac{rp}{1-p}\cdot\frac{K}{U}$ (1)
where $K$ is the total number of clusters and $U$ is the total number of clients. For our selected parameters, for each real query, the client is required to make just under two fake queries. Therefore, the total number of queries within an epoch is 24 million. In Table 2, we included fake query overhead in query request, response and total computation plus estimated network time in Wally. The epoch size in Wally is the total time it takes to process all the queries (real and fake).
For Tiptoe, we assume there are eight million queries the server processes one by one. Tiptoe requires running experiments on real compute nodes; due to resource limitations, we carefully estimate their results using Table 7 in [26].
Before discussing Wally’s performance, we emphasize that Wally requires the following two extra assumptions not required in Tiptoe.
  • Wally requires sufficiently many honest clients during an epoch.
  • It also needs an anonymization network that removes all identifying information from incoming messages and randomly permutes messages over a short time, e.g., one-second time slots.
In other words, Wally trades off privacy guarantees for scalability.
Wally has significantly better performance than Tiptoe across all the metrics. Specifically, the query size (including overhead due to fake queries) in Wally 19-23 times smaller than in Tiptoe. This is because the query in Wally only covers a single cluster, while Tiptoe covers all clusters. Similarly, the response size is 3.1-3.6 smaller. Wally’s efficient server computation makes it an ideal candidate for practical deployments in comparison to Tiptoe. For example, to process eight million queries, the server with 320 cores requires around 117 minutes in Wally. In Tiptoe, the same server would require more than two years to process these queries, which is unsuitable for most practical search services.
Scaling DP with $\epsilon$.
Figure 4 shows how the expected fake queries per cluster for each client scales with $\epsilon$ for one million clients and four million clients. This is derived by dividing Equation 1 by the number of clusters $K$. Note that the total number of clients has a linear effect while the dependence on $\epsilon$ is nonlinear as we decrease $\epsilon$.
Figure 4: The relationship between the number of queries per cluster per client for $500$k and $1$M total clients. We fix the number of real queries to $\Delta=2$.
Precision.
Table 3 shows how each query grows as we increase precision, i.e., as we increase the BFV parameters using our plaintext CRT optimization described in Section A. The six bit precision parameters are the same as we used in Table 2 while the $13$ bit parameters use the plaintext CRT on $\log_{2}(t_{1})=15.3$ and $\log_{2}(t_{2})=16$, $\log_{2}(t)=\log_{2}(t_{1}t_{2})=31.3$. Lastly, the 29 precision parameters use two $n=8192$ BFV parameter sets with $\log_{2}t_{1}=29$ and $\log_{2}t_{2}=30$.
Precision (bits) 6 13 29
Query (KB) 226 283 901
Response (KB) 53 63 196
Server comp. (ms) 19 35 110
Table 3: Here we list the query and response size per each query (real and fake). Each response size includes $44$kB of metadata. Multiply all numbers by $\phi+1$ if DP necessitates $\phi$ fake queries per real query.
Scaling with the number of clients.
Table 4 shows how the fake query overhead varies with the number of messages per epoch $\Delta$ and the number of clients in the system. We emphasize that the number of fake queries increases linearly with the number of buckets and decreases proportionally to the number of clients. As mentioned above, this can be offset by increasing the size of the clusters to a multiple of $n$ and we can perform the similarity computation in parallel with more threads. Table 4 emphasizes the importance of having many clients in a system since the overhead vanishes as the number of clients becomes large.
Number of Clients .5M 1M 2M 10M
Fake queries ($\Delta=1$) $7.5$ $4.8$ $2.9$ $1.4$
Exp. MB ($\Delta=1$) $1.9$ $1.08$ $0.65$ $3.11$
Fake queries ($\Delta=4$) $35.3$ $18$ $9.6$ $2.7$
Exp. MB ($\Delta=4$) $8$ $4.1$ $2.2$ $0.61$
Table 4: This table shows how the fake query overhead changes with the number of clients and the number of messages. We fix the number of clusters to $K=8545$ to match our 35M dataset, Table 2.
7.3 Keyword Query Performance
Table 5 shows single-threaded keyword PIR server performance, measured on an Intel Xeon w3-2423 CPU with 32GB of RAM using swift-homomorphic-encryption [6]. Each entry used a BFV parameter set of $n=4096$, $\log_{2}Q=83$ with three ciphertext moduli, and $\log_{2}t=6$ using MulPIR [4]. Each response size was 27KB. Note, however, that each dataset size requires manual tuning for optimal encryption and cuckoo hashing parameters.
Database entries Entry Size (Bytes) Query size (KB) Time (ms)
100k 10 368 166
100k 100 539 465
1 million 10 539 565
1 million 100 624 1700
Table 5: Keyword-PIR, or exact query, server run times for Wally. These times include the optimizations in Section A.
8 Related Work
Anonymization networks alone do not provide query privacy. Further, the server can learn the distribution of queries to each cluster when these protocols are applied to clustered databases.
Recently, there has been a number of works making private search engines using homomorphic encryption, and potentially PIR, in order to obtain full cryptographic obliviousness. Tiptoe [26], the most recent and performant, has been discussed above in-detail. Other less efficient solutions include Coeus [1] as well as Hers [17], with the latter providing privacy for the server’s database as well. Adjusting Wally to provide security for the server would decrease performance since we would need to add circuit privacy to the server’s SHE computation, raising the RLWE parameters to the next largest ring dimension $n$ for noise flooding[7, 32], and including a client public key in each query [20].
Another related area of work is differentially-private multi-server PIR schemes [34, 2]. In both works, servers learn differentially private, noised histograms over queries. The former, [34], has clients generate fake queries whereas the latter, [2], has the servers inject fake queries. Both require a secure multiparty computation over the entire database. Notably, [2] can serve one million queries in about two seconds on a 100K entry database with eight servers in a dishonest majority security model and a ten minute pre-computation. Their offline and online computation grows linearly with the size of the server’s database.
Differential privacy has also been applied to multiple anonymous messaging systems [37, 35, 31]. These are large (many server) mix-net-like protocols where the servers add in fake queries to hide user messages with a differential privacy guarantee. Though they are for a different application, their use of differential privacy is similar to Wally’s.
9 Conclusion
Wally demonstrates that private search systems can be scalable to systems with many users. Wally carefully balances somewhat homomorphic encryption together with differential privacy to hide the client’s query and reveal a differentially private histogram over all clients’ traffic to database clusters. Further, Wally’s overhead from differential privacy vanishes as the number of users in the system increases. This vanishing overhead is crucial to scalability and is potentially useful in many private systems deployable at-scale.
Appendix A Optimizations
Here we present our optimizations for both use cases: approximate nearest neighbor search and exact retrieval (PIR). Our optimizations revolve around the BFV [9, 18] and BGV [10] SHE schemes. Further, we describe several optimizations to the MulPIR scheme [4]. First, we describe the standard RNS optimization in SHE [21] and a method to compress ciphertexts by dropping LSBs as described in Cheetah [28].
RNS form.
Often $Q$ is larger than machine-size words and we use $Q$ as the product of machine-sized primes $Q=q_{0}q_{1}\cdots q_{l}$. This allows us to store a ciphertext as two polynomials each stored as an $n\times(l+1)$ matrix of integers by the Chinese remainder theorem (CRT) on $Q$. Further, we require each $q_{i}$ to be NTT friendly, $q_{i}=1\mod 2n$. This allows us to use the number theoretic transform (NTT) to efficiently switch between each evaluation form and coefficient form in time $O(n\log_{2}n)$, since the NTT is a modular version of the fast Fourier transform over a prime modulus. We can efficiently compress a ciphertext by modulus-switching down to a smaller modulus, $Q^{\prime}$ where $Q^{\prime}|Q$, e.g., $Q^{\prime}=q_{0}$, without affecting the ciphertext’s noise budget.
Dropping ciphertext LSBs.
We use Cheetah’s [28] method to compress ciphertexts after homomorphic computation: we modulus switch down to the smallest RNS modulus, $Q^{\prime}=q_{0}$, then drop the least significant bits of the ciphertext. This technique compresses the response size further than simply modulus-switching down, but it also adds noise. Cheetah incorrectly models the ciphertext as noiseless (Appendix F in [27]), but we derive the proper analysis as follows. Dropping the LSBs of a BFV ciphertext $\mathsf{ct}=(c_{0},c_{1})$ can be analyzed by expressing the polynomials as high and low bit polynomials $c_{0}=2^{b_{0}}c_{0}^{h}+c_{0}^{l},\ c_{1}=2^{b_{1}}c_{1}^{h}+c_{1}^{l}$. Then, the new decryption equation is
$2^{b_{0}}c_{0}^{h}+2^{b_{1}}c_{1}^{h}s=\lfloor Q/t\rceil m+e+e_{drop}$
for
$e_{drop}=-c_{0}^{l}-c_{1}^{l}s.$
We have correctness as long as $\|e+e_{drop}\|_{\infty}<Q^{\prime}/2t$. Concretely, we are able to drop around $8$kB from the response size in approximate nearest neighbors search to a response size of $47$kB when combined with our BFV plaintext CRT optimization described in the next subsection.
A.1 Approximate Nearest Neighbors Optimizations
The homomorphic computation in the approximate nearest neighbors use case is $C$ inner products over real embedding vectors satisfying $\mathbf{e}$, normalized to $\|\mathbf{e}\|_{2}=1$ in our use case, where $C$ is the maximum cluster size. We scale $\mathbf{e}$’s entries to integers, round them to a vector $\mathbf{\widetilde{e}}$, and perform the inner product over the integers. The computation is correct as long as the inner product does not wrap around mod $t$. We define plaintext precision as $\approx\log_{2}(t)/2$ since the inner product has a multiplicative depth of one.
Minimal rotation keys and bandwidth.
Wally uses the “baby-step giant-step” (BSGS) optimization [24] to minimize the number of rotation keys: for $d$-dimensional inner products, BSGS requires $2\sqrt{d}$ rotations for step sizes of $1$ (baby) and $\sqrt{d}$ (giant). See Algorithm 1 in [24] for the BSGS algorithm. Additionally, we are able to use two rotation keys since $d<n/2$ in our use case. This optimization leads us to a query size of one ciphertext $\mathsf{ct}$ and two rotation keys $\mathsf{evk}=(\mathsf{evk}_{1},\mathsf{evk}_{\sqrt{d}})$. For key-switching, and relinearization in PIR below, we use the hybrid-GHS [21] key-switching strategy described in-detail by Kim et al. [29].
Say we require a plaintext modulus of $t\approx 2^{29}$ for $\sim 14$ bits of plaintext precision. The smallest BFV ring dimension supporting this is $n=8192$ which requires a evaluation key modulus of $Q\approx 2^{165}$, ciphertext modulus of $Q^{\prime}\approx 2^{110}$ for 128 bits of post-quantum security [3] [^3] https://github.com/malb/lattice-estimator. A query consists of $2$ polynomials for the ciphertext and $2$ polynomials for the evaluation keys, or $225$kB and $676$kB respectively, for a total query size of $901$kB for these parameters.
Re-using BFV evaluation keys via the plaintext CRT.
Wally achieves minimal query bandwidth per client query by a novel use of the Chinese remainder theorem (CRT) on the BFV plaintext space. The CRT on plaintexts allows increase the plaintext precision without increasing evaluation key size since BFV’s evaluation keys are independent of the plaintext modulus. Therefore, we encrypt a vector $\mathbf{\widetilde{v}}$ as $\mathbf{\widetilde{v}}\mod t_{0}$ and $\mathbf{\widetilde{v}}\mod t_{1}$ for NTT-friendly primes $\{t_{0},t_{1}\}$. The computation is correct if the inner products do not wrap around modulo $t=t_{0}t_{1}$, but the individual ciphertexts can have their computation wrap around modulo each $t_{i}$.
In general, plaintext CRT is preferable to increasing the ring dimension since the evaluation key sizes dominate the ciphertext size: 6 polynomials versus 2 polynomials. This is true in the bandwidth-optimal key-switching strategy for hybrid GHS key-switching. Another advantage of the plaintext CRT is that it allows for very high precision computations which do not wrap around mod $t$. One limitation of this optimization is that we run out of NTT-friendly plaintext primes for small parameter sets, like $n=4096$ or $2048$.
A.2 PIR Optimizations
Uneven dimensions.
In general, the cost of a ciphertext-ciphertext multiplication is much higher than a ciphertext rotation and both are much higher than addition or plaintext multiplication. For example, Table 1 shows ciphertext addition and plaintext multiplication as less than $100\mu s$ while ciphertext rotation is $0.5ms$ and ciphertext multiplication is $2.5ms$.
Therefore, we structure the database (cluster) as a rectangle $\mathbf{D}\in R_{t}^{d_{1}\times d_{2}}$ with uneven dimensions $d_{1}$ and $d_{2}$. For any fixed HE parameter set, the server’s compute time is dominated by ciphertext multiplications and rotations and that the former is a constant factor $\gamma$ more than the latter. This simplification yields a compute time of $2d_{1}t_{r}+2d_{2}t_{r}+d_{2}t_{\times}$ where $t_{\times}$ and $t_{r}$ are the times to multiply and rotate ciphertexts, respectively. More simply, the total compute time, measured in terms of rotations, is
$(\gamma+2)d_{1}+2d_{2}=(\gamma+2)C/d_{2}+2d_{2}$
rotations. Minimizing this as a function in $d_{2}$ gives $d_{2}=\sqrt{1+\gamma/2}\sqrt{C}$.
This optimization saw an improvement of $30-40\%$ in MulPIR run times. We chose this concrete analysis since SHE parameter regimes are restricted per use case and it is much simpler than an asymptotic analysis.
Linearizing parts of query expansion
MulPIR is an optimized query expansion which originated with Angel et al. [5]. In short, a query index is encoded in a plaintext polynomial then expanded into $k$ ciphertexts by calling $k$ ciphertext rotation operations (Galois automorphisms followed by key-switching). We noticed that some of these expansions can be substituted with linear operations: call $\mathsf{ct}\pm\mathsf{ct}^{\prime}$ where $\mathsf{ct}^{\prime}$ is $\mathsf{ct}$ rotated. In general, we saw a $10-25\%$ improvement in MulPIR’s expansion from our optimization.
Lazy rescaling in BFV multiplication.
We apply the lazy rescaling technique of Kim et al [29] in MulPIR’s last step, the inner product between ciphertexts $\mathbf{r}^{T}\mathbf{a}$. The main idea here is that BFV multiplications first multiply polynomials over the integers, then scale and round back to integers modulo $Q$. Kim et al [29, Appendix F] noticed that one can multiply over the integers, add over the integers, then scale and round after the additions, from the dimension number of scaling operations to one. Lazy rescaling yielded a $15-20\%$ improvement in our MulPIR implementation.
Optimization for very large entries
In MulPIR the database is represented as a plaintext matrix $\mathbf{D}\in R_{t}^{d_{1}\times d_{2}}$. Conventionally, the response is computed as
[…]
This approach results in $d_{1}\cdot d_{2}\cdot n$ ciphertext-plaintext multiplications and $d_{2}\cdot t$ ciphertext-ciphertext multiplications, where $t$ is the number of plaintexts used to encode one entry. The number of ciphertext-ciphertext multiplications grows linearly with $t$ and remains a concretely small value when entries are not large. However, it will blow up with extremely large entries (e.g., when they are high-resolution photos).
To overcome this challenge, we can swap the order of the computation as
[…]
The outer-product takes $d_{1}\cdot d_{2}$ ciphertext-ciphertext multiplications to compute, which is independent of the entry size, while the inner products still take $d_{1}\cdot d_{2}\cdot n$ ciphertext-plaintext multiplications. Therefore, this optimization can reduce the computation overhead when the entry size ($t$ times the plaintext size) is larger than the first dimension size $d_{1}$.
Keyword PIR.
MulPIR’s keyword PIR functionality is given by using cuckoo hashing to find the entry’s proper index. We noticed for our cluster sizes $C$, we are able to split the cluster into two hash tables and perform $1$-hash cuckoo hashing on each without increasing the (empirical) failure probability. This resulted in a 2×2$\times$2 × improvement in query size and server computation time.
Low-level SHE Optimizations
BFV’s multiplication step over the integers is done by extending the modulus $Q$ to a larger modulus $PQ$ so the multiplication does not wrap around. We use Bajard et al.’s method to extend the basis, “fast basis conversion” in [8] and re-use the basis elements in $Q$ for a $\sim 15\%$ improvement in BGV multiplication.
We add a conditional, lazy modular reduction to Harvey’s NTT [25]. The main idea is that we only reduce during the butterfly if the integers would exceed the machine word size. This saw a $20-25\%$ improvement in forward NTT times and $15-20\%$ improvement in inverse NTT times for $61$-bit primes. Note, we saw no improvement on primes near $64$-bits.