### v6

# Scalable Private Search with Wally

Abstract
This paper presents Wally, a private search system that supports efficient search queries against large databases. When sufficiently many clients are making queries, Wally’s performance is significantly better than previous systems while providing a standard privacy guarantee of $(\epsilon,\delta)$-differential privacy.
Specifically, for a database with 3.2 million entries, Wally’s queries per second (QPS) is 7-28x higher, and communication is 6.69-31x smaller than Tiptoe, a state-of-the-art private search system. In Wally, each client adds a few fake queries and sends each query via an anonymous network to the server at independently chosen random instants. We also use somewhat homomorphic encryption (SHE) to reduce the communication size.
The number of fake queries each client makes depends inversely on the number of clients making queries. Therefore, the overhead of fake queries vanishes as the number of honest clients increases, enabling scalability to millions of queries and large databases.
I Introduction
Consider a scenario where a client holding a potentially private search query wants to search information from a large database hosted on a server. This scenario reflects the flow of current search engines such as Google, Bing, and DuckDuckGo. These search engines improve the quality of our lives by promptly providing helpful information. Nevertheless, they require learning the client’s query in order to respond accurately, hence providing no query privacy. In some situations, learning the query could reveal sensitive and personal information about the client. For example, consider the queries “High blood pressure with late-stage HIV?”, “Closest diabetes clinic from Albuquerque”, and “Best divorce lawyers”. Revealing these queries to search engines could be damaging to the client. However, even if the queries do not appear sensitive, they may allow the server to learn personal information like medical information, marital status, sexuality, etc. Therefore, enabling query privacy is critical for search engines.
Private search systems prioritize user privacy. There is a long list of works that have explored protecting query privacy, by periodically injecting fake queries that resemble real queries [1, 2, 3, 4, 5] or obfuscating users’ original queries using decoy terms [6, 7, 8, 9]. These approaches are quite efficient. However, they lack a standard privacy guarantee and often rely on assumptions that limit adversarial capabilities. Many systems assume adversaries have restricted knowledge—whether unaware of the obfuscation algorithm, lacking background data, or using a predefined inference strategy. As a result, when these assumptions are removed or a more capable adversary is considered, fake queries become identifiable, ultimately breaking query privacy [10, 11, 12, 13].
Recent privacy search systems do provide the strongest privacy guarantee - the server does not learn any information about the client query irrespective of knowledge and capability, but these schemes incur very high server computation and communication which makes them impractical. For example, Tiptoe [14], a state-of-the-art private search system, requires a communication of around 17 MB and provides queries per second (QPS) of only 909, with 10,000 cores, for a database with 3.2 million entries. The high computational cost is because, for each client query, the server must scan the entire database; otherwise, it will learn which database entries the client is not interested in. Similarly, the client must send cryptographic material for each database entry, which results in a high communication cost [^1] In reality, these systems trade off response size with request size by using a standard database clustering trick. We detail this trade-off in a later section..
The world of private search thus is in an undesirable state. On one hand, traditional approaches do not provide a standard privacy guarantee, making them vulnerable; on the other hand, schemes that provide a strong privacy guarantee have very high computational and communication requirements, due to which they cannot be used in large-scale systems.
Our system.
We present Wally, a privacy preserving search system that provides standard privacy guarantees while maintaining efficiency. Our key observation is that in search engines, there are always enough clients making queries. Using this observation we explore a new trade-off between efficiency and privacy. Specifically, in Wally some small information is leaked to the server, but this leakage is proven to preserve differential privacy for the participating users. In return, Wally’s computational and communication overhead reduces significantly compared to the state of the art. Specifically, for a database with 3.2 million entries, Wally’s QPS is 7$\sim$28x higher and communication is 6.69$\sim$31x smaller than Tiptoe. Additionally, the search quality (presented as MRR@@@@100 scores) of Wally is decently better than Tiptoe.
Wally relaxes its privacy notion to $(\epsilon,\delta)$-differential privacy. The privacy guarantee of schemes like Tiptoe is stronger than Wally. However, differential privacy is an accepted standard for strong privacy [15]. Intuitively, every client in Wally is guaranteed that leakage to the server does not change by too much when their input is modified or removed. More formally, the server’s ability to infer their queries does not improve by more than a factor of $e^{\epsilon}$ , except with probability at most $\delta$.
We emphasize that we do not make any other assumption about the client queries or the adversary’s capabilities. Additionally, the privacy guarantee only gracefully degrades when the adversary sees query history.
Summary of techniques.
Like Tiptoe, Wally also uses standard search techniques based on semantic embeddings. At a high level, the embedding model maps any text string into a floating-point vector of dimension $d$ with a semantic guarantee: embeddings corresponding to similar texts will have a high inner product score (cosine similarity). This way, the search problem is reduced to the problem of private nearest neighbors: The client must find the embedding vector that maximizes the inner product with the query vector.
The server divides the database into $K$ clusters and sends cluster centroids to each client. The system works in epochs. The epoch length is picked to ensure that enough honest clients are present to make queries and there are enough server resources to process all the queries by the end of an epoch. Within each epoch, a large number of honest users make queries. We discuss this assumption in more detail towards the end of the introduction.
For each query, the client finds $\Delta$ closest clusters and forwards them to the server. The server responds by sending all the embeddings in those clusters. The queries arrive at the server via an anonymous network, ensuring that all identifiable information, such as IP address, is removed from them. The server then generates a response to each query independently and sends it back to the respective client via an anonymous network. Although queries are anonymized, this does not guarantee query privacy. The server can infer information about the client’s query by exploiting the access pattern (if a particular cluster is accessed) or arrival time (when the server receives queries). To prevent this, Wally implements the following two changes at the client.
First, each client in Wally makes fake queries. To generate a single fake query, the client picks a random cluster. The client does not have to interact with the server to generate random queries and generates them independently. The number of fake queries is sampled from a negative binomial distribution to ensure $(\epsilon,\delta)$-differential privacy. Looking ahead, to maintain privacy, the client must send only a single cluster per query. We emphasize that fake queries do not affect the accuracy or correctness seen by any client, as the client can discard the responses due to these queries.
Second, Wally also ensures that neither the arrival time nor the order of the queries leak any information about the client to the server. To achieve this, instead of sending a query instantly, each client sends queries in randomly sampled one-second slots within an epoch. Additionally, an anonymous network collects the requests received within a particular slot and sends them as a batch to the server.
We show that anonymizing the queries, adding fake queries, and randomizing the query schedule is equivalent to sharing a noised histogram of queries from all the clients with the server. We then prove that this histogram guarantees $(\epsilon,\delta)$ privacy.
As mentioned above, for each query, the server response contains all the entries in the cluster. For large clusters, this would result in a high response overhead. We utilize lattice-based, homomorphic encryption (SHE) to reduce the response overhead. Specifically, for each real query, the client freshly encrypts the query embedding under SHE and sends it along with the corresponding cluster to the server. For a fake query, the client encrypts just a zero embedding. To compute the distance score, the server computes the dot-product between the encrypted embeddings and the cluster embeddings under SHE. This reduces the response because the size of encrypted distance scores is significantly smaller than the embedding size.
Formal contributions. Our contributions are as follows:
  • An efficient private search system, named Wally. A novel system that meticulously achieves $(\epsilon,\delta)$ differential privacy for client queries. It achieves this by carefully stitching various steps to enhance privacy. Moreover, we provides a formal security proof that demonstrates the protocol’s security against a semi-honest adversary.
  • Novel optimizations enable efficient computation of distance scores under SHE.
  • A standalone open-source Swift library that includes an implementation of the BFV SHE scheme and various privacy primitives, including Private dot-product and Keyword Private Information Retrieval. This library is highly versatile and can be easily integrated into various privacy applications.
  • We compare Wally with two state-of-art schemes and demonstrate how Wally provides a better trade-off between privacy and performance.
Epoch size in Wally.
Wally works in epochs, and we assume there are enough clients present in an epoch. Note that large epochs allow considering more clients. This assumption aligns with current search engines. For example, hundreds of millions of people make queries in current search systems daily [16]. So we could consider a day-size epoch with millions of users. However, note that epoch size also dictates the client experience; long epochs mean some clients will have to wait till the end of epoch to get response to their queries. Picking the very short epoch is not desirable as well. As few clients will present, this means high overhead due to fake queries. Concretely, within an epoch, on average, each client makes $O(\frac{C\Delta\log(1/\delta)}{U\epsilon})$ fake queries, where $U$ is the number of clients within an epoch, and $C$ is the number of clusters on the server. Moreover, $\Delta$ is the maximum number of clusters each client can probe in an epoch.
Our experiments assume epochs of length 1-10 minutes, with that at least 100,000 to 500,000 clients are present to make queries within any epoch. Malicious clients can affect privacy by not submitting the fake queries. In our privacy calculation, we assume that at least half the clients are honest and exclude fake queries from malicious clients.
We remark that Tiptoe does not have this requirement. Therefore, for systems with few clients (like a few hundred users), the overhead from fake queries will be relatively high; therefore, for such a situation, Tiptoe would be more suitable.
We further discuss these assumptions in Section VIII.
II Background
Let $[n]:=\{1,\dots,n\}$ be the set of the first $n$ natural numbers starting from one. Matrices and vectors are represented as capital boldface letters and capital lowercase, respectively, e.g., $\mathbf{M}$ and $\mathbf{v}$. The entry-wise, Hadamard, product of two vectors is denoted as $\mathbf{v}\odot\mathbf{v}^{\prime}$. All logarithms are the natural log unless noted otherwise. We denote a random variable $a$ being sampled from a probability distribution $D$ as $a\leftarrow D$. All distributions in this work are over finite sets and $a\leftarrow S$ denotes uniformly sampling from $S$ when $S$ is a discrete set. We denote the integers as $\mathbb{Z}$ and the integers modulo a positive integer $Q$ as $\mathbb{Z}_{Q}$. Let $\mathcal{NB}(r,p)$ denote the negative binomial distribution parameterized by $r,p$. Its probability mass function is $Pr(X=k)=\binom{k+r-1}{k}(1-p)^{k}p^{r}$, $X\sim\mathcal{NB}(r,p)$, for every non-negative integer $k$. An important fact about the negative binomial distribution is that it is infinitely divisible, Definition 1.
Definition 1.
A distribution $\mathcal{D}$ is infinitely divisible if for all $n\in\mathbb{N}$, it can be expressed as the sum of $n$ i.i.d. variables. That is, there exists a distribution $\mathcal{D}_{n}$ such that $\mathcal{X}_{1}+\cdots+\mathcal{X}_{n}\sim\mathcal{D}$ and $\mathcal{X}_{i}\sim\mathcal{D}_{n}$.
II-A (Somewhat) Homomorphic Encryption
Fully homomorphic encryption (FHE) is a special cryptosystem allowing arbitrary computation over ciphertexts. FHE for arbitrary computation is still very expensive. To achieve practical performance, somewhat homomorphic encryption (SHE, also called leveled FHE) is often used, which only supports a limited number of computations.
We focus on SHE schemes based on the ring learning with errors (RLWE) problem. Concretely, Wally can be implemented using any SHE scheme such as BFV [17] or BGV [18] that share the following structure. A plaintext $m$ is a polynomial in a ring $R_{t}=\mathbb{Z}_{t}[X]/(X^{n}+1)$ (with degree at most $n-1$) and plaintext modulus $t$ is a prime or prime power. The secret key $s$ is a polynomial of degree $n-1$ with small coefficients, in $\{0,\pm 1\}$. A ciphertext is a pair of polynomials $\mathsf{ct}=(c_{0},c_{1})=(a,as+m+e)\in R_{Q}^{2}$ where $Q$ is ciphertext modulus and $R_{Q}:=\mathbb{Z}_{Q}[X]/(X^{n}+1)$. Here $a$ is picked uniform randomly and $e$ is a noise polynomial with coefficients sampled from a bounded Gaussian distribution. A scheme satisfies its decryption formula: $c_{0}+c_{1}s\mod Q=m+e$. This noise polynomial’s coefficients grow as we compute more homomorphic operations on the ciphertext, and we can only decrypt correctly if $\|e\|_{\infty}<Q/2t$. We call $\log_{2}(Q)-\log_{2}(2t)$ the parameter’s noise budget, measured in bits.
SHE parameters.
We choose parameters by first choosing a homomorphic computation, plaintext modulus, and a security parameter. This leads to a ring dimension and modulus $(n,Q)$ which 1) satisfies the security requirement and 2) allows for enough noise budget to allow decryption after homomorphic computation. Larger $Q$ gives more noise budget, but it also necessitates a larger dimension $n$ for security. The ratio of the ring dimension to the modulus bits, $n/\log_{2}Q$, stays roughly constant for a fixed security level.
Vectorized SHE.
When a plaintext modulus $t$ is a prime satisfying $t=1\mod 2n$, these SHE schemes support operations on plaintext vectors as well as polynomials. In the vectorized case, each ciphertext encrypts a plaintext vector $\mathbf{v}\in\mathbb{Z}_{t}^{n}$, and the homomorphic operations are single-instruction multiple-data (SIMD) instructions, that is, the same operation is performed on each component of the plaintext vectors. Vectorized SHE is crucial for applications requiring homomorphic linear algebra.
SHE operations and costs.
SHE schemes we consider support the following operations. Given $\mathsf{ct}=\mathsf{Enc}(\mathbf{v})$, $\mathsf{ct}^{\prime}=\mathsf{Enc}(\mathbf{v})$,
  • $\mathsf{PtCtAdd}(\mathsf{ct},\mathbf{v}^{\prime})$ returns a ciphertext encrypting $\mathbf{v}+\mathbf{v}^{\prime}$.
  • $\mathsf{CtCtAdd}(\mathsf{ct},\mathsf{ct}^{\prime})$ returns a ciphertext encrypting $\mathbf{v}+\mathbf{v}^{\prime}$.
  • $\mathsf{PtCtMult}(\mathsf{ct},\mathbf{v}^{\prime})$ returns a ciphertext encrypting $\mathbf{v}\odot\mathbf{v}^{\prime}$.
  • $\mathsf{CtCtMult}(\mathsf{ct},\mathsf{ct}^{\prime})$ returns a ciphertext encrypting $\mathbf{v}\odot\mathbf{v}^{\prime}$.
  • $\mathsf{CtRotate}(\mathsf{ct},r)$ for $r\in[0,n/2)$ returns a ciphertext encrypting
    (((( $\mathsf{Rot}(\mathbf{v}),\mathsf{Rot}(\mathbf{\bar{v}}))\in\mathbb{Z}_{p}^{n/2}\times\mathbb{Z}_{p}^{n/2}$
    where $\mathsf{Rot}(\mathbf{v})$ ($\mathsf{Rot}(\mathbf{\bar{v}})$) is $\mathbf{v}$ ($\mathbf{\bar{v}}$) cyclically shifted to the left by r positions, if $\mathsf{ct}$ originally encrypted $\mathbf{v}^{\prime}=(\mathbf{v},\mathbf{\bar{v}})$. We can homomorphically swap $(\mathbf{v},\mathbf{\bar{v}})$ with the same operation, called “conjugation”.
In Table I we show concrete run times and noise growth for BFV, an efficient SHE scheme. In general, addition and plaintext multiplication are efficient and addition has minimal noise growth. The latter has larger growth proportional to $t$. Ciphertext-ciphertext multiplication has large noise growth and is somewhat inefficient since it requires multiplying the ciphertext polynomials over the integers then rounding down back to $\mod Q$. Ciphertext rotation adds little noise but it is somewhat slow due to key-switching. Therefore, minimizing the number of ciphertext multiplications and ciphertext rotations is crucial for scalable applications.
Evaluation keys.
BFV multiplication, rotation, and substitution operations require evaluation keys that depend on the client’s secret key. The server could store the evaluation keys once in typical SHE applications and reuse them across various client requests. However, the server learns the requests belonging to the same client. Looking ahead, this will break the privacy guarantee in Wally  therefore, the client must send fresh evaluation keys with each request. Therefore, these keys contribute to the client’s request size.
Operation Time (ms) Noise added (bits)
$\mathsf{CtCtAdd}$ $0.004$ $0.5$
$\mathsf{PtCtMult}$ $0.02$ $20$
$\mathsf{CtCtMult}$ $2.5$ $26$
$\mathsf{CtRotate}$ $0.5$ $0.5$
TABLE I: Experimental computation cost and noise growth of each BFV homomorphic operation. The polynomial degree $n$ is $4096$, the ciphertext modulus $Q$ has $83$ bits, and the plaintext modulus $t=40961$ (16 bits). $\mathsf{CtRotate}$ noise is measured on repeated rotations and the others are from two input ciphertexts with the same noise levels. Time costs are measured with the swift-homomorphic-encryption library [19] (git commit b70d927) on an Intel Xeon w3-2423 using a single core.
0
Differential privacy.
We use the standard notion of differential privacy (DP) [20, 21]: a randomized mechanism $M$ is $(\epsilon,\delta)$-differentially private for $\epsilon\geq 0$ and $\delta\in[0,1]$, if for any two datasets $X,X^{\prime}$ differing on one element, and for any subset $S$ of possible transcripts output by $M$, $\Pr[M(X)\in S]\leq e^{\epsilon}\cdot\Pr[M(X^{\prime})\in S]+\delta$ where the probability is over the randomness of $M$. Let $\zeta$ be an output of a mechanism $M$, then the privacy loss in observing $\zeta$ between two datasets $X,X^{\prime}$ as input is
$\log\left(\frac{\Pr[M(X)=\zeta]}{\Pr[M(X^{\prime})=\zeta]}\right).$
Essentially, standard DP ensures that neighboring datasets have a privacy loss of at most $\epsilon$ with probability at least $1-\delta$ for all observations. We use basic composition in DP [22, Lemma 2.3].
Lemma 1.
If $M_{1}(X),\dots,M_{k}(X)$ are each $(\epsilon,\delta)$-DP with independently sampled randomness, then $(M_{1}(X),\dots,M_{k}(X))$ is $(k\epsilon,k\delta)$-DP.
Anonymous networks.
Anonymous networks (ANs), also called anonymization networks and anonymous communication protocols, are protocols which offer user anonymity by using cryptographic techniques such as onion routing [23], or mixnets [24]. Throughout the paper, we assume that ANs 1) strip all identifying information from incoming queries from clients, and 2) collect all of them received in a second and send them in a batch to the server.
Private information retrieval.
Though the paper is mostly presented for semantic search, Wally applies to keyword search as well by using keyword variant of private information retrieval (PIR). PIR protocols are cryptographic client-server protocols where a server holds a large database and the client can privately retrieves an entry corresponding to a keyword from the database. In this paper, we focus on protocols based on a single server [25]. There is also a notion of differentially-private PIR in the multi-server setting [26, 27], where the database is stored across non-colluding servers while guaranteeing differential privacy. We can not directly use these protocols in Wally which stores the database on one server or multiple servers controlled by a single entity.
II-B Threat Model
We break Wally’s entities into three groups: clients, the anonymous network’s (AN) hops, and the server. The server and AN hops are modeled as semi-honest adversaries attempting to learn a client’s query. Clients can behave maliciously, intentionally deviating from the protocol, in an attempt to learn an honest client’s query. The server’s database is assumed to be public.
Assumptions.
  • Clients: We assume there are at least two honest clients. Note that malicious client could break privacy of the protocol by looking ahead, as described in Theorem 1 and in VIII. Wally needs many more honest clients for better performance, as looking ahead since Wally relies on differential privacy. Malicious clients can collude with each other, an anonymous network hop, or the server. If they collude with an anonymous network hop, then we assume they collude with the server and a hop that is not the first hop.
  • Anonymous network hops: We assume there are at least two hops, and that at least one hop is honest. This is a standard assumption in anonymous networks. For example, oblivious HTTP [28] has two hops where the second hop is assumed to be operated by the server and the first hop is assumed to be operated independently by a different party.
  • Server: The server can collude with any anonymous network hop besides the entry hop, as well as clients. We assume the entry hop cannot collude with the server since the server would learn which client sent which queries, fake and real.
  • Cryptographic assumptions: Lastly, we assume the CPA security of RLWE-based SHE [29, 17].
Client privacy.
Under the above assumptions, Wally protects client query utilizing two privacy notions:
  • The client’s queries in Wally are protected under CPA-security of Wally’s SHE scheme.
  • Which clusters the client queries is protected under a differential privacy guarantee, Theorem 1.
Overall, Wally provides a differential privacy guarantee on the client’s queries because of the latter bullet above. Further, we make no assumptions on the server’s knowledge of the expected query distribution across all clients.
III Private search with inefficient communication
Figure 1: High-level overview of Wally. Throughout an epoch, clients send encrypted queries, real and fake, at random time slots to an anonymization network which “removes” identifying information.
In this section we provide a warm-up protocol that provides the required privacy but has high communication overhead. This protocol act as stepping stone towards our final protocol discussed in Section V.
Embedding.
Similar to state of the art insecure semantic search and previous secure search, This protocol utilizes embeddings for search. Embeddings are a machine learning-based technique that map unstructured objects like text, images, and videos to $d$-dimensional vectors of floating-point numbers. These embeddings retain the semantic relationship between objects by placing similar objects close together in a vector space. Hence, they are used to find relevant entries in search systems. Embeddings can be generated for different data types, including text, image, and video. Therefore, our protocol can support all these data types. Wally is compatible with recent embedding models, such as BERT [30].
Clustering based nearest neighbors search.
The server has a database of $N$ entries, where each entry consists of a document such as text or image and a bit string of associated metadata. In the offline phase, the server maps each document to the corresponding embedding. At the time of a query, the client embeds the query into an embedding. Once the client’s query and the server’s documents are represented as embeddings, the client’s objective is to identify the nearest server embedding and its associated metadata. Finding the exact nearest neighbor is computationally expensive even in an insecure environment. Therefore, we utilize an approximate nearest neighbor algorithm using clustering to find the most relevant documents to the client’s query. The protocol has the following flow:
  • Initialization. At initialization, the server divides the document embeddings into $K$ clusters using $K$-means clustering, a heuristic-based technique that partitions data into $K$ disjoint clusters $C_{1},\cdots,C_{K}$, and outputs these clusters and their centroid embeddings $c_{1},\cdots,c_{K}$. The server also divides the document metadata into the same clusters. The server sends the centroid embeddings to the client.
  • Query. To make a query, the client first locally finds $\Delta$ clusters nearest to the query embedding $C^{\prime}_{1},\cdots,C^{\prime}_{\Delta}$. For this, the client computes a cosine similarity between the query and each of the centroid embeddings and picks clusters whose centroid has the highest similarity.
  • Response. The client then sends the nearest clusters $C^{\prime}_{1},\cdots,C^{\prime}_{\Delta}$ to the server. The server responds by sending the embeddings and their metadata in clusters $C^{\prime}_{1},\cdots,C^{\prime}_{\Delta}$ to the client.
  • Local computation. The client locally computes cosine similarity between the query embedding and each embedding in clusters $C^{\prime}_{1},\cdots,C^{\prime}_{1}={\Delta}$, sorts the scores to find the closest embedding, and the corresponding metadata.
The clustering provides a trade-off between performance and accuracy; generally, increasing the number of clusters improves the performance (reduces server computation and response size). However, it degrades the accuracy because there is a higher chance that a potential nearest neighbor is missed. Therefore, finding a balance is important for the overall usability of the system.
This protocol is not private. To prevent the protocol from disclosing any information about the client’s query, it must conceal the closest clusters from the server. Remember that clustering groups highly correlated documents together in a cluster. Consequently, knowing the nearest clusters could reveal the information about the client’s query.
III-A Hiding nearest clusters using differential privacy
Hiding the nearest clusters is a significant performance bottleneck in previous schemes like Tiptoe. For instance, to conceal a single nearest cluster, Tiptoe requires the client to send an encrypted query for each cluster; similarly, it requires the server to perform SHE operations against all the clusters. This is because omitting clusters will indicate to the server that the client is not interested in them. Computing over the entire database hides the nearest cluster completely, but this results in a prohibitively high communication and computation.
We propose an approach based on differential privacy. The key advantage is that touching all the clusters is not required, which results in a significant improvement in communication and server computation than Tiptoe. We emphasize that differential privacy is a different privacy notion than the full obliviousness Tiptoe achieves; however, it is still a standard privacy notion. Intuitively, our proposed approach provides $(\epsilon,\delta)$-differential privacy. This approach guarantees to every client that the maximum privacy loss due to its queries is essentially bounded by $e^{\epsilon}$ with all but $\delta$ probability. In other words, a client’s queries do not significantly influence the outcome of any statistical inference run by the server.
The starting point of our solution is an observation that in large-scale search engines, sufficiently many clients are making queries at any given time. Therefore, we can hide a particular client query among the batch of queries from all the clients. Our approach require following steps.
Epochs.
As the first, we require the system to run in epochs of time, each one second long; within each epoch, a fixed number of clients make queries. This does not require any synchronization across clients. Each client can independently decide to participate or skip an epoch. However, we assume a minimum number of honest clients stay online throughout an epoch and that a participating client can only make at most $\Delta$ queries within each epoch. We also require that the client send $\Delta$ clusters as independent queries to the server instead of sending them together. The worst privacy loss for a particular client is the sum of the loss in each epoch in which the client has participated. The correctness guarantee is that all the queries within an epoch are processed by the end of the epoch.
Anonymized queries.
We also require all the queries to be anonymized. For this, each client will send the queries via an anonymization network.
Note that even though queries are anonymized, by observing traffic patterns, the server can still learn information about the queries. For example, if the server has side information that a particular client is interested in a specific cluster and all the other clients are known not to access the cluster, then by observing that the cluster is accessed, the server can infer whether a request is from the client or not. Therefore, we require deniability of queries. Similarly, the server can exploit the arrival times of queries to infer if queries belong to a particular client. For example, the queries towards each epoch’s start could belong to a particular user. Therefore, we need to suppress leakage through the timing channel.
To achieve the desired properties, we make the following two additional changes, highlighted in Figure 3:
  • Fake queries. In addition to real queries, each client makes a few fake queries. Specifically, for each cluster, the client makes $Z\leftarrow\mathcal{D}$ fake queries, where distribution $\mathcal{D}$ is chosen to guarantee differential privacy after fake queries from many clients are aggregated. In expectation, the client makes $K\leftarrow\mathbb{E}[\mathcal{D}]$ fake queries..
  • Random query schedule. The client makes queries at random instances within the epoch. For this, we assume that each epoch is divided into slots, each one second long. For each query (real or fake), the client independently picks a random slot and only makes the query at that slot.
Intuitively, these two changes suppress the leakage, as mentioned above. In Section VI, we show that using the negative binomial distribution for fake queries $\mathcal{D}$ along with random query schedule guarantees $(\epsilon,\delta)$ differential privacy.
Drawback of warm-up protocol.
Even though the protocol conceals the closest clusters, it incurs high communication costs. Concretely, the server transmits embeddings and metadata for all the nearest clusters back to the client, leading to substantial communication overhead.
IV Optimizing the communication using SHE
We use SHE to reduce the communication of our warm-up protocol. In this section, we discuss the details of our efficient SHE-based instantiation.
IV-A Reducing communication in nearest neighbor
[…]
We rely on vectorized, RLWE-based SHE described in Section II to hide a query embedding. At a high level, the client sends an SHE encryption of its embedding to the server, and the server homomorphically (under encryption) computes similarity scores between the query embedding and all the embeddings in the nearest cluster. The server then returns encrypted scores to the client. SHE computation is inherently expensive; we utilize various techniques to make similarity computation efficient.
Server database.
Recall that the server database is divided into K clusters. Each cluster consists of at most $N^{\prime}$ embeddings. For simplicity, we assume that for a given cluster $C_{i}$, the dataset is represented as a cube $D_{C_{i}}$ of dimensions $N^{\prime}/d\times d\times d$, recall that $d$ is the embedding size. Each cube consists of $N^{\prime}/d$ slices, each of which is a matrix of dimensions $d\times d$, and each column within a slice is an embedding. Denote elements of $i$’th embedding as $e_{i}=[e_{i}^{0},\cdots,e_{i}^{d-1}]$. For now, we assume that the plaintext dimension size $n=d$ and each diagonal within a slice is a separate plaintext. That is each slice is represented as $d$ plaintexts vectors, and $j$’th vector consists of entries $p_{j}=[e_{0}^{j},e_{1}^{j+1},e_{2}^{j+2},\cdots,e_{d-1}^{j+d-1}]$. This means each plaintext vector consists of a single index from each embedding in a slice in increments of one. The client encrypts the query embedding $\hat{q}=[q^{0},q^{1},\cdots,q^{d-1}]$ in a single SHE ciphertext.
Figure 2: A single slice computation in Wally using SHE operations. The server only require aligning the query once for all the slices.
Secure dot-product computation.
From Section II, to calculate cosine similarity between two normalized embeddings, the server needs to compute a dot-product between them. To achieve that, the server must align the query to ensure that elements of a query vector only get multiplied with corresponding element of plaintext vectors, i.e., every element-wise multiplication should be of the form $e_{j}^{i}*q^{i}$. The server will use homomorphic rotation to move the elements within the encrypted query vector to the correct alignment. Specifically, for a given slice, starting with the first plaintext vector $j=0$, the server performs vectorized ciphertext-plaintext multiplication between the query ciphertext and the first plaintext vector $j=0$. This results in encrypted $p_{0}*\hat{q}=[e_{0}^{0}*q^{0},e_{1}^{1}*q^{1},e_{2}^{2}*q^{2},\cdots,e_{d-1}^{d-1}*q^{d-1}]$. No rotation is required for first multiplication because the query is already aligned with the first vector. The elements of the second plaintext vector $j=1$ are $p_{1}=[e_{0}^{1},e_{1}^{2},e_{2}^{3},\cdots,e_{d-1}^{0}]$. Hence, the server cannot directly multiply it with the query vector. The server first homomorphically rotates the query vector one slot to the left, resulting in query vector $\hat{q}^{1}=[q^{1},q^{2},\cdots,q^{0}]$ then multiplies the result: $p_{1}*\hat{q}^{1}=[e_{0}^{1}*q^{1},e_{1}^{2}*q^{2},e_{2}^{3}*q^{3},\cdots,e_{d-1}^{0}*q^{0}]$. The server then repeats this for all the remaining plaintext vectors in a slice: rotating the query vector in increments of one and multiplying it with the plaintext vector. By the end the server has $p_{0}*\hat{q},p_{1}*\hat{q}^{1},\cdots,p_{d-1}*\hat{q}^{d-1}$ encrypted multiplications. The server then sums all of these vectors together, resulting in a single encrypted vector in which each element is a dot product between the client query vector and one of the embeddings in the slice. Figure 2 shows an example server computation for a single slice for $n=d=4$.
Database and query packing.
The above description assumes that the polynomial dimension $n$ equals embedding dimension $d$. However, $n$ is often significantly larger than $d$. In this setting, each plaintext vector could hold multiple diagonals. Our scheme packs diagonals across the slices in a single plaintext vector. Concretely, a cube’s $j$-th plaintext vector holds $(j\mod d)$-th diagonals of $\lfloor{n/d}\rfloor$ slices. In all the practical databases that we have considered, the number of slices per cube $N^{\prime}/d$ is smaller than $\lfloor{n/d}\rfloor$. Therefore, by using packing, the entire cube can be encoded with only $d$ plaintext vectors.
The client query encrypts a plaintext vector of $\lfloor{n/d}\rfloor$ repetitions of the query embedding. This query still allows the server to align the entries of query embeddings by homomorphically rotating left in increments of one. The number of ciphertext-plaintext multiplications is now $\lceil N^{\prime}/n\rceil*d$ while the response consists of $\lceil N^{\prime}/n\rceil$ ciphertexts. Again, for the databases we consider, $\lceil N^{\prime}/n\rceil\approx 1$ results in $d$ multiplications, and each response is a single ciphertext.
Recall from Section II that homomorphic rotation is costly. In the above explanation, the server performs $d$ rotations even after packing to align the query. In Wally  we use the standard baby-step giant-step (BSGS) optimization [32] to minimize the number of rotations: for $d$-dimensional inner products, BSGS requires only $\sqrt{d}$ rotations for step sizes of $1$ (baby step) and $\sqrt{d}$ (giant step). See Algorithm 1 in [32] for the BSGS algorithm.
Figure 3: Differential privacy related changes to hide real queries. Changes are highlighted in gray.
IV-B Metadata Fetch
After determining the index of the nearest entry, the client retrieves metadata related to it. When the metadata per entry is small, the server responds by returning the metadata for the entire cluster with each query. It’s important to note that this doesn’t leak any additional information since the server already knows the query cluster. The client then locally selects the metadata associated with the most relevant entry. However, if the metadata is large (in kilobytes), it becomes infeasible to download the entire cluster. In this case, the client obliviously retrieves the metadata of only the most relevant entry within a single cluster. To achieve this, we use Keyword private information retrieval (PIR) scheme based on cuckoo hashing and index PIR.
Concretely, after getting encrypted scores, the client finds the most relevant embedding locally. Use the identity of this entry to generate a Keyword PIR query. The metadata query is then clustered in which this embedding falls, as well as the Keyword PIR query.
Protecting privacy leakage.
Note that, it is still required to protect the query cluster. Therefore, clients will require participating in an additional epoch with $\Delta=1$. For this epoch also, we required queries to be anonymized, clients generate fake queries, and execute queries at random slots within an epoch.
Details of Keyword PIR.
The server encode each entry into Keyword-value pair. Where keyword equals entry ID and value is the metadata. For ease of explanation we assume that each cluster on average consist of $N^{\prime}$ entries. The server uses the cuckoo hashing to map the $N^{\prime}$ entries into a table of size $O(N^{\prime})$ using two random hash functions. The guarantee is that if a keyword is present in the database, it must be at entry index $h_{1}(\textit{keyword})$ or $h_{2}(\textit{keyword})$ in the table. The client then uses index PIR to privately fetch entries at these two indices and locally find the index containing the entry and associated data. Wally uses MulPIR [33] as the underlying index PIR, an efficient single-server PIR scheme based on SHE. We emphasize that even though we initialize Wally with MulPIR, it is compatible with any recent efficient PIR scheme.
Details of MulPIR.
The server represent cuckoo table into $\sqrt{N^{\prime}}\times\sqrt{N^{\prime}}$ matrix For this explanation, we assume that each metadata entry has the same size as SHE plaintext. If entries are larger than a plaintext, then we split each entry across multiple plaintexts, and if entries are smaller, then we pack multiple entries into each plaintext polynomial when applicable. A client query in MulPIR is a single ciphertext $\hat{q}$ encrypting row and column indices. Given a query, the server does the following.
  1. 1.
    First expand $\hat{q}$ to two $\sqrt{N^{\prime}}$-length ciphertext vectors. Encrypting row indicator vector $\mathbf{r}$ and column indicator vector $\mathbf{c}$. The server uses the oblivious expansion algorithm given in [33] to expand. This algorithm requires $\sqrt{N^{\prime}}$ homomorphic substitutions.
  2. 2.
    Then compute the matrix-vector product $\mathbf{a}:=\mathbf{D}\mathbf{c}$ with $N$ plaintext-ciphertext multiplications and ciphertext additions.
  3. 3.
    Finish by computing $\mathbf{r}^{T}\mathbf{a}$ that results a single ciphertext that is a response. This step uses $\sqrt{N^{\prime}}$ ciphertext-ciphertext multiplications and additions.
Section A discusses various optimizations we implemented for MulPIR. Overall, our optimizations resulted in a three-fold reduction in server computation and a three-fold reduction in request size.
Algorithm 1 $\mathsf{ServerInit}$
Input: Semantic search database $\{(\mathbf{e}_{i},m_{i})\}_{l\in[N]}$, number of clusters $K$
Output: Processed semantic $D_{\text{sem}}$, cluster centroids $c$.
1:$(\mathbf{c},\mathbf{C})\leftarrow\textsf{K-MeansClustering}(\{(\mathbf{e}_{i},m_{i})\}_{l\in[N]},K)$triangleright dividing semantic data into $K$ clusters
2:for $i\in[K]$ do
3:     for $j\in[d]$ do $\triangleright$ Diagonal packing
4:         for $l\in[|C[i]|/d]$ do
5:              $\mathbf{p}[{j}]=\mathbf{p}[{j}]||[C[i,0]_{l*C[i]/d+j}^{j},$
6:
7:              $\cdots,C[i,0]_{l*C[i]/d+j+d-1}^{d-1}]$
8:         end for
9:         $\mathbf{p}[j]=\mathsf{EncodePtVec}(\mathbf{p}[j])$
10:     end for
11:     for $l\in[h]$ do $\triangleright$ Rotations for BSGS
12:         for $j\in[g]$ do
13:              $\mathbf{p}[gl+j]=\textsf{PtxtRotate}(\mathbf{p}[gl+j],-gk)$
14:         end for
15:     end for
16:     $D_{\text{sem}}[i]=p$
17:end for
18:Output $D_{\text{sem}},c$
V Wally Protocol
In this section, we describe the complete Wally protocol for a single epoch in semantic search. We assume that $U$ honest clients are present to query within that epoch. Due to the space limitation, we only describe the semantic search protocol in detail, but keyword search follows a similar protocol with changes mentioned in Section A.
Server initialization.
The server uses Algorithm 1 to encode the database with $N$ embeddings into $K*d$ vectorized plaintexts. Specifically, the algorithm divides the input embeddings into $K$ clusters using K-means clustering. The algorithm then iterates over each cluster separately. Each cluster consists of $|C[i]|/d$ column-wise $d\times d$ slices. The algorithm then fills $d$ plaintext vectors for each cluster. In a plaintext vector $j$, the algorithm packs $j$’th diagonal of all $|C[i]|/d$ matrices. Once $d$ plaintext vectors are generated, the algorithm iterates over a group of $g=\sqrt{d}$ plaintexts, rotating each group in increasing multiples of $-g$. These rotations are required for the baby-step giant step (BSGS) optimization [32]. Note that these rotations are performed on plaintexts, so their cost are negligible.
Query generation.
Each client uses Algorithm 2 to generate queries and their schedule at the start of an epoch. The client locally has a set of cluster centroids $c$ that the server generates during initialization. The algorithm first generates independent queries for at most $\Delta$ real query embeddings. To generate a $j$’th real query, the algorithm picks a centroid $\mathsf{id}_{l}$ nearest to $j$’th query embedding. Then, the algorithm copies the embedding $n/d$ times into a plaintext vector $p$ of size $n$. This is done to take advantage of SHE packing. The algorithm then generates a fresh SHE secret key $\mathsf{sk}$ and rotation evaluation key $\mathsf{evk}$, and encrypts $p$ using $\mathsf{sk}$.
Next, the algorithm generates fake queries. Here, the algorithm iterates over each cluster, sampling a number of fake queries from the negative binomial distribution $\mathcal{NB}(r/U,p)$. To generate a single fake query, the algorithm picks a random centroid, generates a fresh SHE secret key $\mathsf{sk}$ and evaluation keys $\mathsf{evk}$, and encrypts $\mathbf{0}$ using $\mathsf{sk}$. Note that the client must generate a fresh evaluation key to keep each query anonymous. Also, each evaluation key must include a rotation key for steps $1$ and $g$, because the server computation involves rotating by these steps.
The algorithm then permutes the real and fake queries list $\mathbf{Q}$ and generates a random schedule for the queries. To generate the schedule $\mathbf{S}$, for each query $i\in Q$, the algorithm picks a random slot $t$ independently and appends $i$ to $S[t]$. In other words, the schedule $\mathbf{S}$ maps slot indices to the query indices. At a particular time slot $t$ within an epoch, the client will make queries (independently) indicated by $S[t]$.
Server computation.
The server uses Algorithm 4 to generate a response for every query received. The algorithm first aligns the query ciphertext $\mathsf{ct}$. That is, first it copies the query ciphertext $g$ times and then rotates each copy left in increments of one, in total $g$ rotations. Then, the algorithm performs a dot-product between rotated query ciphertexts and the plaintext vectors for cluster $\mathsf{id}$. Recall that we ensure each cluster is encoded into $d$ plaintexts. The algorithm iterates in groups of $g$ plaintexts; within each group, multiply the $i$-th plaintext with the $i$-th copy of the rotated query and sum the resulting ciphertexts. These iterations yield $h=\sqrt{d}$ ciphertexts. After that, the algorithm iterates over the $h$ resulting ciphertexts, rotating right each in increments of $g$ and summing all of them into a single ciphertext $R$. This step requires a total of $h$ rotations. The algorithm returns $R$ and cluster $\mathsf{id}$’s metadata.
Algorithm 2 $\mathsf{ClientQuery}$
Input: client queries $q=\{q_{i}\}_{i\in\Delta}$
Clusters centroids $c=\{\mathbf{c}_{l}\}_{l\in[K]}$
$U$ is the number of clients in the system
Output: Real and fake queries list $\mathbf{Q}$.
1:for $l\in[\Delta]$ do $\triangleright$ generate real queries
2:     $\mathbf{p}[1:N/K]=\mathbf{e}_{l}||\mathbf{e}_{l}||\cdots||\mathbf{e}_{l}$
3:     $(\mathsf{sk}_{l},\mathsf{evk}_{l})\leftarrow\mathsf{KeyGen}(1^{\lambda})$
4:     $\mathsf{ct}_{l}\leftarrow\mathsf{EncSemPt}({\mathsf{sk}_{l}},\mathbf{p})$
5:     Pick cluster $\mathsf{id}_{l}$ nearest to $\mathbf{e}_{l}$ from $c$
6:     $Q\leftarrow Q||(\mathsf{ct}_{l},\mathsf{evk}_{l},\mathsf{id}_{l})$
7:end for
8:for $i\in[K]$ do
9:     $F_{i}\leftarrow\mathcal{NB}(r/U,p)$triangleright number of fake queries
10:     for $j\in[F_{i}]$ do $\triangleright$ generate fake queries
11:         $(\mathsf{sk}^{\prime}_{j},\mathsf{evk}^{\prime}_{j})\leftarrow\mathsf{KeyGen}(1^{\lambda})$
12:         $\mathsf{ct}_{j}^{\prime}\leftarrow\mathsf{EncSemPt}(\mathsf{sk}_{j}^{\prime},\mathbf{0})$
13:         Pick random cluster $\mathsf{id}_{j}$ from $c$
14:         $Q\leftarrow Q||(\mathsf{ct}_{j}^{\prime},\mathsf{evk}_{j}^{\prime},\mathsf{id}_{j})$
15:     end for
16:end for
17:Randomly permute $Q$
18:$\mathbf{S}\leftarrow\mathsf{RandScheduleGen}(Q)$triangleright generate schedule
Algorithm 3 $\mathsf{RandScheduleGen}$
Input: List $\mathbf{Q}$ of real and fake queries
1:Epoch length is $T$
2:Initialize a schedule $\mathbf{S}$triangleright $T$ empty lists
3:for $j\in[|\mathbf{Q}|]$ do
4:     $i\leftarrow[T]$triangleright Sample random slot
5:     $\mathbf{S}[i]\leftarrow\mathbf{S}[i]||j$
6:end for
7:return $\mathbf{S}$
Algorithm 4 $\mathsf{ServerComputation}$
Input: Query $(\mathsf{ct},\mathsf{evk},\mathsf{id})$
Output: Encrypted scores and data $r$.
1:for $j\in[g]$ do $\triangleright$ Align query
2:     $\mathsf{ct}^{\prime}[j]=\mathsf{SHERotate}_{\mathsf{evk}}(\mathsf{ct},j)$
3:end for
4:$D=D_{\text{sem}}[\mathsf{id}]$triangleright Dot-Product
5:$R=\sum_{k\in[h]}\mathsf{SHERotate}_{\mathsf{evk}}(\sum_{j\in[g]}D[j+kg]*\mathsf{ct}^{\prime}[j],gk)$
6:$\textsf{data}\leftarrow\textsf{data}_{\mathsf{id}}$triangleright Cluster $\mathsf{id}$’s metadata
7:return $r:=(R,\textsf{data})$
VI Security
In this section, we prove that if honest clients follow Algorithm 2 to generate queries, then the overall system achieves $(2\epsilon,2\Delta\delta)$-differential privacy in each epoch. Recall from Algorithm 2 that each query consists of the nearest centroid and the embedding. We only focus on proving the privacy of centroids because the IND-CPA security of SHE protects query embeddings. Therefore, we model each query as a nearest centroid without an embedding in this section. During each epoch, the server receives multiple queries for each cluster. Towards the end of the epoch, the server’s view can be considered as a noisy histogram over all clusters. We consider this noisy histogram as an output of the DP mechanism.
We prove this argument in two steps: First, we show that a noisy histogram output by a curator, which gets users’ queries as input and uses negative binomial $\mathcal{NB}$ mechanism to sample fake queries, is $(2\epsilon,2\Delta\delta)$-DP in the central model. Second, we show that the view of the server in an epoch can be simulated only using the curator’s output.
VI-A DP Security in the Central Model
We define the central model as follows:
Central model.
The curator performs the following:
  1. 1.
    Collects the real queries from all the clients, $\{R_{u}\}_{u\in[U]}$, where $U$ is the total number of clients, and $R_{u}$ is each client’s set of queries and will be at most $\Delta$. Let $R$ denote the list of all real queries.
  2. 2.
    For each cluster $b$, samples the number of fake queries $F_{j}\leftarrow\mathcal{NB}(r,p)$. Add it to the fake queries list $F$.
  3. 3.
    Given the list of real and fake queries, randomly permute them: $\mathsf{list}\leftarrow\mathsf{permute}(R,F)$.
  4. 4.
    Send $\mathsf{list}$ to the server.
Another way to view it is that the curator sends a noisy histogram of queries to each cluster to the server. Recall that each client can contribute at most $\Delta$ queries.
We then prove the following theorem:
Theorem 1.
For any query database of size $N$, number of clusters $K$, differential privacy parameters $\epsilon,\delta\in(0,1)$, and $\Delta\in\mathbb{N}$, let $p=e^{-0.2\epsilon/\Delta}$ and $r=3(1+\log(1/\delta))$. Then, the negative binomial mechanism $\mathcal{NB}(r,p)$ is $(2\epsilon,2\Delta\delta)$-DP for $\Delta$-histogram.
Proof.
We define the curator as Algorithm $M$ that takes input from a database $X$ of requests from $N$ clients and outputs a noisy histogram. We prove the theorem by investigating the affected buckets when we change the input $X$ by replacing one client with another.
For a particular cluster $b$ in the histogram, we define $M_{b}(\cdot):=R_{b}+\mathcal{NB}(r,p)$, where $R_{b}$ is the total real queries for the cluster $b$ in the database. Note that this is the curator’s exact output for each cluster. We break down the proof into two edge cases: 1) the simpler case where $X$ and $X^{\prime}$ differ by a client which sends all their messages to a single cluster and 2) they differ by a client which sends $\Delta$ messages to $\Delta$ different clusters.
Simple case. Consider two neighboring databases $X$ and $X^{\prime}$ which differ in queries contributed by a single client. Concretely, in $X$ the client contributes $\Delta$ queries to a cluster $b$ and in $X^{\prime}$ instead contributes to another cluster $b^{\prime}\not=b$. Note that when $j\notin\{b,b^{\prime}\}$ then $M_{j}(R)$ and $M_{j}(R^{\prime})$ are identically distributed. Thus, the privacy provided by $M_{j}$ is 0-DP for all $j\notin\{b,b^{\prime}\}$.
Now consider the cluster $b$; the two databases differ by $\Delta$ real queries for this cluster. By the following, lemma $M_{b}$ provides $(\epsilon,\delta)$ privacy on cluster $b$.
Lemma 2 (Theorem 13 [34]).
[…]
We can make a similar argument for a cluster $b^{\prime}$. Therefore, by basic composition, Lemma 1, the total privacy provided by $M$ for these neighboring databases is $(2\epsilon,2\delta)$—DP.
General case. Now, we consider a general case where two neighboring databases that differ in queries from a single client. In database $X$ the client contributes queries $\gamma_{j}$ for cluster $j$ and in database $X^{\prime}$ the client contributes queries $\gamma_{j}^{\prime}$ for the cluster $j$. Define $\Delta_{j}:=|\gamma_{j}-\gamma_{j}^{\prime}|$. The total difference between two neighboring databases is given by $\sum_{j}^{B}\Delta_{j}\leq 2\Delta$.
Let $\epsilon_{j}:=\frac{\epsilon\Delta_{j}}{\Delta}\leq\epsilon$. Further, we only consider the clusters where $\Delta_{j}>0$ since the clusters where $\Delta_{j}=0$ are already 00-DP. (Equivalently, $\epsilon_{j}=0$ for these $j$.) Then, Lemma 2 implies cluster $j$ is $(\epsilon_{j},\delta)$-DP for $\Delta_{j}$-summation for the same mechanism since
$\frac{\epsilon}{\Delta}=\frac{\epsilon}{\Delta}\cdot\frac{\Delta_{j}}{\Delta_{j}}=\frac{\epsilon_{j}}{\Delta_{j}}.$
Simple composition over all the clusters yields
$\sum_{j}\epsilon_{j}=\frac{\epsilon}{\Delta}\sum_{j}\Delta_{j}\leq 2\epsilon.$
Basic composition also yields that the composed protocol has $\delta^{\prime}:=2\Delta\delta$ since at most $2\Delta$ clusters differ. ∎
We emphasize that the theorem is stated for a particular epoch $S$. If clients participate in many epochs, say $l$, then basic composition yields $(2l\epsilon,2l\cdot\Delta\delta)$-DP.
VI-B Generating view of the server
We show how the curator’s output is sufficient to simulate the server’s view generated by distributed DP algorithm defined in Algorithm 2 together with an anonymization network (AN). We also require pseudorandom ciphertexts and evaluation keys, a common assumption for RLWE-based SHE schemes. The latter is also why we cannot reuse evaluation keys.
Theorem 2.
Let $\mathcal{AN}$ be an anonymous network that 1) removes all identifying information from messages it receives and 2) randomly permutes all the messages it receives over a second-long slot. Let each honest client run the distributed mechanism in Algorithm 2 with messages sent to the server via $\mathcal{AN}$. Then, the view of the server in an epoch can be simulated using the curator noisy histogram output described above in the central model.
Proof.
We use infinite divisibility of $\mathcal{NB}$ distribution to produce a standard cryptography hybrid argument.
Hybrid 0: Every client is the same as the real client given in Algorithm 2 except it does not include encrypted embeddings with each request (real or fake). Hybrid 0 is equivalent to the real client finding the nearest centroid for each real query and a fake random centroid for each fake query. The client then picks a random slot for each query and send associated queries at each slot.
Hybrid 1: In this hybrid, we replace all the clients with a single simulator. The input of a simulator is a list $R=\{R_{u}\}_{u}$, where $R_{u}$ are the real queries for $u$-th client. The simulator then simulates each client $u$ with input $R_{u}$ as in Hybrid 0.
As the simulator internally runs each client $u$ as in Hybrid 0 with input $R_{u}$, from the adversary’s point of view, the distribution of requests is the same as Hybrid 0.
Hybrid 2: In this Hybrid, the simulator’s input is the same as in the previous Hybrid 1. The simulator is defined as follows.
  1. 1.
    Collects the real queries from all the clients, $\{R_{u}\}_{u\in[U]}$. Call it a list of real queries $R$.
  2. 2.
    For each cluster $b$, samples the number of fake queries $F_{j}\leftarrow\mathsf{NB}(p,r)$. Add it to the fake queries list $F$.
  3. 3.
    Given the list of real queries $R$ and fake queries $F$, randomly permutes them: $\mathsf{list}\leftarrow\mathsf{permute}(R,F)$
  4. 4.
    Initialize schedule list $S$ of size $z$ (where z is the total number of slots in an epoch). For each query $q\in\mathsf{list}$, pick a random slot $p\leftarrow z$ and set $S[p]=S[p]\cup q$.
  5. 5.
    For each slot $p\in S$ send queries in $S[p]$ to the server.
There are two differences between Hybrid 2 and Hybrid 1:
  1. 1.
    In Hybrid 1, each simulated client samples fake queries for each cluster from a distribution $\mathcal{NB}(r/U,p)$. While in Hybrid 2 (Line 2) the simulator samples fake queries per cluster from $\mathcal{NB}(r,p)$.
  2. 2.
    In Hybrid 1, each simulated client generates a schedule for its queries independently, while in Hybrid 2, the simulator generates a schedule for queries.
To show that Hybrid 2 and 1 are indistinguishable, we use the definition of infinitely divisible distribution, Definition 1.
Note that negative binomial distribution is infinite divisible because $\mathcal{NB}(r,p)=\sum_{i=1}^{U}\mathcal{NB}(r/U,p)$. Thus, the distribution of per cluster fake queries per cluster in Hybrid 2 (Line 2) is the same as in Hybrid 1. Observe that a random slot is picked for each query independently in both Hybrids. Therefore, the query schedule generated in both Hybrids is indistinguishable.
To complete the proof, observe that in Line 3 of Hybrid 2, the simulator generates $\mathsf{list}$ in a manner similar to the curator in the central model. However, the key difference lies in the scheduling of queries. The simulator sends queries from $\mathsf{list}$ at random intervals, while the curator in the central model sends $\mathsf{list}$ all at once to the server. Despite the adversary potentially recognizing its interaction with the simulator, it gains no additional advantage because the random schedule remains independent of secret inputs from the user.
VII Implementation
We have implemented Wally in Swift and Python languages. We next discuss implementation details for each component of Wally
VII-A Open Source SHE library
Components of Wally outlined in Section IV are instantiated using the Brakerski-Fan-Vercauteren (BFV) encryption scheme [17]. We have released the SHE components of Wally as an open-source Swift library named Apple Swift Homomorphic Encryption Library available at https://github.com/apple/swift-homomorphic-encryption. It includes the BFV scheme, a secure dot-product for search, and Keyword PIR for metadata fetch. We have implemented many optimizations within the library, which significantly improve the overall performance and can be of independent interest. The library API is simple and flexible, and can be used for other applications. We also provide an easy to follow tutorial and examples to use the API.
We discuss a few key optimizations and implementation decisions:
  • Residue Number System (RNS) variant of BFV. In BFV scheme, we manipulate elements modulo ciphertext modulus $Q$, which is an integer with hundreds of bits. Implementing this scheme requires slow multi-precision modulo arithmetic. To address this, we instead implemented the RNS variant of BFV [35]. In this variant, $Q$ is chosen as $Q=q_{1},\ldots,q_{l}$, and each $q_{i}$ is a single-precision integer of typically 28 or 55 bits. By using the RNS, any integer $x\in Z_{Q}$ can be represented as $\{x_{1}=x\mod q_{1},\cdots,x_{l}=x\mod q_{l}\}$ and the operation on $x\mod Q$ can be performed by performing the same operation on each RNS component $x_{i}$ modulus $q_{i}$. At the end we can use CRT to extract original ciphertext in $Z_{Q}$. We also implemented the Hybrid RNS variant of key-switching [36] and fast base conversion [37] methods required for $\mathsf{CtCtMult}$ and $\mathsf{CtRotate}$ operations.
  • Reducing the cost of BFV operations. Number Theoretic Transformation (NTT) speeds up polynomial multiplications in BFV operations. We implemented NTT using Harvey’s butterfly [38] and optimized it using lazy modular reduction. In every butterfly operation, we track multiplications and only perform modulus when the result overflows the machine word size. This optimization improves forward and backward NTT by $20-25\%$. In BFV multiplication, we perform tensor product and addition in large basis, then scale it back to modulus $Q$. When performing many BFV multiplications, we skip the scaling step after each multiplication and perform it only once at the end [39]. This results in $10-20\%$ improvement in MulPIR used for fetching metadata.
  • Reducing the size of response ciphertexts. At the end of the Wally protocol, the server sends response ciphertexts back to the client. The client just decrypts these ciphertexts and will not use them for further computation. Due to this, we can use the standard modulus switching technique to reduce the size of these ciphertexts by scaling down to the RNS limb $q_{l}$. In BFV after mod switch, the message is encoded into higher $\log_{2}{t}$ bits. Therefore, we can drop the least significant $\log_{2}{q_{l}/t}$ bits of the response ciphertexts and still correctly decrypt the message. However, doing so does add some noise in the resulting ciphertext. We carefully monitor this noise and drop the bits that still allow correct decryption. In [40], the authors suggested that we can drop $l_{0}$ and $l_{1}$ bits from the first and second ciphertext polynomials as long as $z\sqrt{2\frac{n}{9}}2^{l_{1}}+2^{l_{0}}<2^{\log_{2}{(q_{l}/t)}}$, where $z$ represents a score of uniform distribution. The decryption error is bounded with z score probability. We picked $z=8$ that gives decryption error of $2^{-37.5}$. We set $l_{0}=\log_{2}{(q_{l}/t)}-2$, and calculates $l_{1}\leq\lfloor\log_{2}(q/t)-2-\log_{2}(z\sqrt{2n/9})\rfloor$. If $l_{1}\leq 1$ we update $l_{0}=l-1$ and $l_{1}=0$.
    Note that we assume that the original error in the ciphertext is smaller than the error introduced by dropping bits.
  • Reducing the request size. As mentioned in Section II, evaluation keys are essential for SHE operations. In more detail, during the search, the server utilizes evaluation keys to rotate and merge ciphertexts after multiplication. To minimize key size, the client transmits only two keys in two rotations: one for the giant step and another for the baby step (the server repeatedly uses the same keys for all other rotations). Furthermore, we introduce an RNS-based plaintext decomposition method that maintains a constant key size while achieving high precision. Specifically, we select the modulus $t^{\prime}$ (which allows high precision without wrap-around; more details in the next section) as a multiple of two smaller NTT-friendly primes, $t_{0}$ and $t_{1}$. The server decomposes each database vector into two RNS components. The client decomposes the query into two RNS components, encrypts each separately using the same secret key $s$ but different plaintext modulus as $t_{1}$. The server performs exact dot-product computations using each query ciphertext, and the server returns both of them to the client. After decryption, the client reconstructs the original plaintext in $Z_{t}$ using the Chinese Remainder Theorem (CRT). Our main observation is that the evaluation keys do not depend on the plaintext modulus used in the ciphertext; therefore, the server will use the same evaluation key for both queries.
    In metadata fetch, to expand a query, the server performs repeated substitution operations on the query ciphertext. Typically, the server requires one substitution key for each required power. However, we propose a method where the client sends keys for only a few substitutions, and the server computes the remaining substitutions using the same keys. We provide detailed explanations in the Appendix.
VII-B Embedding and Clustering
We map each document in the server database into dense embeddings that retain semantic information. For this, we use the sentence embedding model ‘msmarco-distilbert-base-tas-b’ with 66 million parameters and optimized for semantic search [^3] https://huggingface.co/sentence-transformers/msmarco-distilbert-base-tas-b. In particular, we use the implementation provided by the ‘sentence-transformers’ Python library [^4] https://www.sbert.net. The model outputs dense 768-dimensional floating point vectors. We then use PCA to reduce the dimensionality of each vector to 192.
The next step is clustering these vectors. To do so, we first find $K$ cluster centroids by running k-mean clustering. Then we assign each embedding the cluster with the nearest centroid. Finally, all points and their assigned cluster are compiled into an index.
We implemented this pipeline using the open-source library FAISS. FAISS provides fast, GPU-based algorithms for k-means clustering and indexing. For a comprehensive list of optimizations implemented by FAISS, refer to https://github.com/facebookresearch/faiss/wiki#research-foundations-of-faiss. FAISS exposes a flexible clustering API that allows us to experiment with various clustering parameters, such as the number of data points sampled, the number of iterations of the basic k-means algorithm, and whether to run on GPU or CPU. For indexing, we utilize one of the Flat indexes provided by FAISS to perform fast intra-cluster nearest neighbor search.
VII-C Parameters Selection
SHE parameters.
Performance of Wally directly depends on our choice of BFV parameters. We tried multiple parameter sets and selected the one that yielded the optimal performance. For search and metadata fetch, the selected parameters are as follows: polynomial degree $n=4096$, ciphertext modulus of $\log_{2}{Q}=83$ bits, with CRT limbs of $27,28$  and $28$ bits. We sample error from a centered binomial distribution with standard deviation $3.2$ and secret key from a ternary distribution modulus $Q$. This achieves 128 bits of quantum security.
For search, we set plaintext modulus of $\log_{2}{t}=16$ bits. Recall that the embedding model outputs floating-point vectors of dimension $d$, however BFV operates over integer vectors. We convert each float into a fixed-point integer by multiplying it by the appropriate scaling factor $p$ and then rounding down to the nearest integer. We select $p$ to ensure the dot products do not wrap around modulo $t$. We say that precision is equal to $\log_{2}(p)$ bits. To avoid wrap-around, it is important that $p<\sqrt{(t-1)/2}-\sqrt{d}/2$, which implies that $\log_{2}(p)<1/2\log_{2}t\approx 7$. Therefore, we obtain at least 7 bits of precision. To increase the precision, we use the plaintext CRT decomposition. For this, we pick two plaintext CRT limbs of plaintext modulus of $\log_{2}{t_{0}}=16\text{ and }\log_{2}{t_{1}}=17$ bits. Total plaintext bits are 31 bits, which gives us a precision of 15 bits without increasing evaluation key size.
As mentioned above, we reduce the size of the response ciphertext by first mod switching down to the last CRT limb of 27 bits and then further dropping the least significant bits. To avoid decryption error due to dropping bits, we set $l_{0}=9$ and $l_{1}=0$. This drops the size of the single response ciphertext from 55 KB to 23.5 KB.
For MulPir to handle larger noise growth, we use a plaintext modulus of $\log_{2}{t}=5$ bits. Having a smaller plaintext modulus allows us to set $l_{0}=19$ and $l_{1}=13$, which reduces each response ciphertext to 12 KB.
Epoch and number of users.
As described in Section III, Wally operates in discrete epochs. In each epoch, a total of $U$ users submit queries. However, malicious users can undermine Wally’s privacy guarantees by refraining from submitting fake queries. For our privacy analysis, we assume that half of the total clients in any epoch behave maliciously. We compute the privacy loss by excluding fake queries from these malicious clients. In other words we consider worst-case behavior of all the malicious clients.
Longer epochs lead to a larger $U$, which in turn reduces the number of fake queries required per user. However, extending the epoch duration may negatively impact service quality, since some queries might only be processed at the epoch’s end. For our evaluation, we use a epoch size of one and ten minutes and assume that at least 100,000 and 500,$0$ users submit queries during each epoch, of which half are honest. Given that search engines like Google and Bing process millions of queries per minute and serve billions of users, 500k is a conservative estimate in the context of modern search engines [41].
We also assume that each user will participate in around 400 epochs, one or two epochs each day of a year.
Differential privacy parameters.
For a fixed failure probability $\delta$, a smaller $\epsilon$ offers stronger privacy but compromises performance since each client must execute more fake queries. We set $\delta=2^{-30}$, ensuring that DP will fail only once in a billion queries, and $\epsilon=1/800$, providing a reasonable trade-off between privacy and performance. In each epoch, we assume that each participating client makes $\Delta$ queries. These queries represent the number of nearest clusters fetched for each client’s query. As mentioned in Section VI, this guarantees $(1/400,2^{-26})$ differential privacy per epoch. Consequently, the total differential privacy for each client across 400 epochs is $(1,2^{-26})$.
For each cluster, a client samples fake queries from $\mathcal{NB}(r/U,p)$. Which means in expectation the number of fake queries is
$\mathbb{E}[\text{Fake queries per client}]=\frac{rpK}{(1-p)U}$ (1)
$=\frac{3(1+\log_{2}(1/2^{-30}))e^{-0.0005/\Delta}K}{1-e^{-0.0005/\Delta}U}$ (2)
where $K$ is the total number of clusters and $U$ is the total number of participating honest clients. A little bit work shows that the above equation is always $\leq 102,700\Delta K/U$. As mentioned above considering honest users $U=250,000$, a modest estimate given modern search engines’ scale [41]. Therefore number of fake queries is upper bounded by $\leq 0.205\Delta K$
Search parameters and metric
The product of clusters $K$ and the number of clusters probed $\Delta$ determine the performance and accuracy of Wally. A smaller $K$ and a smaller $\Delta$ both enhance performance by reducing the overhead of fake queries. Further notice that very small $K$ will result in large response size. Similarly, reducing $\Delta$ may lead to a substantial drop in accuracy due to the increased probability of missing the true nearest neighbor.
Our objective is to minimize the fake queries overhead, which is proportional to the product $K\Delta$, while balancing search quality and response size. However, to achieve this using exhaustive search over a candidate set of parameters would require considerable time.
We experimented with different parameters and selected those that provide best accuracy while keeping $K$ moderately small and minimizing $\Delta$ as much as possible. For each probe, the client generates a new query with a large evaluation key. Consequently, large $\Delta$ will result in large total request size.
Dataset
We use the MSMARCO document ranking dataset  [42]. The dataset contains approximately 3.2 million passages, where each passage comprises a document identifier, document URL, document title, and document body.
VIII Evaluation
In this section, we demonstrate Wally’s concrete performance and compare it with baselines.
Experimental setup.
We ran server-side computations on an Intel Xeon w3-2423 instance with 32GB of RAM and 6 cores.
Metrics
To compare the performance, we use following metrics
  • Queries per second (QPS): this metric reflects how many queries the server can process in a second for fixed infrastructure. We assume that the server has at least 10,000 cores. We measure the latency for a single query on a single core and divide the total number of cores by that number.
  • Bandwidth: We calculate the size of the client query and the server response for each query.
  • Accuracy: To evaluate the accuracy of our search results, we used the Mean Reciprocal Rank (MRR@100) metric, a standard measure for assessing the quality of search results. Given a query $q$ and a ground truth index $i$ such that $DB[i]$ is the most relevant entry, we merge all scores (along with entry indices) obtained from probing $\Delta$ clusters, sort them, and retain the first 100 scores. If the list contains $i$ at the j-th rank (where $j\leq 100$), the reciprocal rank score is calculated as $1/j$. In the absence of $i$ in the list, MRR@100 is set to 0.
Baselines.
We compared Wally with following baselines. Both baselines offer full obliviousness but have high performance overhead:
  • Tiptoe (simulated) [14], which is a state of the art private search system based on linear homomorphic encryption (LWE) . Similar to Wally, Tiptoe uses a clustering approach to find the nearest candidate. However for each query they only probe a single cluster. Unfortunately, due to resource constraints we are unable to run the complete Tiptoe system. We therefore simulate it by using following estimates
    • Latency: For latency we estimate the minimum time the server will take to process a single query. We exclude network cost, which gives advantage to Tiptoe as their network bandwidth is quite high. For each query, the computation is dominated by dot-products between vectors. The server first generates a token by multiplying RLWE encryption of secret key with precomputed LWE ciphertext. For this $16\sqrt{N}$ dot products are performed between vectors of dimension $n=2048$. Note that even though this computation can be performed in offline phase, the server still has to perform it for each query. Then to privately fetch and rank candidates from a particular cluster, the server performs $N$ dot-products between vectors of dimension $d$. For each vector we assume that each element is of 32 bits.
    • Accuracy: To simulate their accuracy results we divide the database into $\sqrt{N}$ clusters, probe a nearest cluster and then calculate MRR@100 of the resulting candidates.
    • Communication: To estimate communication we calculate minimum number of 32 bit elements the client and server have to transfer per query. To generate a token, the client sends at least $n=2048$ RLWE ciphertexts or $n^{2}$ elements. For ranking, the client sends a query consisting of $d\sqrt{N}$ elements. For tokens, the server response consists of $16$ RLWE ciphertexts (each consisting of at least 2048 elements) and for ranking the response consists of $\sqrt{N}$ elements.
  • Pacmann, [43] which is a concurrent work. Like Wally, this scheme also employs sublinear server computation, but it offers full obliviousness. Pacmann’s approach differs from Wallyin two key aspects. Firstly, to locate the nearest neighbor, Pacmann utilize a graph-based search algorithm. Secondly, Pacmann employs sublinear Offline/Online PIR to retrieve graph nodes and traverse the graph.
    For each query, the scheme necessitates multiple round trips and per-query maintenance. We compare our results with the ones provided in the Pacmann paper [43], which is sufficient since we conducted our own experiments in the matching environment.
For Wally, we calculate the QPS and bandwidth by summing the cost of $\Delta$ real queries and fake queries.
Comparison with Tiptoe.
The results are presented in Table II. Wally significantly outperforms Tiptoe across all metrics. Notably, Wally’s QPS is 11~27 times better, this is because for each query the server has to perform at least one expensive operation per entry in the database. Similarly, the communication per query is 10~34 times smaller than Tiptoe, for varying numbers of clusters probed. Tiptoe can only probe a single cluster per query, and to probe more clusters, a new query must be initiated. For $\Delta=1$, a single probe case, Wally’s MRR100 score is only slightly better than Tiptoe’s score of 0.11. However, as we increase $\Delta=5$, Wally’s MRR100 score improves to 0.18, which is considerably better than Tiptoe’s score.
Comparison with Pacmann.
As given in the table, Pacmann’s scheme has better QPS and MRR100 than Wally. Specifically, the better QPS is because Pacmann is based on offline/online sublinear PIR and their MRR100 is better due to use of graph based search data structure which is superior than clustering technique. However, their scheme requires large client storage and communication per query. Specifically, for 3.5 million database, the client is required to store a hint of 614 MB and per query communication is 61.6 MB, which is prohibitively high, specifically for cellular networks. Another downside is that hint-based protocol can not directly handle database updates. Finally, to gain better MRR100 the scheme requires multiple round trips which adds to their network delay. We did not consider the cost of network delay in their QPS calculation. Specifically, due to their use of offline/online PIR, even though QPS of Pacmann is 1.35.6×1.3$\sim$ 5.6$\times$1.3 ∼ 5.6 × better than Wally but the communication is 23123×23$\sim$ 123$\times$23 ∼ 123 × as large for varying $\Delta$.
Client Storage (MB) Communication (MB) Queries Per Second (QPS) MRR100
Tiptoe [14] 0.61 17.4 909 0.11
Pacmann [43] 614 61.6 34,482 0.26
Wally ($K=256$, $\Delta=1$) 0.04 0.56 25,974 0.12
Wally ($K=256$, $\Delta=3$) 0.04 1.7 9,881 0.16
Wally ($K=256$, $\Delta=5$) 0.04 2.6 6,667 0.18
TABLE II: Comparison of Wally  Tiptoe and Pacmann for MSMARCO [42] database. Communication and QPS for Wally includes overhead due to fake queries. Worst value in each metric is marked with a red color.
Benchmarking Wally.
In Table III, we evaluate the performance of Wally across various database sizes ranging from one million entries to hundred million entries and cluster sizes ranging from 128 to 1,024. For all these experiments, we assume that the client probes the nearest cluster $\Delta=1$. Additionally, we consider epoch sizes of one and ten minutes, with 100,000 and 500,000 thousand honest users, respectively.
Firstly, we notice that increasing the epoch size (i.e., the number of honest users) leads to improved overall performance across all database sizes. Specifically, the QPS increases by 2-3 times, while the request and response sizes improve by 1.5-2.4 times. This demonstrates that Wally is an ideal choice for applications where moderate to extended delays are acceptable.
Secondly, we observe that the expected number of fake queries is nearly proportional to the product of the number of honest users ($U$) and inverse of the number of clusters ($K$). For each database, as the epoch size increases from one minute to ten minutes, we increase the number of users by 5×5$\times$5 × (which reduces fake queries) but also increase the number of clusters by 2×2$\times$2 × (which increases fake queries). Consequently, we see that fake queries almost decrease by a factor of $5/2=2.5$, indicating a linear dependency. Alternatively, we can consider that reducing the number of clusters $K$ for a fixed database will lead to a decrease in fake queries.
Thirdly, we observe that as the database size varies, the request and response sizes linearly increase. However, the QPS decreases. This is because for larger databases, each cluster size becomes substantial, resulting in increased server operations per query.
Entries in database N 1 Million 16 Million 100 Million
Epoch size 1 Minute 10 minutes 1 Minute 10 minutes 1 Minute 10 minutes
Number of users $U$ 100,000 500,000 100,000 500,000 100,000 500,000
Number of clusters (K) 128 256 256 512 512 1,024
Expected fake queries 1.7 0.3 3.3 1.7 6.7 2.6
Request size (MB) 0.76 0.36 1.18 0.76 2.17 1.01
Response size (MB) 0.17 0.04 2.3 1.5 13.9 3.03
Queries per second (in thousand) 222 666 14 32 4.7 13.9
TABLE III: Performance of Wally across varying database sizes. Each client will probe a single cluster $\Delta=1$. Each setting provides $(\epsilon=1,\delta=2^{-26})$-DP guarantee.
Number of clusters (K) 64 128 256 512
Expected fake queries 0.3 0.6 1.2 2.4
Request size (MB) 64 128 256 512
Response size (MB) 64 128 256 512
QPS (in thousand) 64 128 256 512
TABLE IV: Performance of Wally for different values of $K$. Each epoch is of ten minutes and database has one million entries. Number clusters probed $\Delta=3$. Each setting provides $\epsilon=1,\delta=2^{-26}$ DP guarantee.
Clusters probed ($\Delta$) 1 3 5 10
Expected fake queries 0.3 0.6 0.8 1.3
Request size (MB) 0.36 0.45 0.59 0.65
Response size (MB) 0.08 0.10 0.11 0.14
QPS (in thousand) 370 312 277 217
TABLE V: Performance of Wally for different values of $\Delta$. Each epoch is of ten minutes and database has one million entries. Number of clusters $K=125$. Each setting provides $\epsilon=1,\delta=2^{-26}$ DP guarantee.
IX Related Work
Early private search systems.
Early private search systems heuristically add dummy queries or obfuscate queries by changing the query client-side, but they ultimately send queries in the clear to the server [44, 45, 46, 47, 6, 48, 49, 50, 51]. These systems are broken by a semi-honest server, i.e., the server alone can de-anonymize queries [52, 53, 54, 55] or they have strict distribution requirements on queries, e.g., i.i.d. queries [47]. This contrasts with Wally  which has a provable differential privacy guarantee: the server’s view between the histograms over clusters with and without a client’s queries is differentially private. Further, the client’s queries in Wally are encrypted under homomorphic encryption to prevent the trivial attack shown in [26, Section 3].
Homomorphic encryption.
Homomorphic encryption alone is a popular technique which has been used in private search systems [14, 56, 57, 58]. A major downside of these systems is that they require encrypted computation over the entire database, whereas Wally computes on a few clusters per query. Wally achieves this by utilizing differential privacy to hide the requested cluster index from the server, which makes Wally practically deployable in settings with many honest clients. On the other hand, these previous systems based only on homomorphic encryption have the strongest possible privacy notion: fully obliviousness from the server’s point of view assuming standard cryptographic hardness assumptions. An in-depth comparison with Tiptoe [14] in Section I. A related system is presented in [6] which sends text queries in the clear but uses additively homomorphic encryption [59] to compute an encrypted similarity score. The queries are noised, or “embellished”, beforehand by replacing each word in the query with a set of unrelated words. Embellishing heuristics provide no provable security guarantee, unlike the differential privacy guarantee in Wally.
Secure multiparty computation.
Secure multi-party computation (MPC) is a common technique for private search systems. These systems incur large query sizes or high latency due to many communication rounds. SANNS [60] uses garbled circuits and homomorphic encryption for private nearest neighbors search in the semi-honest client-server model. As a result, they have complete cryptographic obliviousness for each query and the server’s database entries are private besides the query’s response, but each query sends gigabytes of data for databases of 1M entries or more due to their use of garbled circuits [60, Table 1]. Preco [61] is a two-server maliciously-secure MPC private nearest neighbor search system which requires an linearly-sized computation by each server ($N$), a $\sqrt{N}$ computation by the client, and $O(N)$ communication for each query [61, Table 1]. They achieve fully cryptographic obliviousness via distributed point functions [62] and probabilistic batch codes [63]. Concretely, they achieve 10MB of communication and 10s latency for a database with 10M entries.
Differential privacy.
Differentially-private multi-server (information-theoretic MPC) PIR schemes [26, 27] are not immediately applicable to private search since PIR itself only supports exact matches, but these works’ use of DP is related to our use of DP since they use DP to lower their online computation while leaking information to the servers. In both works, servers learn differentially private, noised histograms of queries over database entries. Note, this compares to Wally where the server learns noised histograms of queries over clusters and only requires one server. The latter [27] is more efficient and requires a multi-server PIR over the entire database once on the database for a batch of queries. Notably, [27] can serve one million queries in about two seconds on a 100K entry database with eight servers in a dishonest majority security model and a ten minute pre-computation. Their offline and online computation grows linearly with the size of the server’s database. Wally, on the other hand, has no offline computation and is sublinear in the database size.
Anonymization networks.
Anonymization networks and mix-nets [64, 24] anonymize the user’s identity but send the query in the clear. Only applying these methods to search systems would not hide users’ private information and are broken by a semi-honest server [51].
Trusted hardware.
Trusted hardware is another tool used in private search systems [65, 66, 67]. However, trusted hardware search systems are only as strong as their underlying hardware [68, 69, 70, 71, 72].
Concurrent works.
Pacmann [43] efficiently reduces graph-based private nearest neighbors search to many PIR queries on a graph. Pacmann uses offline-online PIR which requires the client to stream the database beforehand, then each PIR computation is sublinear since it uses a recent sublinear PIR protocol [73]. As a result, it is a cryptographic, fully-oblivious private search system and offers strong accuracy results from their use of graph-based nearest neighbors search. However, Pacmann requires each client stream the entire database offline whereas Wally requires no offline computation. Further, Pacmann requires many rounds per query, unlike Wally which only requires one round for small database entries and two rounds for large database entries.
X Conclusion
Wally demonstrates that private search systems can be scalable to systems with many users. Wally carefully balances somewhat homomorphic encryption together with differential privacy to hide the client’s query and reveal a differentially private histogram over all clients’ traffic to database clusters. Further, Wally’s overhead from differential privacy vanishes as the number of users in the system increases. This vanishing overhead is crucial to scalability and is potentially useful in many private systems deployable at-scale.
Appendix A Optimizations
Here we present our optimizations for both use cases: private nearest neighbor search and exact retrieval (PIR). Our optimizations revolve around the BFV [29, 17] and BGV [18] SHE schemes. Further, we describe several optimizations to the MulPIR scheme [33]. First, we describe the standard RNS optimization in SHE [74] and a method to compress ciphertexts by dropping LSBs as described in Cheetah [75].
Often $Q$ is larger than machine-size words and we use $Q$ as the product of machine-sized primes $Q=q_{0}q_{1}\cdots q_{l}$. This allows us to store a ciphertext as two polynomials each stored as an $n\times(l+1)$ matrix of integers by the Chinese remainder theorem (CRT) on $Q$. Further, we require each $q_{i}$ to be NTT friendly, $q_{i}=1\mod 2n$. This allows us to use the number theoretic transform (NTT) to efficiently switch between each evaluation form and coefficient form in time $O(n\log_{2}n)$, since the NTT is a modular version of the fast Fourier transform over a prime modulus. We can efficiently compress a ciphertext by modulus-switching down to a smaller modulus, $Q^{\prime}$ where $Q^{\prime}|Q$, e.g., $Q^{\prime}=q_{0}$, without affecting the ciphertext’s noise budget.
Dropping ciphertext LSBs.
We use Cheetah’s [75] method to compress ciphertexts after homomorphic computation: we modulus switch down to the smallest RNS modulus, $Q^{\prime}=q_{0}$, then drop the least significant bits of the ciphertext. This technique compresses the response size further than simply modulus-switching down, but it also adds noise. Cheetah incorrectly models the ciphertext as noiseless (Appendix F in [40]), but we derive the proper analysis as follows. Dropping the LSBs of a BFV ciphertext $\mathsf{ct}=(c_{0},c_{1})$ can be analyzed by expressing the polynomials as high and low bit polynomials $c_{0}=2^{b_{0}}c_{0}^{h}+c_{0}^{l},\ c_{1}=2^{b_{1}}c_{1}^{h}+c_{1}^{l}$. Then, the new decryption equation is
$2^{b_{0}}c_{0}^{h}+2^{b_{1}}c_{1}^{h}s=\lfloor Q/t\rceil m+e+e_{drop}$
for
$e_{drop}=-c_{0}^{l}-c_{1}^{l}s.$
We have correctness as long as $\|e+e_{drop}\|_{\infty}<Q^{\prime}/2t$. Concretely, we are able to drop around $8$ kB from the response size in private nearest neighbor search to a response size of $47$ kB when combined with our BFV plaintext CRT optimization described in the next subsection.
A-A Private Nearest Neighbor Search Optimizations
The homomorphic computation in the private nearest neighbor search use case is $C$ inner products over real embedding vectors satisfying $\mathbf{e}$, normalized to $\|\mathbf{e}\|_{2}=1$ in our use case, where $C$ is the maximum cluster size. We scale $\mathbf{e}$’s entries to integers, round them to a vector $\mathbf{\widetilde{e}}$, and perform the inner product over the integers. The computation is correct as long as the inner product does not wrap around mod $t$. We define plaintext precision as $\approx\log_{2}(t)/2$ since the inner product has a multiplicative depth of one.
Minimal rotation keys and bandwidth.
Wally uses the “baby-step giant-step” (BSGS) optimization [32] to minimize the number of rotation keys: for $d$-dimensional inner products, BSGS requires $2\sqrt{d}$ rotations for step sizes of $1$ (baby) and $\sqrt{d}$ (giant). See Algorithm 1 in [32] for the BSGS algorithm. Additionally, we are able to use two rotation keys since $d<n/2$ in our use case. This optimization leads us to a query size of one ciphertext $\mathsf{ct}$ and two rotation keys $\mathsf{evk}=(\mathsf{evk}_{1},\mathsf{evk}_{\sqrt{d}})$. For key-switching, and relinearization in PIR below, we use the hybrid-GHS [74] key-switching strategy described in-detail by Kim et al. [39].
Say we require a plaintext modulus of $t\approx 2^{29}$ for $\sim 14$ bits of plaintext precision. The smallest BFV ring dimension supporting this is $n=8192$ which requires a evaluation key modulus of $Q\approx 2^{165}$, ciphertext modulus of $Q^{\prime}\approx 2^{110}$ for 128 bits of post-quantum security [76] [^5] https://github.com/malb/lattice-estimator. A query consists of $2$ polynomials for the ciphertext and $2$ polynomials for the evaluation keys, or $225$ kB and $676$ kB respectively, for a total query size of $901$ kB for these parameters.
Re-using BFV evaluation keys via the plaintext CRT.
Wally achieves minimal query bandwidth per client query by a novel use of the Chinese remainder theorem (CRT) on the BFV plaintext space. The CRT on plaintexts allows increase the plaintext precision without increasing evaluation key size since BFV’s evaluation keys are independent of the plaintext modulus. Therefore, we encrypt a vector $\mathbf{\widetilde{v}}$ as $\mathbf{\widetilde{v}}\mod t_{0}$ and $\mathbf{\widetilde{v}}\mod t_{1}$ for NTT-friendly primes $\{t_{0},t_{1}\}$. The computation is correct if the inner products do not wrap around modulo $t=t_{0}t_{1}$, but the individual ciphertexts can have their computation wrap around modulo each $t_{i}$.
In general, plaintext CRT is preferable to increasing the ring dimension since the evaluation key sizes dominate the ciphertext size: 6 polynomials versus 2 polynomials. This is true in the bandwidth-optimal key-switching strategy for hybrid GHS key-switching. Another advantage of the plaintext CRT is that it allows for very high precision computations which do not wrap around mod $t$. One limitation of this optimization is that we run out of NTT-friendly plaintext primes for small parameter sets, like $n=4096$ or $2048$.
A-B PIR Optimizations
Uneven dimensions.
In general, the cost of a ciphertext-ciphertext multiplication is much higher than a ciphertext rotation and both are much higher than addition or plaintext multiplication. For example, Table I shows ciphertext addition and plaintext multiplication as less than $100\mu s$ while ciphertext rotation is $0.5ms$ and ciphertext multiplication is $2.5ms$.
Therefore, we structure the database (cluster) as a rectangle $\mathbf{D}\in R_{t}^{d_{1}\times d_{2}}$ with uneven dimensions $d_{1}$ and $d_{2}$. For any fixed HE parameter set, the server’s compute time is dominated by ciphertext multiplications and rotations and that the former is a constant factor $\gamma$ more than the latter. This simplification yields a compute time of $2d_{1}t_{r}+2d_{2}t_{r}+d_{2}t_{\times}$ where $t_{\times}$ and $t_{r}$ are the times to multiply and rotate ciphertexts, respectively. More simply, the total compute time, measured in terms of rotations, is
$(\gamma+2)d_{1}+2d_{2}=(\gamma+2)C/d_{2}+2d_{2}$
rotations. Minimizing this as a function in $d_{2}$ gives $d_{2}=\sqrt{1+\gamma/2}\sqrt{C}$.
This optimization saw an improvement of $30-40\%$ in MulPIR run times. We chose this concrete analysis since SHE parameter regimes are restricted per use case and it is much simpler than an asymptotic analysis.
Linearizing parts of query expansion
MulPIR is an optimized query expansion which originated with Angel et al. [63]. In short, a query index is encoded in a plaintext polynomial then expanded into $k$ ciphertexts by calling $k$ ciphertext rotation operations (Galois automorphisms followed by key-switching). We noticed that some of these expansions can be substituted with linear operations: call $\mathsf{ct}\pm\mathsf{ct}^{\prime}$ where $\mathsf{ct}^{\prime}$ is $\mathsf{ct}$ rotated. In general, we saw a $10-25\%$ improvement in MulPIR’s expansion from our optimization.
Lazy rescaling in BFV multiplication.
We apply the lazy rescaling technique of Kim et al [39] in MulPIR’s last step, the inner product between ciphertexts $\mathbf{r}^{T}\mathbf{a}$. The main idea here is that BFV multiplications first multiply polynomials over the integers, then scale and round back to integers modulo $Q$. Kim et al [39, Appendix F] noticed that one can multiply over the integers, add over the integers, then scale and round after the additions, from the dimension number of scaling operations to one. Lazy rescaling yielded a $15-20\%$ improvement in our MulPIR implementation.
Optimization for very large entries
In MulPIR the database is represented as a plaintext matrix $\mathbf{D}\in R_{t}^{d_{1}\times d_{2}}$. Conventionally, the response is computed as
[…]
This approach results in $d_{1}\cdot d_{2}\cdot n$ ciphertext-plaintext multiplications and $d_{2}\cdot t$ ciphertext-ciphertext multiplications, where $t$ is the number of plaintexts used to encode one entry. The number of ciphertext-ciphertext multiplications grows linearly with $t$ and remains a concretely small value when entries are not large. However, it will blow up with extremely large entries (e.g., when they are high-resolution photos).
To overcome this challenge, we can swap the order of the computation as
[…]
The outer-product takes $d_{1}\cdot d_{2}$ ciphertext-ciphertext multiplications to compute, which is independent of the entry size, while the inner products still take $d_{1}\cdot d_{2}\cdot n$ ciphertext-plaintext multiplications. Therefore, this optimization can reduce the computation overhead when the entry size ($t$ times the plaintext size) is larger than the first dimension size $d_{1}$.
Keyword PIR.
MulPIR’s keyword PIR functionality is given by using cuckoo hashing to find the entry’s proper index. We noticed for our cluster sizes $C$, we are able to split the cluster into two hash tables and perform $1$-hash cuckoo hashing on each without increasing the (empirical) failure probability. This resulted in a 2×2$\times$2 × improvement in query size and server computation time.
Low-level SHE Optimizations
BFV’s multiplication step over the integers is done by extending the modulus $Q$ to a larger modulus $PQ$ so the multiplication does not wrap around. We use Bajard et al.’s method to extend the basis, “fast basis conversion” in [37] and re-use the basis elements in $Q$ for a $\sim 15\%$ improvement in BGV multiplication.
We add a conditional, lazy modular reduction to Harvey’s NTT [38]. The main idea is that we only reduce during the butterfly if the integers would exceed the machine word size. This saw a $20-25\%$ improvement in forward NTT times and $15-20\%$ improvement in inverse NTT times for $61$-bit primes. Note, we saw no improvement on primes near $64$-bits.